### v1

# Minimizing movements for quasilinear Keller–Segel systems with nonlinear mobility in weighted Wasserstein metrics

Abstract
We prove the global existence of weak solutions to quasilinear Keller–Segel systems with nonlinear mobility by minimizing movements (JKO scheme) in the product space of the weighted Wasserstein space and $L^{2}$ space. While minimizing movements for Keller–Segel systems with linear mobility are adapted in the product space of the Wasserstein space and $L^{2}$ space, due to the nonlinearity of mobility, we need to use the weighted Wasserstein space instead of the Wasserstein space. Moreover, since the mobility function is not Lipschitz, we first find solutions to the Keller–Segel systems whose mobility is apporoximated by a Lipschitz function, and then we establish additional uniform estimates and convergences to derive solutions to the Keller–Segel systems.
Key words and phrases: Minimizing movements, JKO scheme, Weighted Wasserstein distance, Keller–Segel, Gradient flows, Nonlinear mobility.
1 Introduction
We consider the following parabolic system:
$\begin{cases}\partial_{t}u=\Delta u^{p}-\nabla\cdot(\chi u^{\alpha}\nabla v)&\mathrm{in}\ \Omega\times(0,\infty),\\ \partial_{t}v=\Delta v-v+u&\mathrm{in}\ \Omega\times(0,\infty),\\ \nabla u\cdot\boldsymbol{n}=\nabla v\cdot\boldsymbol{n}=0&\mathrm{on}\ \partial\Omega\times(0,\infty),\\ u(\cdot,0)=u_{0}(\cdot),\ v(\cdot,0)=v_{0}(\cdot)&\mathrm{in}\ \Omega,\end{cases}$ (1.1)
where $p\geq 1,\ 0<\alpha<1,\ \chi>0,\ d\geq 2$, $\boldsymbol{n}$ is the outer unit normal vector to $\partial\Omega$ and $\Omega$ is a bounded convex domain in $\mathbb{R}^{d}$ with smooth boundary. In addition, $u_{0}\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega)$ is a nonnegative function and $v_{0}\in H^{1}(\Omega)$ is also a nonnegative function, where $\mathcal{P}(\Omega)$ is the set of Borel probability measures on $\Omega$.
The Keller–Segel system is a model to describe an aggregation phenomenon of cellular slime molds with chemotaxis, where we denote by $u$ the cell density and by $v$ the concentration of the chemical attractant. We thus consider nonnegative solutions to (1.1). In order to take into account a volume-filling effect, the exponent $\alpha\in(0,1)$ is introduced, where the volume-filling effect is the phenomenon that the movement of cells is restricted by the presence of other cells. There are various mathematical analyses about (1.1) (see [20, 23, 22, 11, 21, 24]).
The purpose in this paper is to find global weak solutions to (1.1) by regarding the system (1.1) as a gradient flow of a suitable functional in a suitable metric space. In more detail, we use the time discrete variational method called minimizing movements (JKO scheme [13]). When $\alpha=1$, the system (1.1) can be seen as a gradient flow of the energy functional
$\tilde{E}(u,v)\coloneq\frac{1}{\chi(p-1)}\int_{\Omega}u^{p}\,dx-\int_{\Omega}uv\,dx+\frac{1}{2}\int_{\Omega}\left(|\nabla v|^{2}+v^{2}\right)\,dx$
in the product space of the Wasserstein space and $L^{2}$ space (see [4, 3, 18, 19]). Here, Wassrestein space is the metric space of Borel probability measures with finite second moment $\mathcal{P}_{2}(\Omega)$ endowed with the Wasserstein distance
$W_{2}(\mu_{0},\mu_{1})^{2}\coloneq\inf_{\gamma\in\Gamma(\mu_{0},\mu_{1})}\int_{\Omega\times\Omega}|x-y|^{2}\,d\gamma(x,y)\quad\mathrm{for}\ \mu_{0},\mu_{1}\in\mathcal{P}_{2}(\Omega),$ (1.2)
where $\Gamma(\mu_{0},\mu_{1})$ is the set of $\gamma\in\mathcal{P}(\Omega\times\Omega)$ satisfying
$\gamma(A\times\Omega)=\mu_{0}(A),\ \gamma(\Omega\times A)=\mu_{1}(A)\ \mathrm{for\ all\ Borel\ set}\ A\subset\Omega.$
In [4], Blanchet and Laurençot showed the global exsitence of weak solutions to the Keller–Segel system with $\alpha=1,p=2-2/d$ and small initial data in $\Omega=\mathbb{R}^{d}\ (d\geq 3)$ by minimizing movements. In [18] and [19], Mimura proved the global existence of weak solutions to the Keller–Segel system with $\alpha=1$ and $p\geq 2-2/d$, adding the assumption of small initial data if p=22/dp=2-2/d, in bounded smooth domain $\Omega\subset\mathbb{R}^{d}\ (d\geq 3)$ by minimizing movements.
However, since the mobility function $u^{\alpha}\ (0<\alpha<1)$ is nonlinear, the system (1.1) cannot be seen as a gradient flow of a corresponding energy functional in the product space of the Wasserstein space and $L^{2}$ space. We thus change the Wasserstein space to the weighted Wasserstein space (see [8] and Section 2.2), which is the extension of the Wasserstein space in some sense. Then we will see the system (1.1) as a gradient flow of the energy functional
$E(u,v)\coloneq\frac{p}{\chi(p-\alpha)(p+1-\alpha)}\int_{\Omega}u^{p+1-\alpha}\,dx-\int_{\Omega}uv\,dx+\frac{1}{2}\int_{\Omega}\left(|\nabla v|^{2}+v^{2}\right)\,dx$
in the product space of the weighted Wasserstein space and $L^{2}$ space.
Minimizing movements with the Wasserstein space have been studied a lot, thus the methods of a improved regularity of minimizers and convergences to weak formulation are established (for instance [4, 3, 17, 5, 7, 13, 18, 19]). On the other hand, minimizing movements with the weighted Wasserstein space are used only a few papers ([15, 25, 26]). In [15] and [25], they deal with some fourth-order partial differential equations like the Cahn–Hilliard type equation and mainly with the Lipschitz mobility. While the mobility function $u^{\alpha}$ in (1.1) is not Lipschitz, we show that minimizing movements with the weighted Wasserstein space can be also used for the type of the equations (1.1). Set $X\coloneq(L^{p+1-\alpha}\cap\mathcal{P}(\Omega))\times H^{1}(\Omega)$.
Theorem 1.1.
Let $\alpha\in(0,1),\ p=1+\alpha,\ \chi>0,\ d\geq 2$ and $(u_{0},v_{0})\in X$ be a pair of nonnegative functions. Then for all $T>0$, there exists a nonnegative weak solution $(u,v)$ to (1.1) on the time interval $[0,T]$ satisfying
$\bullet u\in L^{\infty}((0,T);L^{2}(\Omega))\cap L^{2}((0,T);H^{1}(\Omega)),$
$\bullet\|u(t)\|_{L^{1}(\Omega)}=1\quad\forall t\in[0,T],$
$\bullet v\in L^{\infty}((0,T);H^{1}(\Omega))\cap L^{2}((0,T);H^{2}(\Omega))\cap C^{\frac{1}{2}}([0,T];L^{2}(\Omega)),$
$\bullet\lim_{t\to 0}W_{m}(u(t),u_{0})=0\ \mathrm{and}\ \lim_{t\to 0}\|v(t)-v_{0}\|_{L^{2}(\Omega)}=0,\ \mathrm{where}\ m(u)=u^{\alpha},$
and
$\int_{0}^{T}\int_{\Omega}u^{\alpha}[(1+\alpha)\nabla u-\chi\nabla v]\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u_{0}-u(\cdot,T))\varphi\,dx,$
$\int_{0}^{T}\int_{\Omega}\left[\nabla v\cdot\nabla\zeta+v\zeta-u\zeta\right]\,dx\,dt=\int_{\Omega}(v_{0}-v(\cdot,T))\zeta\,dx,$
for all $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ and $\zeta\in H^{1}(\Omega)$.
Theorem 1.1 is only the case $p=1+\alpha$ for all $0<\alpha<1$. When $p<1+\alpha$, adding an assumption for $p$ and $\alpha$, we obtain weak solutions to (1.1).
Theorem 1.2.
Let $\chi>0,\ d\geq 2$ and $(u_{0},v_{0})\in X$ be a pair of nonnegative functions. Assume
$1+\alpha-\frac{2}{d}<p<1+\alpha,\ \alpha\in\bigg[\frac{1+p}{3},1\bigg).$
Then for all $T>0$, there exists a nonnegative weak solution $(u,v)$ to (1.1) on the time interval $[0,T]$ satisfying
$\bullet u\in L^{\infty}((0,T);L^{p+1-\alpha}(\Omega)),\ u^{\frac{p+1-\alpha}{2}}\in L^{2}((0,T);H^{1}(\Omega)),$
$\bullet\|u(t)\|_{L^{1}(\Omega)}=1\quad\forall t\in[0,T],$
$\bullet v\in L^{\infty}((0,T);H^{1}(\Omega))\cap L^{2}((0,T);H^{2}(\Omega))\cap C^{\frac{1}{2}}([0,T];L^{2}(\Omega)),$
$\bullet\lim_{t\to 0}W_{m}(u(t),u_{0})=0\ \mathrm{and}\ \lim_{t\to 0}\|v(t)-v_{0}\|_{L^{2}(\Omega)}=0,\ \mathrm{where}\ m(u)=u^{\alpha},$
and
$\int_{0}^{T}\int_{\Omega}(\nabla u^{p}-\chi u^{\alpha}\nabla v)\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u_{0}-u(\cdot,T))\varphi\,dx,$
$\int_{0}^{T}\int_{\Omega}\left[\nabla v\cdot\nabla\zeta+v\zeta-u\zeta\right]\,dx\,dt=\int_{\Omega}(v_{0}-v(\cdot,T))\zeta\,dx,$
for all $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ and $\zeta\in H^{1}(\Omega)$.
Remark 1.3.
In Theorem 1.2, we have the condition
$p>1+\alpha-\frac{2}{d}.$
We can see this exponent in the point of view of scalling. Let $(u,v)$ satisfies the system:
$\begin{cases}\partial_{t}u=\Delta u^{p}-\nabla\cdot(u^{\alpha}\nabla v)&\mathrm{in}\ \mathbb{R}^{d}\times(0,\infty),\\ \partial_{t}v=\Delta v+u&\mathrm{in}\ \mathbb{R}^{d}\times(0,\infty),\end{cases}$ (1.3)
then $(u_{\lambda},v_{\lambda})$, where $u_{\lambda}(x,t)=\lambda^{\frac{2}{1+\alpha-p}}u(\lambda x,\lambda^{2}t),v_{\lambda}=\lambda^{\frac{2(p-\alpha)}{1+\alpha-p}}v(\lambda x,\lambda^{2}t)$ for $(x,t)\in\mathbb{R}^{d}\times(0,\infty),\ \lambda>0$, also satisfies (1.3). By the first equation of (1.3), the $L^{1}$ norm of $u_{\lambda}$ is preserved for $t\in(0,\infty)$. Thus we focus the $L^{1}$ norm, then by the change of variables, we have
$\|u_{\lambda}(\cdot,t)\|_{L^{1}(\mathbb{R}^{d})}$ $=\int_{\mathbb{R}^{d}}\lambda^{\frac{2}{1+\alpha-p}}u(\lambda x,\lambda^{2}t)\,dx$
$=\int_{\mathbb{R}^{d}}\lambda^{\frac{2}{1+\alpha-p}-d}u(x,\lambda^{2}t)\,dx$
$=\lambda^{\frac{2}{1+\alpha-p}-d}\|u(\cdot,\lambda^{2}t)\|_{L^{1}(\mathbb{R}^{d})}\quad\mathrm{for\ all}\ \lambda>0,\ t>0.$
Hence the exponent that the $L^{1}$ norm is invariant by scalling is
$\frac{2}{1+\alpha-p}-d=0\Leftrightarrow p=1+\alpha-\frac{2}{d}.$
On the other hand, when $\alpha=1$, the case p>22/dp>2-2/d is called sub-critical case and it is known that the Keller–Segel system has global weak solutions in that case, which is proved by various ways including minimizing movements ([12, 4, 18, 19]). When $0<\alpha<1$, it is shown that system (1.1) has global weak solutions if $p>1+\alpha-2/d$ in [22] and [11], where minimizing movements are not used. Moreover if $p\geq 1$ and $\alpha\geq 1$ satisfy the condition $p<1+\alpha-2/d$, there exists a finite time blow-up solution of the Keller–Segel system ([10]). Hence from these facts, we may derive the proper condition for global existence of weak solutions to (1.1) by minimizing movements. Finally, the additional assumption for $\alpha$ in Theorem 1.2 is technical because it is not appeared in [22] and [11].
In the critical case $p=1+\alpha-2/d$, if we assume that $\chi>0$ is sufficiently small, which is equivalent to the smallness of the $L^{1}$ norm of the initial data, then there exist global weak solutions to (1.1).
Theorem 1.4.
Let $d\geq 3$ and $(u_{0},v_{0})\in X$ be a pair of nonnegative functions. Assume
$p=1+\alpha-\frac{2}{d},\ \alpha\in\bigg[\frac{d-1}{d},1\bigg).$
If $\chi>0$ is small enough then the same statement in Theorem 1.2 holds.
Remark 1.5.
In Theorem 1.4, the condition $(d-1)/d\leq\alpha<1$ comes from $(1+p)/3\leq\alpha<1$ in Theorem 1.2. If $d=2$ then $p=\alpha<1$, thus we exclude $d=2$ from the assumption in Theorem 1.4.
Remark 1.6.
In [24], it is shown that if the $L^{1}$ norm of the initial data is small enough then there exists a global solution to the Keller–Segel system (1.1) with $p=1+\alpha-2/d$. Note that minimizing movements are not used in [24].
Remark 1.7.
In Theorems 1.1, 1.2 and 1.4, by a little modification of the proof, it also holds that for all $s_{1},s_{2}]\subset[0,T]$,
$\int_{s_{1}}^{s_{2}}\int_{\Omega}(\nabla u^{p}-\chi u^{\alpha}\nabla v)\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u(\cdot,s_{1})-u(\cdot,s_{2}))\varphi\,dx,$
$\int_{s_{1}}^{s_{2}}\int_{\Omega}\left[\nabla v\cdot\nabla\zeta+v\zeta-u\zeta\right]\,dx\,dt=\int_{\Omega}(v(\cdot,s_{1})-v(\cdot,s_{2}))\zeta\,dx.$
Due to lack of the Lipschitz property of the mobility function $u^{\alpha}\ (0<\alpha<1)$, it is complicated for us to consider the Euler–Lagrange equation. When $\alpha=1$, the mobility function $u$ is a smooth and Lipschitz function. Then the Wasserstein distance has a good property that the perturbation $\mu_{a}$ of a measure $\mu$ can be represented by the push-forward measure of $\mu$ by a map $T_{a}:\mathbb{R}^{d}\ni x\mapsto x+a\xi\in\mathbb{R}^{d}$ for $a>0$ and $\xi\in C_{c}^{\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$, that is, $\mu_{a}={T_{a}}_{\#}\mu$, where ${T_{a}}_{\#}\mu$ is defined by ${T_{a}}_{\#}\mu(A)=\mu(T_{a}^{-1}(A))$ for all Borel set $A\subset\mathbb{R}^{d}$. On the other hand, when $0<\alpha<1$, the mobility function $u^{\alpha}$ is smooth but not Lipschitz because of the singularity of its derivative at $u=0$. Moreover, since it is not known that the weighted Wasserstein distance has a representation such as (1.2), we cannot use the same way to consider the perturbation.
Then we need to apply another way called the flow interchange lemma (see Section 4). The flow interchange lemma is introduced in [17] for the Wasserstein space, and Lisini, Matthes and Savaré apply it for the weighted Wasserstein space in [15]. But, since the Lipschitz property of the mobility function is needed for thier method, the flow interchange lemma does not work directly in the case of the mobility $u^{\alpha}$. To overcome this problem, we approximate the function $u^{\alpha}$ by a $C^{\infty}$ and Lipschitz function $m_{\varepsilon}:[0,\infty)\to[0,\infty)$ for $\varepsilon\in(0,1)$:
$m_{\varepsilon}(r)\coloneq(r+\varepsilon)^{\alpha},\ m_{\varepsilon}^{\prime}(r)=\frac{\alpha}{(r+\varepsilon)^{1-\alpha}}\leq\frac{\alpha}{\varepsilon^{1-\alpha}}\quad\mathrm{for}\ r\in[0,\infty),$
that is, we first consider the system:
$\begin{cases}\partial_{t}u=\nabla\cdot\left(\frac{p}{p-\alpha}m_{\varepsilon}(u)\nabla u^{p-\alpha}\right)-\nabla\cdot(\chi m_{\varepsilon}(u)\nabla v)&\mathrm{in}\ \Omega\times(0,\infty),\\ \partial_{t}v=\Delta v-v+u&\mathrm{in}\ \Omega\times(0,\infty),\\ \nabla u\cdot\boldsymbol{n}=\nabla v\cdot\boldsymbol{n}=0&\mathrm{on}\ \partial\Omega\times(0,\infty),\\ u(\cdot,0)=u_{0}(\cdot),\ v(\cdot,0)=v_{0}(\cdot)&\mathrm{in}\ \Omega.\end{cases}$ (1.1)ε
Thanks to this approximation, we can get the solutions to (1.1)ε, and then we need to obtain uniform estimates with respect to $\varepsilon$ and convergences as $\varepsilon\to 0$. The key point for uniform estimates is the boundedness of the functional $\boldsymbol{U}_{\varepsilon}:L^{p+1-\alpha}\cap\mathcal{P}(\Omega)\to\mathbb{R}$ defined by
$\boldsymbol{U}_{\varepsilon}(u)\coloneq\int_{\Omega}U_{\varepsilon}(u(x))\,dx,$
where $U_{\varepsilon}:[0,\infty)\to\mathbb{R}$ satisfies $U_{\varepsilon}^{\prime\prime}(r)m_{\varepsilon}(r)=1$ and $U_{\varepsilon}^{\prime}(0)=U_{\varepsilon}(0)=0$ (see Lemma 2.9). On the other hand, the key point for the convergences, in particular the pointwise convergence for $t$ weakly in $L^{1}\cap L^{p+1-\alpha}(\Omega)$, is the lower semicontinuity of the weighted Wasserstein distance (see Lemma 2.6 and Lemma 6.3). In order to get the convergence, we use the refined Ascoli–Arzelà theorem ([2, Proposition 3.3.1]), and in more detail, we need the estimate like the equi-continuity with respect to the weighted Wasserstein distance:
$W_{m}(u_{\tau}(t),u_{\tau}(s))\leq C(\sqrt{|t-s|}+\sqrt{\tau})\quad\mathrm{for}\ t,s\in[0,T],$
where $m(u)=u^{\alpha}$, $u_{\tau}$ is a pointwise constant function (Definition 5.1) and $C>0$ is a constant independent of $\varepsilon$ and $\tau$. However, we only obtain such estimate replaced $m$ by $m_{\varepsilon}$, that is, the distance depends on $\varepsilon$ (see Lemma 5.3). Thus combining the lower semicontinuity of the weighted Wasserstein distance with the above estimate, we prove the pointwise convergence for $t$ weakly in $L^{1}\cap L^{p+1-\alpha}(\Omega)$. This approach is original and may be applied for other cases if we use the lower semicontinuity of the weighted Wasserstein distance similarly.
Finally, we remark that the way of the approximation of $u^{\alpha}$ by $(u+\varepsilon)^{\alpha}$ is important in our case. In [15], they also approximate the mobility function by another way such as $u^{\alpha}$ by $(u+\varepsilon)^{\alpha}-\varepsilon^{\alpha}$. However, thier approximation requires that initial data $u_{0}$ belongs to $L^{2}(\Omega)$ in order to obtain the uniform estimate for $\boldsymbol{U}_{\varepsilon}$. If $p<1+\alpha$ then our initial data $u_{0}$ does not belong to $L^{2}(\Omega)$, thus we cannot use thier approximation in our case, especially in Theorem 1.2 and Theorem 1.4. On the other hand, in order to derive uniform estimate for $\boldsymbol{U}_{\varepsilon}$, our approximation requires that initial data $u_{0}$ belongs to $L^{2-\alpha}(\Omega)$, which is naturally satisfied, hence it is effective to use our apporoximation in our case.
This paper is organized as follows. In Section 2, we recall the definition of the weighted Wasserstein distance and some properties of it introduced in [8]. Section 3 is devoted to the time discrete variational scheme. In Section 4, we deal with the flow interchange lemma and prepare some lemmas to adapt it. Fundamentally, we refer to the method in [15], but our functions (minimizers in Section 3) have a lower regularity than thier ones, so we derive a suitable regurality of minimizers to obtain the Euler–Lagrange equation (Lemma 4.8). In Section 5, we consider uniform estimates with respect to $\tau$, which yield that the time discrete solution $(u_{\tau},v_{\tau})$ (Definition 5.1) converges to a weak solution to (1.1)ε. Then in Section 6, we also establish uniform estimates with respect to $\varepsilon$, which is guaranteed by the uniform estimate of $\boldsymbol{U}_{\varepsilon}$ (Lemma 2.9). In addition, the pointwise convergence for $t$ weakly in $L^{1}\cap L^{p+1-\alpha}(\Omega)$ (Lemma 6.3) plays an important role in this section. Using these estimates and convergences, we prove Theorem 1.1 ($p=1+\alpha$). Finally, in Section 7, we obtain additional uniform estimates to show Theorem 1.2 ($1+\alpha-2/d<p<1+\alpha$) and Theorem 1.4 ($p=1+\alpha-2/d$ and small $\chi>0$).
2 Preliminary
2.1 Notations
$\mathcal{P}(\mathbb{R}^{d})=\{\mu:\mu\ \mathrm{is\ a\ Borel\ probability\ measure\ in}\ \mathbb{R}^{d}\}$
$\mathcal{P}(\Omega)=\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\mu(\Omega)=1,\ \mu(\mathbb{R}^{d}\setminus\Omega)=0\}$
$\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})=\{\mu:\mu\ \mathrm{is\ a\ nonnegative\ Radon\ measure\ in}\ \mathbb{R}^{d}\}$
$\mathcal{M}^{+}(\Omega)=\{\mu\in\mathcal{M}_{loc}^{+}(\mathbb{R}^{d}):\mu(\mathbb{R}^{d}\setminus\Omega)=0\}$
$\mathcal{M}_{loc}(\mathbb{R}^{d};\mathbb{R}^{d})=\{\boldsymbol{\nu}:\boldsymbol{\nu}\ \mathrm{is\ a}\ \mathbb{R}^{d}\mbox{-}\mathrm{valued\ Radon\ measure\ in}\ \mathbb{R}^{d}\}$
$\mathcal{M}(\Omega;\mathbb{R}^{d})=\{\boldsymbol{\nu}\in\mathcal{M}_{loc}(\mathbb{R}^{d};\mathbb{R}^{d}):\boldsymbol{\nu}(\mathbb{R}^{d}\setminus\Omega)=0\}$
$C_{c}(\mathbb{R}^{d})=\{f\in C(\mathbb{R}^{d}):\mathrm{supp}(f)\ \mathrm{is\ compact\ in}\ \mathbb{R}^{d}\}$
$B_{R}=\{x\in\mathbb{R}^{d}:|x|<R\}$
Note that since $\Omega$ is bounded, then $\mathcal{P}(\Omega)\subset\mathcal{M}^{+}(\Omega)$. By the Riesz representation theorem, $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ (resp. $\mathcal{M}_{loc}(\mathbb{R}^{d};\mathbb{R}^{d})$) can be identified with the dual space of $C_{c}(\mathbb{R}^{d})$ (resp. $C_{c}(\mathbb{R}^{d};\mathbb{R}^{d})$).
2.2 Weighted Wasserstein distance
We recall the weighted Wasserstein distance which is a (pseudo) distance on the space of nonnegative Radon measures $\mathcal{M}^{+}(\Omega)$ and introduced in [8]. First, we recall the continuity equation for Radon measures.
Definition 2.1.
([8, Definition 4.2], Solutions of the continuity equation) Let $\mu_{0},\mu_{1}\in\mathcal{M}^{+}(\Omega)$. We denote by $CE(0,1;\mu_{0}\to\mu_{1})$ the set of pairs of time dependent measures $\{\mu_{t}\}_{t\in[0,1]}\subset\mathcal{M}^{+}(\Omega)$ and $\{\boldsymbol{\nu}_{t}\}_{t\in[0,1]}\subset\mathcal{M}(\Omega;\mathbb{R}^{d})$ such that
  1. 1.
    $t\mapsto\mu_{t}$ is weakly* continuous in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ : for all $f\in C_{c}(\mathbb{R}^{d})$,
    $\int_{\mathbb{R}^{d}}f(x)\,d\mu_{t}(x)\ \mbox{is continuous with respect to }t\in[0,1],$
  2. 2.
    $\{\boldsymbol{\nu}_{t}\}_{t\in[0,1]}$ is a Borel measureable family with
    $\int_{0}^{1}\int_{\Omega}\,d\boldsymbol{\nu}_{t}(x)\,dt<\infty,$
  3. 3.
    $(\mu_{t},\boldsymbol{\nu}_{t})$ satisfies the continuity equation in the weak sense : for all $\psi\in C_{c}^{1}(\mathbb{R}^{d}\times(0,1))$,
    $\int_{0}^{1}\int_{\mathbb{R}^{d}}\partial_{t}\psi(x,t)\,d\mu_{t}(x)\,dt+\int_{0}^{1}\int_{\mathbb{R}^{d}}\nabla\psi(x,t)\cdot d\boldsymbol{\nu}_{t}(x)\,dt=0,$
  4. 4.
    $\mu_{t}|_{t=0}=\mu_{0},\ \mu_{t}|_{t=1}=\mu_{1}$: for all $f\in C_{c}(\mathbb{R}^{d})$,
    $\lim_{t\to i}\int_{\mathbb{R}^{d}}f(x)\,d\mu_{t}(x)=\int_{\mathbb{R}^{d}}f(x)\,d\mu_{i}(x)\quad i=0,1.$
Next, we define the functional in the space of Radon measures, which is important for the definition of the weighted Wasserstein distance.
Definition 2.2 ([8, Section 3], The action functional).
Let $\mu\in\mathcal{M}^{+}(\Omega)$ and $\boldsymbol{\nu}\in\mathcal{M}(\Omega;\mathbb{R}^{d})$. Let $m:[0,\infty)\to[0,\infty)$ be a concave function. Then we define the action functional $\Psi_{\Omega}:\mathcal{M}^{+}(\Omega)\times\mathcal{M}(\Omega;\mathbb{R}^{d})\to[0,\infty]$ by
$\Psi_{\Omega}(\mu,\boldsymbol{\nu})\coloneq\begin{dcases}\int_{\Omega}\frac{|\boldsymbol{w}|^{2}}{m(\rho)}\,dx&\mathrm{if}\ \boldsymbol{\nu}^{\perp}=0,\\ \infty&\mathrm{if}\ \boldsymbol{\nu}^{\perp}\neq 0,\end{dcases}$
where $\mu=\rho\mathscr{L}^{d}+\mu^{\perp}$ and $\boldsymbol{\nu}=\boldsymbol{w}\mathscr{L}^{d}+\boldsymbol{\nu}^{\perp}$ are Lebesgue decomposition with respect to Lebesgue’s measure $\mathscr{L}^{d}$, that is, $\rho\in L^{1}(\Omega),\ \boldsymbol{w}\in L^{1}(\Omega;\mathbb{R}^{d})$ and $\mu^{\perp}$ (resp. $\boldsymbol{\nu}^{\perp}$) is a singular part of $\mu$ (resp. $\boldsymbol{\nu}$).
Definition 2.3 ([8, Definition 5.1], Weighted Wasserstein distance).
Let $\mu_{0},\mu_{1}\in\mathcal{M}^{+}(\Omega)$ and $m:[0,\infty)\to[0,\infty)$ be a concave function. Then weighted Wasserstein distance between $\mu_{0}$ and $\mu_{1}$ is defined as
$W_{m,\Omega}(\mu_{0},\mu_{1})^{2}$ $\coloneq\inf\left[\int_{0}^{1}\Psi_{\Omega}(\mu_{t},\boldsymbol{\nu}_{t})\,dt:(\mu_{t},\boldsymbol{\nu}_{t})\in CE(0,1;\mu_{0}\to\mu_{1})\right]$ (2.1)
$=\begin{dcases}\inf\left[\int_{0}^{1}\int_{\Omega}\frac{|\boldsymbol{w}_{t}(x)|^{2}}{m(\rho_{t}(x))}\,dx\,dt:(\mu_{t},\boldsymbol{\nu}_{t})\in CE(0,1;\mu_{0}\to\mu_{1})\right]\ &\mathrm{if}\ \boldsymbol{\nu}^{\perp}=0,\\ \infty\ &\mathrm{if}\ \boldsymbol{\nu}^{\perp}\neq 0.\end{dcases}$
We usually omit to write $\Omega$, then $W_{m}(\mu_{0},\mu_{1})$, but if we emphasize the domain $\Omega$, we write $W_{m,\Omega}(\mu_{0},\mu_{1})$.
Next, we collect some properties of the weighted Wasserstein distance ([8, Theorems 5.5, 5.4, 5.6, 2.3], [2, Lemma 8.1.10]).
Proposition 2.4 (Distance and topology).
The functional $W_{m}$ is a (pseudo) distance on $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ which induces a stronger topology than weak* one in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$. Moreover bounded sets with respect to $W_{m}$ are weakly* relatively compact in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$.
Lemma 2.5 (Existence of minimizers).
Let $\mu_{0},\mu_{1}\in\mathcal{M}^{+}(\Omega)$. Whenever the infimun in (2.1) is a finite value, it is attained by a curve $(\mu_{t},\boldsymbol{\nu}_{t})\in CE(0,1;\mu_{0}\to\mu_{1})$.
Lemma 2.6 (Lower semicontinuity).
Let $\{\Omega_{n}\}_{n}$ be a sequence of bounded domain converging to a bounded domain $\Omega$, that is, $\mathscr{L}^{d}|_{\Omega_{n}}\rightharpoonup\mathscr{L}^{d}|_{\Omega}$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ as $n\to\infty$:
$\int_{\mathbb{R}^{d}}f(x)\,d\mathscr{L}^{d}|_{\Omega_{n}}(x)\to\int_{\mathbb{R}^{d}}f(x)\,d\mathscr{L}^{d}|_{\Omega}(x)\quad\mathrm{as}\ n\to\infty,$
for all $f\in C_{c}(\mathbb{R}^{d})$. Moreover, let series of functions $\{m_{n}\}_{n}$ be monotonically decreasing with respect to $n$ and pointwise converging to $m$ as $n\to\infty$. Then, the map $(\mu_{0},\mu_{1})\mapsto W_{m}(\mu_{0},\mu_{1})$ is lower semicontinuous with respect to weak* convergence in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$: if sequences of measures $\{\mu_{0}^{n}\}_{n}$ and $\{\mu_{1}^{n}\}_{n}$ satisfy $\mu_{0}^{n}\rightharpoonup\mu_{0},\ \mu_{1}^{n}\rightharpoonup\mu_{1}$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ as $n\to\infty$ then
$W_{m,\Omega}(\mu_{0},\mu_{1})\leq\liminf_{n\to\infty}W_{m_{n},\Omega_{n}}(\mu_{0}^{n},\mu_{1}^{n}).$ (2.2)
Lemma 2.7 (Convolution).
Let $\psi:(0,\infty)\times\mathbb{R}^{d}\to[0,\infty)$ be a convex and lower semicontinuous function satisfying $\psi(\cdot,0)=0$. Let $\mu\in\mathcal{M}^{+}(\Omega)$ and $\boldsymbol{\nu}\in\mathcal{M}(\Omega;\mathbb{R}^{d})$ be such that $\mu=\rho\mathscr{L}^{d}+\mu^{\perp}$ and $\boldsymbol{\nu}=\boldsymbol{w}\mathscr{L}^{d}+\boldsymbol{\nu}^{\perp}$ where $\rho\in L^{1}(\Omega)$, $\boldsymbol{w}\in L^{1}(\Omega;\mathbb{R}^{d})$ and $\mu^{\perp}$, $\boldsymbol{\nu}^{\perp}$ are singular parts of Lebesgue decomposition. We define $\Psi(\mu,\boldsymbol{\nu})\coloneq\int_{\mathbb{R}^{d}}\psi(\rho(x),\boldsymbol{w}(x))\,dx=\int_{\Omega}\psi(\rho(x),\boldsymbol{w}(x))\,dx$ and let $k\in C_{c}^{\infty}(\mathbb{R}^{d})$ be a nonnegative convolution kernel with $\int_{\mathbb{R}^{d}}k(x)\,dx=1$. Then
$\Psi(\mu*k,\boldsymbol{\nu}*k)\leq\Psi(\mu,\boldsymbol{\nu}),$ (2.3)
where $\mu*k$ (resp. $\boldsymbol{\nu}*k$) is the measure defined by the (density) function
$x\mapsto\mu*k(x)\coloneq\int_{\mathbb{R}^{d}}k(x-y)\,d\mu(y)\ \left(\mbox{resp.}\ \boldsymbol{\nu}*k\coloneq\int_{\mathbb{R}^{d}}k(x-y)\,d\boldsymbol{\nu}(y)\right).$
Next lemma implies that the minimizer of the weighted Wasserstein distance can be approximated by smooth denisity functions. In [6, Lemma 3.6], they showed the existence of smooth functions ($\rho_{n}$ and $\boldsymbol{w}_{n}$ in next lemma) approximating the minimizer of the weighted Wasserstein distance, however we will prove that the weighted Wasserstein distance can be approximated by another smooth functions ($\rho_{n}$ and $\phi_{n}$ in next lemma). Note that, in [15, Proposition 2.2], they stated similar approximation lemma, but they did not give the proof, so we give rigorous one. Thanks to this approximation, we can do calculations simply and adapt the flow interchange lemma (see Section 4).
Lemma 2.8.
Let $m\in C^{\infty}(0,\infty)$ be a positive concave function such that $\inf_{r\in(0,\infty)}m(r)>0$. Let $\mu_{0},\mu_{1}\in L^{q}\cap\mathcal{P}(\Omega)$ for $q\in[1,\infty)$ with $W_{m}(\mu_{0},\mu_{1})<\infty$. Then for every decreasing sequence of smooth bounded sets $\Omega_{n}$ converging to $\Omega$ as $n\to\infty$, that is, $\Omega_{n+1}\subset\Omega_{n}$ for $n\in\mathbb{N}$ and $\mathscr{L}^{d}|_{\Omega_{n}}\rightharpoonup\mathscr{L}^{d}|_{\Omega}$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ as $n\to\infty$, there exists a vanishing sequence $\{b_{n}\}_{n}$ such that $\Omega_{[b_{n}]}\subset\Omega_{n}$, where $\Omega_{[b_{n}]}\coloneq\Omega+b_{n}B_{1}=\{x+b_{n}y:x\in\Omega,y\in B_{1}\}$, and there exist a nonnegative function $\rho_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1])$ and a function $\phi_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1])$ with $\nabla\phi_{n}\cdot\boldsymbol{n}=0$ on $\partial\Omega_{n}\times[0,1]$, satisfying the following:
  1. 1.
    $\|\rho_{n}(t)\|_{L^{1}(\Omega_{n})}=1$ for $n\in\mathbb{N},\ t\in[0,1]$,
  2. 2.
    $\|\rho_{n}(0)-\mu_{0}\|_{L^{q}(\mathbb{R}^{d})}+\|\rho_{n}(1)-\mu_{1}\|_{L^{q}(\mathbb{R}^{d})}\to 0$ as $n\to\infty$,
  3. 3.
    $(\rho_{n},m(\rho_{n})\nabla\phi_{n})$ satisfies the continuity equation:
    $\partial_{t}\rho_{n}(x,t)=-\nabla\cdot(m(\rho_{n}(x,t))\nabla\phi_{n}(x,t))\quad\mathrm{for\ all}\ (x,t)\in\Omega_{n}\times(0,1)$
    and moreover
    $W_{m}(\mu_{0},\mu_{1})^{2}=\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}m(\rho_{n}(x,t))|\nabla\phi_{n}(x,t)|^{2}\,dx\,dt.$
Proof.
Let $\{\Omega_{n}\}_{n}$ be a decreasing sequence of smooth convex bounded sets converging to $\Omega$ as $n\to\infty$. Since it is decreasing and bounded, we can find a sequence $\{b_{n}\}_{n}$ such that $b_{n}\to 0$ as $n\to\infty$ and $\Omega_{[b_{n}]}\subset\Omega_{n}$. Let $(\mu_{t},\boldsymbol{\nu}_{t})\in CE(0,1;\mu_{0}\to\mu_{1})$ be a minimizer of $W_{m}(\mu_{0},\mu_{1})$ (Lemma 2.5). Let us extend $(\mu_{t},\boldsymbol{\nu}_{t})$ outside the interval $[0,1]$ by setting $\boldsymbol{\nu}_{t}=0$ if $t<0$ or $t>1$, and $\mu_{t}=\mu_{0}$ if $t<0$, $\mu_{t}=\mu_{1}$ if $t>1$. Then $(\mu_{t},\boldsymbol{\nu}_{t})$ still satisfies the continuity equation. For $\{b_{n}\}_{n}$ as the above, let $k_{n}\in C_{c}^{\infty}(\mathbb{R}^{d})$ be a nonnegative mollifier such that $\mathrm{supp}(k_{n})\subset B_{b_{n}}$. Then define measures $\mu_{t}*k_{n}$ and $\boldsymbol{\nu}_{t}*k_{n}$ which have spatial smooth densities
$\tilde{\mu}_{n,t}(x)\coloneq\int_{\mathbb{R}^{d}}k_{n}(x-y)\,d\mu_{t}(y),$
$\tilde{\boldsymbol{\nu}}_{n,t}(x)\coloneq\int_{\mathbb{R}^{d}}k_{n}(x-y)\,d\boldsymbol{\nu}_{t}(y).$
Note that $\mathrm{supp}(\tilde{\mu}_{n,t}),\ \mathrm{supp}(\tilde{\boldsymbol{\nu}}_{n,t})\subset\Omega_{n}$. Moreover let $h_{\frac{1}{n}}\in C_{c}^{\infty}(\mathbb{R})$ be a nonnegative mollifier in $\mathbb{R}$ such that $\mathrm{supp}(h_{\frac{1}{n}})\subset[-\frac{1}{n},\frac{1}{n}]$ and define functions
$\tilde{\rho}_{n}(x,t)\coloneq\int_{\mathbb{R}}\tilde{\mu}_{n,z}(x)h_{\frac{1}{n}}(t-z)\,dz\quad\mathrm{for}\ (x,t)\in\mathbb{R}^{d}\times\mathbb{R},$
$\tilde{\boldsymbol{w}}_{n}(x,t)\coloneq\int_{\mathbb{R}}\tilde{\boldsymbol{\nu}}_{n,z}(x)h_{\frac{1}{n}}(t-z)\,dz\quad\mathrm{for}\ (x,t)\in\mathbb{R}^{d}\times\mathbb{R}.$
Then $\tilde{\rho}_{n}\in C^{\infty}(\mathbb{R}^{d}\times\mathbb{R})$ and $\tilde{\boldsymbol{w}}_{n}\in C^{\infty}(\mathbb{R}^{d}\times\mathbb{R};\mathbb{R}^{d})$, in addition, their spatial supports are included in $\Omega_{n}$. Notice that $\tilde{\rho}_{n}(\cdot,-\frac{1}{n})=\tilde{\mu}_{n,0}$ and $\tilde{\rho}_{n}(\cdot,1+\frac{1}{n})=\tilde{\mu}_{n,1}$. Indeed, for all $x\in\mathbb{R}^{d}$, we have
$\tilde{\rho}_{n}\left(x,-\frac{1}{n}\right)$ $=\int_{\mathbb{R}}\tilde{\mu}_{n,z}(x)h_{\frac{1}{n}}\left(-\frac{1}{n}-z\right)\,dz=\int_{-\frac{2}{n}}^{0}\tilde{\mu}_{n,z}(x)h_{\frac{1}{n}}\left(-\frac{1}{n}-z\right)\,dz$
$=\int_{-\frac{2}{n}}^{0}\tilde{\mu}_{n,0}(x)h_{\frac{1}{n}}\left(-\frac{1}{n}-z\right)\,dz=\tilde{\mu}_{n,0}(x).$
The other equality can be proved similarly. By the convexity of $|a|^{2}/m(b)$ for each $(a,b)\in\mathbb{R}^{d}\times(0,\infty)$ and Jensen’s inequality, we have
$\frac{|\tilde{\boldsymbol{w}}_{n}(x,t)|^{2}}{m(\tilde{\rho}_{n}(x,t))}\leq\int_{\mathbb{R}}\frac{|\tilde{\boldsymbol{\nu}}_{n,z}(x)|^{2}}{m(\tilde{\mu}_{n,z}(x))}h_{\frac{1}{n}}(t-z)\,dz\quad\mathrm{for}\ (x,t)\in\mathbb{R}^{d}\times\mathbb{R},$
then
$\int_{\mathbb{R}^{d}}\frac{|\tilde{\boldsymbol{w}}_{n}(x,t)|^{2}}{m(\tilde{\rho}_{n}(x,t))}\,dx\leq\int_{\mathbb{R}^{d}}\int_{\mathbb{R}}\frac{|\tilde{\boldsymbol{\nu}}_{n,z}(x)|^{2}}{m(\tilde{\mu}_{n,z}(x))}h_{\frac{1}{n}}(t-z)\,dz\,dx\quad\mathrm{for}\ t\in\mathbb{R}.$ (2.4)
Since $\tilde{\boldsymbol{w}}_{n}(\cdot,t)=0$ if t<1/nt<-1/n or $t>1+1/n$ and $\tilde{\boldsymbol{w}}_{n}(x,\cdot)=0$ if $x\in\mathbb{R}^{d}\setminus\Omega_{n}$, by (2.4) and Fubini’s theorem, it follows
$\int_{-\frac{1}{n}}^{1+\frac{1}{n}}\int_{\Omega_{n}}\frac{|\tilde{\boldsymbol{w}}_{n}(x,t)|^{2}}{m(\tilde{\rho}_{n}(x,t))}\,dx\,dt$ $=\int_{\mathbb{R}}\int_{\mathbb{R}^{d}}\frac{|\tilde{\boldsymbol{w}}_{n}(x,t)|^{2}}{m(\tilde{\rho}_{n}(x,t))}\,dx\,dt$
$\leq\int_{\mathbb{R}}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}}\frac{|\tilde{\boldsymbol{\nu}}_{n,z}(x)|^{2}}{m(\tilde{\mu}_{n,z}(x))}h_{\frac{1}{n}}(t-z)\,dz\,dx\,dt$
$=\int_{\mathbb{R}}\int_{\mathbb{R}^{d}}\frac{|\tilde{\boldsymbol{\nu}}_{n,z}(x)|^{2}}{m(\tilde{\mu}_{n,z}(x))}\,dx\,dz$
$\leq W_{m}(\mu_{0},\mu_{1})^{2},$ (2.5)
where last inequality is followed by Lemma 2.7. Thus we set
$\rho_{n}(x,t)\coloneq\tilde{\rho}_{n}\left(x,c_{n}t-\frac{1}{n}\right),\ \boldsymbol{w}_{n}(x,t)\coloneq c_{n}\tilde{\boldsymbol{w}}_{n}\left(x,c_{n}t-\frac{1}{n}\right)\quad\mbox{where}\ c_{n}\coloneq 1+\frac{2}{n}.$
Then $\rho_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1]),\ \boldsymbol{w}_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1];\mathbb{R}^{d})$ and it holds
$\partial_{t}\rho_{n}(x,t)+\nabla\cdot\boldsymbol{w}_{n}(x,t)=0\quad(x,t)\in\Omega_{n}\times(0,1).$
By (2.2) and the change of variables, we have
$\frac{1}{c_{n}}\int_{0}^{1}\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}(x,t)|^{2}}{m(\rho_{n}(x,t))}\,dx\,dt$ $=\frac{1}{c_{n}}\int_{0}^{1}\int_{\Omega_{n}}\frac{c_{n}^{2}|\tilde{\boldsymbol{w}}_{n}\left(x,c_{n}t-\frac{1}{n}\right)|^{2}}{m\left(\tilde{\rho}_{n}\left(x,c_{n}t-\frac{1}{n}\right)\right)}\,dx\,dt$
$=\int_{-\frac{1}{n}}^{1+\frac{1}{n}}\int_{\Omega_{n}}\frac{|\tilde{\boldsymbol{w}}_{n}(x,t)|^{2}}{m(\tilde{\rho}_{n}(x,t))}\,dx\,dt$
$\leq W_{m}(\mu_{0},\mu_{1})^{2}.$ (2.6)
Since $\mu_{0},\mu_{1}\in L^{q}\cap\mathcal{P}(\Omega)$, using the property of the mollifier, we have
$\rho_{n}(i)\to\mu_{i}\ \mathrm{in}\ L^{q}(\mathbb{R}^{d})\quad\mathrm{as}\ n\to\infty\quad i=0,1,$
in particular
$\rho_{n}(i)\rightharpoonup\mu_{i}\ \mathrm{weakly^{*}\ in}\ \mathcal{M}_{loc}^{+}(\mathbb{R}^{d})\quad\mathrm{as}\ n\to\infty\quad i=0,1.$
We infer from Lemma 2.6 that
$W_{m}(\mu_{0},\mu_{1})^{2}\leq\liminf_{n\to\infty}W_{m}(\rho_{n}(0),\rho_{n}(1))^{2}\leq\liminf_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}|^{2}}{m(\rho_{n})}\,dx\,dt.$
Combining the above with (2.2) and letting $n\to\infty$, we have
$W_{m}(\mu_{0},\mu_{1})^{2}=\lim_{n\to\infty}W_{m}(\rho_{n}(0),\rho_{n}(1))^{2}=\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}|^{2}}{m(\rho_{n})}\,dx\,dt.$ (2.7)
Next, for fixed $t\in[0,1]$ we consider the following equation:
$\begin{cases}\nabla\cdot\boldsymbol{w}_{n}(x,t)=\nabla\cdot(m(\rho_{n}(x,t))\nabla\phi_{n}(x))&x\in\Omega_{n},\\ \nabla\phi_{n}(x)\cdot\boldsymbol{n}=0&x\in\partial\Omega_{n}.\end{cases}$ (2.8)
Since $\mathrm{supp}(\boldsymbol{w}_{n}(\cdot,t))\subset\Omega_{n}$, then $\boldsymbol{w}_{n}(x,t)=0$ for $x\in\partial\Omega_{n}$. By the Gauss–Green theorem, it follows
$\int_{\Omega_{n}}\nabla\cdot\boldsymbol{w}_{n}(x,t)\,dx=0.$
Hence (2.8) has a unique weak solution $\phi_{n}\in H^{1}(\Omega_{n})$ such that
$\int_{\Omega_{n}}m(\rho_{n})\nabla\phi_{n}\cdot\nabla\psi\,dx=\int_{\Omega_{n}}\boldsymbol{w}_{n}\cdot\nabla\psi\,dx\quad\forall\psi\in H^{1}(\Omega_{n}).$ (2.9)
Due to the elliptic regurality theorem, $\phi_{n}$ can be a smooth function and satisfies $\nabla\cdot\boldsymbol{w}_{n}=\nabla\cdot(m(\rho_{n})\nabla\phi_{n})$ in $\Omega_{n}$. Taking $\psi=\phi_{n}$ in (2.9) and using Hölder’s inequality, we have
$\int_{\Omega_{n}}m(\rho_{n})|\nabla\phi_{n}|^{2}\,dx=\int_{\Omega_{n}}\boldsymbol{w}_{n}\cdot\nabla\phi_{n}\,dx\leq\left(\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}|^{2}}{m(\rho_{n})}\,dx\right)^{\frac{1}{2}}\left(\int_{\Omega_{n}}m(\rho_{n})|\nabla\phi_{n}|^{2}\,dx\right)^{\frac{1}{2}},$
then
$\int_{\Omega_{n}}m(\rho_{n})|\nabla\phi_{n}|^{2}\,dx\leq\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}|^{2}}{m(\rho_{n})}\,dx.$
In addition, since it holds
$\partial_{t}\rho_{n}+\nabla\cdot(m(\rho_{n})\nabla\phi_{n})=\partial_{t}\rho_{n}+\nabla\cdot\boldsymbol{w}_{n}=0\quad\mathrm{in}\ \Omega_{n}\times(0,1),$
we have $(\rho_{n},m(\rho_{n})\nabla\phi_{n})\in CE(0,1;\rho_{n}(0),\rho_{n}(1))$. Combining this with (2.7), we can conclude
$W_{m}(\mu_{0},\mu_{1})^{2}$ $=\lim_{n\to\infty}W_{m}(\rho_{n}(0),\rho_{n}(1))^{2}\leq\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}m(\rho_{n})|\nabla\phi_{n}|^{2}\,dx\,dt$
$\leq\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}\frac{|\boldsymbol{w}_{n}|^{2}}{m(\rho_{n})}\,dx\,dt=W_{m}(\mu_{0},\mu_{1})^{2},$
then
$W_{m}(\mu_{0},\mu_{1})^{2}=\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}m(\rho_{n})|\nabla\phi_{n}|^{2}\,dx\,dt.$
The proof is completed. ∎
2.3 Properties of the functional $\boldsymbol{U}_{\varepsilon}$ and a boundary estimate
First, we establish the uniform estimate for the functional $\boldsymbol{U}_{\varepsilon}$ with respect to $\varepsilon$. This estimate plays an important role in Sections 5 and 6.
Lemma 2.9.
Let $p\geq 1,0<\alpha<1$ and $U_{\varepsilon}:[0,\infty)\to\mathbb{R}$ be a function such that $U_{\varepsilon}^{\prime\prime}(r)m_{\varepsilon}(r)=1$ for $r\in[0,\infty)$ and $U_{\varepsilon}^{\prime}(0)=U_{\varepsilon}(0)=0$, where $m_{\varepsilon}(r)=(r+\varepsilon)^{\alpha}$. Then setting $\boldsymbol{U}_{\varepsilon}:L^{p+1-\alpha}\cap\mathcal{P}(\Omega)\to\mathbb{R}$ by
$\boldsymbol{U}_{\varepsilon}(u)\coloneq\int_{\Omega}U_{\varepsilon}(u(x))\,dx,$
it hold $\boldsymbol{U}_{\varepsilon}\geq 0$ and
$\boldsymbol{U}_{\varepsilon}(u)\leq\frac{1}{1-\alpha}\|u\|_{L^{2-\alpha}(\Omega)}^{2-\alpha}\quad\mathrm{for}\ u\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega).$ (2.10)
In addition, $\boldsymbol{U}_{\varepsilon}$ is lower semicontinuous with respect to weak* convergence in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$.
Proof.
First, the function $U_{\varepsilon}$ can be explicitly represented as
$U_{\varepsilon}(r)=\frac{1}{(2-\alpha)(1-\alpha)}[(r+\varepsilon)^{2-\alpha}-\varepsilon^{2-\alpha}]-\frac{\varepsilon^{1-\alpha}}{1-\alpha}r.$
Since $U_{\varepsilon}^{\prime\prime}\geq 0$ and $U_{\varepsilon}^{\prime}(0)=U_{\varepsilon}(0)=0$, we see $U_{\varepsilon}\geq 0$ and then $\boldsymbol{U}_{\varepsilon}\geq 0$. Using the mean value theorem and the inequality
$a^{\beta}-b^{\beta}\leq|a-b|^{\beta}\quad\mathrm{for}\ a,b\geq 0,\ 0<\beta<1,$ (2.11)
we have for $r\in[0,\infty)$,
$U_{\varepsilon}(r)$ $=\frac{1}{(2-\alpha)(1-\alpha)}[(r+\varepsilon)^{2-\alpha}-\varepsilon^{2-\alpha}]-\frac{\varepsilon^{1-\alpha}}{1-\alpha}r$
$\leq\frac{1}{1-\alpha}r(r+\varepsilon)^{1-\alpha}-\frac{1}{1-\alpha}r\varepsilon^{1-\alpha}$
$=\frac{r}{1-\alpha}[(r+\varepsilon)^{1-\alpha}-\varepsilon^{1-\alpha}]\leq\frac{1}{1-\alpha}r^{2-\alpha},$
then
$\boldsymbol{U}_{\varepsilon}(u)\leq\frac{1}{1-\alpha}\|u\|_{L^{2-\alpha}(\Omega)}^{2-\alpha}\quad\mathrm{for}\ u\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega).$
Finally, since the function $U_{\varepsilon}$ is convex and continuous, by [1, Theorem 2.3.4], $\boldsymbol{U}_{\varepsilon}$ is lower semicontinuous with respect to weak* convergence in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$. ∎
Next lemma is about the estimate on the smooth boundary of convex domain ([9, Lemma 5.1]), and using this lemma later, we need to assume that domain $\Omega$ is convex.
Lemma 2.10.
Let $\Omega$ be a smooth convex set in $\mathbb{R}^{d}$ and $\varphi\in C^{3}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$. Then
$\nabla^{2}\varphi\nabla\varphi\cdot\boldsymbol{n}=\sum_{i,j=1}^{d}\partial_{ij}^{2}\varphi\,\partial_{i}\varphi\,n_{j}\leq 0\quad\mathrm{on}\ \partial\Omega,$
where $\nabla^{2}\varphi$ is the Hessian matrix and $\boldsymbol{n}=(n_{1},\cdots,n_{d})$ is the outer unit normal vector to $\partial\Omega$.
3 Existence of minimizers
In this section, we consider the following discrete scheme:
let $X\coloneq(L^{p+1-\alpha}\cap\mathcal{P}(\Omega))\times H^{1}(\Omega)$ and $m_{\varepsilon}(r)=(r+\varepsilon)^{\alpha}$ for $r\geq 0$. For a fixed time step $\tau>0$,
$\mbox{find}\ (u_{\tau}^{k},v_{\tau}^{k})\in X\ \mbox{satisfying}\ F_{\tau}(u_{\tau}^{k},v_{\tau}^{k})=\inf_{(u,v)\in X}F_{\tau}(u,v)\ \mbox{for each}\ k\in\mathbb{N},\$(3.1)
where $(u_{\tau}^{0},v_{\tau}^{0})=(u_{0},v_{0})$,
$F_{\tau}(u,v)\coloneq\frac{1}{2\tau}\left(\frac{W_{m_{\varepsilon}}(u,u_{\tau}^{k-1})^{2}}{\chi}+\|v-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\right)+E(u,v)\quad(u,v)\in X,$ (3.2)
and $W_{m_{\varepsilon}}$ is the weighted Wasserstein distance on the space of nonnegative Radon measures
$\mathcal{M}^{+}(\Omega)$ (see Definition 2.3). Notice that $\mathcal{P}(\Omega)\subset\mathcal{M}^{+}(\Omega)$. In order to show the existence of minimizers, we apply direct method, then we check that the functional $F_{\tau}$ is bounded below in $X$, the sublevel set of $F_{\tau}$ is relatively compact in $X$ and $F_{\tau}$ is lower semicontinuous with respect to weak topology in $X$.
3.1 Boundedness from below of the functional $F_{\tau}$
Lemma 3.1.
Let $p\geq 1+\alpha-2/d$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then the energy functional $E$ is bounded below in $X$. In particular the functional $F_{\tau}$ is also bounded below.
Proof.
If $p\geq 1+\alpha$ then $p+1-\alpha\geq 2$. Thus we infer from Hölder’s inequality and the interpolation inequality that
$\|uv\|_{L^{1}(\Omega)}$ $\leq\|u\|_{L^{2}(\Omega)}\|v\|_{L^{2}(\Omega)}$
$\leq\|u\|_{L^{p+1-\alpha}(\Omega)}^{\frac{p+1-\alpha}{2(p-\alpha)}}\|u\|_{L^{1}(\Omega)}^{\frac{p-1-\alpha}{2(p-\alpha)}}\|v\|_{H^{1}(\Omega)}$
$=\|u\|_{L^{p+1-\alpha}(\Omega)}^{\frac{p+1-\alpha}{2(p-\alpha)}}\|v\|_{H^{1}(\Omega)}.$
Note that $(p+1-\alpha)/(p-\alpha)\leq p+1-\alpha$ because of $p\geq 1+\alpha$.
On the other hand, when $1+\alpha-2/d<p<1+\alpha$, since $d\geq 2$, it follows
$p>1+\alpha-\frac{2}{d}\geq 1+\alpha-\frac{4}{d+2},$
then
$p+1-\alpha>\frac{2d}{d+2}.$
When $d=2$, we can choose $q\in(1,p+1-\alpha)$ satisfying $q<2/(2-p+\alpha)$ and fix it. By Hölder’s inequality, the interpolation inequality and the Sobolev embedding, we have for $(u,v)\in X$,
$\|uv\|_{L^{1}(\Omega)}$ $\leq\|u\|_{L^{q}(\Omega)}\|v\|_{L^{q^{*}}(\Omega)}$
$\leq\|u\|_{L^{p+1-\alpha}(\Omega)}^{\theta_{2}}\|u\|_{L^{1}(\Omega)}^{1-\theta_{2}}\|v\|_{L^{q^{*}}(\Omega)}$
$\leq C\|u\|_{L^{p+1-\alpha}(\Omega)}^{\theta_{2}}\|v\|_{H^{1}(\Omega)},$
where $C$ is a constant and
$q^{*}=\frac{q}{q-1},\quad\theta_{2}\coloneq\frac{(p+1-\alpha)(q-1)}{(p-\alpha)q}\in(0,1).$
Note that $2\theta_{2}<p+1-\alpha$ since $q<2/(2-p+\alpha)$. On the other hand, when $d\geq 3$, we infer from the similar estimate that
$\|uv\|_{L^{1}(\Omega)}$ $\leq\|u\|_{L^{\frac{2d}{d+2}}(\Omega)}\|v\|_{L^{\frac{2d}{d-2}}(\Omega)}$
$\leq\|u\|_{L^{p+1-\alpha}(\Omega)}^{\theta_{d}}\|u\|_{L^{1}(\Omega)}^{1-\theta_{d}}\|v\|_{L^{\frac{2d}{d-2}}(\Omega)}$
$\leq C\|u\|_{L^{p+1-\alpha}(\Omega)}^{\theta_{d}}\|v\|_{H^{1}(\Omega)},$
where
$\theta_{d}\coloneq\frac{(p+1-\alpha)(d-2)}{(p-\alpha)2d}\in(0,1).$
Observe that $2\theta_{d}<p+1-\alpha$ because of $p>1+\alpha-2/d$. Hence, when $p>1+\alpha-2/d$ and $d\geq 2$, by Young’s inequality, it follows
$\|uv\|_{L^{1}(\Omega)}$ $\leq\frac{1}{\chi(p+1-\alpha)}\|u\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}+C(\alpha,p,d,\chi),$ (3.3)
where $C(\alpha,p,d,\chi)$ is a constant. Then for $(u,v)\in X$, we obtain
$E(u,v)$ $\geq\frac{p}{\chi(p-\alpha)(p+1-\alpha)}\|u\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}-\frac{1}{\chi(p+1-\alpha)}\|u\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}$
$\quad-\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}+\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+\frac{1}{2}\|v\|_{L^{2}(\Omega)}^{2}-C(\alpha,p,d,\chi)$
$\geq\frac{\alpha}{\chi(p-\alpha)(p+1-\alpha)}\|u\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}-C(\alpha,p,d,\chi)$ (3.4)
$\geq-C(\alpha,p,d,\chi)>-\infty.$
If $p=1+\alpha-2/d$ then $p+1-\alpha=2-2/d$ and $\theta_{d}=1-1/d$. Hence by the same argument in the above, it follows
$\|uv\|_{L^{1}(\Omega)}$ $\leq C\|u\|_{L^{2-\frac{2}{d}}(\Omega)}^{1-\frac{1}{d}}\|v\|_{H^{1}(\Omega)}$
$\leq C\|u\|_{L^{2-\frac{2}{d}}(\Omega)}^{2-\frac{2}{d}}+\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}.$ (3.5)
Then for $(u,v)\in X$, we obtain
$E(u,v)$ $\geq\frac{1+\alpha-\frac{2}{d}}{\chi\left(1-\frac{2}{d}\right)\left(2-\frac{2}{d}\right)}\|u\|_{L^{2-\frac{2}{d}}(\Omega)}^{2-\frac{2}{d}}-C\|u\|_{L^{2-\frac{2}{d}}(\Omega)}^{2-\frac{2}{d}}$
$\quad-\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}+\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+\frac{1}{2}\|v\|_{L^{2}(\Omega)}^{2}$
$\geq\frac{1}{\chi}\left[\frac{1+\alpha-\frac{2}{d}}{\left(1-\frac{2}{d}\right)\left(2-\frac{2}{d}\right)}-\chi C\right]\|u\|_{L^{2-\frac{2}{d}}(\Omega)}^{2-\frac{2}{d}}+\frac{1}{4}\|v\|_{H^{1}(\Omega)}^{2}.$ (3.6)
If $\chi>0$ is small enough then the first term of the right hand side is positive and $E$ is bounded below in $X$. Since $\chi>0$, $W_{m_{\varepsilon}}\geq 0$ and $\|\cdot\|_{L^{2}(\Omega)}\geq 0$, $F_{\tau}$ is also bounded below. We complete the proof. ∎
3.2 Compactness
Lemma 3.2.
Let $\{(u_{n},v_{n})\}_{n\in\mathbb{N}}$ be a minimizing sequence of $F_{\tau}$ in $X$. Then there exist a subsequence $\{(u_{n_{l}},v_{n_{l}})\}_{l\in\mathbb{N}}$ and a pair of functions $(u,v)\in X$ such that
$u_{n_{l}}\rightharpoonup u\quad\mathrm{weakly\ in}\ L^{1}\cap L^{p+1-\alpha}(\Omega)\ \mathrm{as}\ l\to\infty,$
$v_{n_{l}}\rightharpoonup v\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ l\to\infty,$
$v_{n_{l}}\to v\quad\mathrm{strongly\ in}\ L^{q}(\Omega)\ \mathrm{as}\ l\to\infty\ \mathrm{for\ all}\ q\in[2,2^{*}),$
where
$2^{*}\coloneq\begin{cases}\infty\quad&\mathrm{if}\ d=2,\\ \frac{2d}{d-2}&\mathrm{if}\ d\geq 3.\end{cases}$
Proof.
Let $\{(u_{n},v_{n})\}\subset X$ be a minimizing sequence of $F_{\tau}$, that is, $F_{\tau}(u_{n},v_{n})$ is bounded in $\mathbb{R}$. Combining this with the estimate (3.1) or (3.1), we get the boundedness of $\|u_{n}\|_{L^{p+1-\alpha}(\Omega)}$ and $\|v_{n}\|_{H^{1}(\Omega)}$: there exists a constant $C=C(\alpha,p,d,\chi)$ such that
$\|u_{n}\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}\leq C,$ (3.7)
$\|v_{n}\|_{H^{1}(\Omega)}^{2}\leq C.$
Since $p+1-\alpha>1$, by the Banach–Alaoglu theorem, there exist subsequences $\{u_{n_{l}}\}_{l\in\mathbb{N}},\{v_{n_{l}}\}_{l\in\mathbb{N}}$ and functions $u\in L^{p+1-\alpha}(\Omega),v\in H^{1}(\Omega)$ such that
$u_{n_{l}}\rightharpoonup u\quad\mathrm{weakly\ in}\ L^{p+1-\alpha}(\Omega)\ \mathrm{as}\ l\to\infty,$ (3.8)
$v_{n_{l}}\rightharpoonup v\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ l\to\infty.$
Moreover, by the Rellich–Kondrachov theorem, we can take a subsequence, still denote $\{v_{n_{l}}\}$, satisfying
$v_{n_{l}}\to v\quad\mathrm{strongly\ in}\ L^{q}(\Omega)\ \mathrm{as}\ l\to\infty\ \mathrm{for\ all}\ q\in[2,2^{*}).$
Since $\{u_{n_{l}}\}\subset\mathcal{P}(\Omega)$, $\|u_{n_{l}}\|_{L^{1}(\Omega)}$ is bounded. For $c>0$, we see
$\int_{\{u_{n_{l}}\geq c\}}u_{n_{l}}\,dx\leq\int_{\{u_{n_{l}}\geq c\}}\frac{u_{n_{l}}^{p+1-\alpha}}{c^{p-\alpha}}\,dx\leq\frac{1}{c^{p-\alpha}}\int_{\Omega}u_{n_{l}}^{p+1-\alpha}\,dx.$
From (3.7), we have
$\limsup_{c\to\infty}\sup_{l\in\mathbb{N}}\int_{\{u_{n_{l}}\geq c\}}u_{n_{l}}\,dx=0.$
Thus $\{u_{n_{l}}\}$ is equi-integrable. By the Dunford–Pettis theorem, there exist a subsequence (not relabeled) and a function $\tilde{u}\in L^{1}(\Omega)$ such that
$u_{n_{l}}\rightharpoonup\tilde{u}\quad\mathrm{weakly\ in}\ L^{1}(\Omega)\ \mathrm{as}\ l\to\infty.$ (3.9)
Here, for all $f\in C_{c}^{\infty}(\Omega)$, from (3.8) and (3.9), we have
$\int_{\Omega}u(x)f(x)\,dx=\int_{\Omega}\tilde{u}(x)f(x)\,dx.$
Therefore we see $u=\tilde{u}$ a.e. in $\Omega$. By (3.9), we have
$1=\int_{\Omega}u_{n_{l}}(x)\,dx\to\int_{\Omega}u(x)\,dx\quad\mathrm{as}\ l\to\infty.$
Hence we obtain $u\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega)$ then $(u,v)\in X$. The proof is completed. ∎
3.3 Lower semicontinuity
Lemma 3.3.
Let $\{(u_{n},v_{n})\}\subset X$ and $(u,v)\in X$ satisfying $u_{n}\rightharpoonup u$ weakly in $L^{1}\cap L^{p+1-\alpha}(\Omega)$ and $v_{n}\to v$ weakly in $H^{1}(\Omega)$ and strongly in $L^{q}(\Omega)$ as $n\to\infty$ for all $q\in[2,2^{*})$. Then
$F_{\tau}(u,v)\leq\liminf_{n\to\infty}F_{\tau}(u_{n},v_{n}).$
Proof.
Fix $(u_{\tau}^{k-1},v_{\tau}^{k-1})\in X$. Let $\{(u_{n},v_{n})\}\subset X$ and $(u,v)\in X$ satisfying $u_{n}\rightharpoonup u$ weakly in $L^{1}\cap L^{p+1-\alpha}(\Omega)$ and $v_{n}\to v$ weakly in $H^{1}(\Omega)$ and strongly in $L^{q}(\Omega)$ for all $q\in[2,2^{*})$ as $n\to\infty$. Then, for all $f\in C_{c}(\mathbb{R}^{d})$, it follows
$\int_{\mathbb{R}^{d}}f(x)u_{n}(x)\,dx=\int_{\Omega}f(x)u_{n}(x)\,dx\to\int_{\Omega}f(x)u(x)\,dx=\int_{\mathbb{R}^{d}}f(x)u(x)\,dx\quad\mathrm{as}\ n\to\infty,$
thus $u_{n}\rightharpoonup u$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ as $n\to\infty$. Since $W_{m_{\varepsilon}}$ is lower semicontinuous with respect to weak* convergence in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ (see Lemma 2.6), we have
$W_{m_{\varepsilon}}(u,u_{\tau}^{k-1})^{2}\leq\liminf_{n\to\infty}W_{m_{\varepsilon}}(u_{n},u_{\tau}^{k-1})^{2}.$
In addition, since the norm is lower semicontinuous with respect to the weak topology, we have
$\|v-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\leq\liminf_{n\to\infty}\|v_{n}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2},$
$\|u\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}\leq\liminf_{n\to\infty}\|u_{n}\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha},$
$\|v\|_{H^{1}(\Omega)}^{2}\leq\liminf_{n\to\infty}\|v_{n}\|_{H^{1}(\Omega)}^{2}.$
We will show
$\lim_{n\to\infty}\|u_{n}v_{n}\|_{L^{1}(\Omega)}=\|uv\|_{L^{1}(\Omega)}.$ (3.10)
By Hölder’s inequality, it follows
$\left|\int_{\Omega}(uv-u_{n}v_{n})\,dx\right|$ $\leq\left|\int_{\Omega}(u-u_{n})v\,dx\right|+\int_{\Omega}|u_{n}||v-v_{n}|\,dx$
$\leq\left|\int_{\Omega}(u-u_{n})v\,dx\right|+\|u_{n}\|_{L^{p+1-\alpha}(\Omega)}\|v-v_{n}\|_{L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)}.$
Note that $2\leq(p+1-\alpha)/(p-\alpha)<2^{*}$ and $v\in L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)$ because of $p+1-\alpha>2/d,p\geq 1$ and $0<\alpha<1$. Since $\|u_{n}\|_{L^{p+1-\alpha}(\Omega)}$ is bounded, using asummptions, we obtain
$\lim_{n\to\infty}\left|\int_{\Omega}(uv-u_{n}v_{n})\,dx\right|=0,$
then (3.10) holds. From these, we complete the proof. ∎
3.4 Conclusion
Proposition 3.4.
Let $p\geq 1+\alpha-2/d$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Let $(u_{0},v_{0})\in X$ be a pair of nonnegative functions. Then for each $k\in\mathbb{N}$, there is at least one minimizer $(u_{\tau}^{k},v_{\tau}^{k})\in X$ in (3.1) and the following inequalities hold:
$\frac{1}{2\tau}\left(\frac{W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{\chi}+\|v_{\tau}^{k}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\right)+E(u_{\tau}^{k},v_{\tau}^{k})$ (3.11)
$\hskip 14.22636pt\leq\frac{1}{2\tau}\left(\frac{W_{m_{\varepsilon}}(\tilde{u},u_{\tau}^{k-1})^{2}}{\chi}+\|\tilde{v}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\right)+E(\tilde{u},\tilde{v})\quad\forall(\tilde{u},\tilde{v})\in X,\ \forall k\in\mathbb{N},$
$E(u_{\tau}^{k},v_{\tau}^{k})\leq E(u_{\tau}^{k-1},v_{\tau}^{k-1})\quad\forall k\in\mathbb{N}.$ (3.12)
Proof.
Let $(u_{\tau}^{k-1},v_{\tau}^{k-1})\in X$ for $k\in\mathbb{N}$. By Lemma 3.1, there exists a minimizing sequence $\{(u_{n},v_{n})\}\subset X$ such that $F_{\tau}(u_{n},v_{n})\to\inf_{(u,v)\in X}F_{\tau}(u,v)>$-$\infty$ as $n\to\infty$. By Lemma 3.2, there exist a subsequence $\{(u_{n_{l}},v_{n_{l}})\}_{l\in\mathbb{N}}$ and a pair of functions $(u_{\tau}^{k},v_{\tau}^{k})\in X$ such that
$u_{n_{l}}\rightharpoonup u_{\tau}^{k}\quad\mathrm{weakly\ in}\ L^{1}\cap L^{p+1-\alpha}(\Omega)\ \mathrm{as}\ l\to\infty,$
$v_{n_{l}}\rightharpoonup v_{\tau}^{k}\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ l\to\infty,$
$v_{n_{l}}\to v_{\tau}^{k}\quad\mathrm{strongly\ in}\ L^{q}(\Omega)\ \mathrm{as}\ l\to\infty\ \mathrm{for\ all}\ q\in[2,2^{*}).$
By Lemma 3.3, we have
$F_{\tau}(u_{\tau}^{k},v_{\tau}^{k})\leq\liminf_{l\to\infty}F_{\tau}(u_{n_{l}},v_{n_{l}}).$
As a result, we see
$\inf_{(u,v)\in X}F_{\tau}(u,v)\leq F_{\tau}(u_{\tau}^{k},v_{\tau}^{k})\leq\liminf_{l\to\infty}F_{\tau}(u_{n_{l}},v_{n_{l}})=\inf_{(u,v)\in X}F_{\tau}(u,v).$
Thus $(u_{\tau}^{k},v_{\tau}^{k})$ is the minimizer of $F_{\tau}$ in $X$ and (3.11) holds obviously. In particular, choosing $(\tilde{u},\tilde{v})=(u_{\tau}^{k-1},v_{\tau}^{k-1})$ in (3.11), we obtain (3.12) and the proof is completed. ∎
Remark 3.5.
In Proposition 3.4, we may take $v_{\tau}^{k}\in H^{1}(\Omega)$ which is not nonnegative, however, we can choose a pair of nonnegative functions as a minimizer of $F_{\tau}$ in $X$. Indeed, let $(u_{\tau}^{k},v_{\tau}^{k})$ be the minimizer in Proposition 3.4 and assume that $v_{\tau}^{k-1}$ is nonnegative. Then since $u_{\tau}^{k}$ is nonnegative, $|v_{\tau}^{k}|\in H^{1}(\Omega)$ and $||v_{\tau}^{k}|-v_{\tau}^{k-1}|\leq|v_{\tau}^{k}-v_{\tau}^{k-1}|$, we have
$F_{\tau}(u_{\tau}^{k},v_{\tau}^{k})$ $\leq F_{\tau}(u_{\tau}^{k},|v_{\tau}^{k}|)$
$=\frac{1}{2\tau}\left[\frac{W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{\chi}+\||v_{\tau}^{k}|-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\right]+E(u_{\tau}^{k},|v_{\tau}^{k}|)$
$\leq\frac{1}{2\tau}\left[\frac{W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{\chi}+\|v_{\tau}^{k}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\right]+E(u_{\tau}^{k},v_{\tau}^{k})=F_{\tau}(u_{\tau}^{k},v_{\tau}^{k}).$
Thus we can choose $(u_{\tau}^{k},|v_{\tau}^{k}|)\in X$ as a minimizer of $F_{\tau}$ in $X$ instead of $(u_{\tau}^{k},v_{\tau}^{k})$. In the rest of the paper, we call $(u_{\tau}^{k},|v_{\tau}^{k}|)$ a minimizer of $F_{\tau}$ in $X$ for $k\in\mathbb{N}$ and denote by $(u_{\tau}^{k},v_{\tau}^{k})$.
4 Euler–Lagrange equations
4.1 Flow interchange lemma
First, we show the existence of a solution to the other equation which is used later. The important properties of this solution are the mass conservation law and nonnegativity.
Proposition 4.1.
Let $1+\alpha-2/d\leq p\leq 1+\alpha$. Let $w_{0}\in L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega)$ be a nonnegative function. Then, for $\delta>0$ and $\varphi\in C^{\infty}(\bar{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$, there exist $T_{0}=T_{0}(\alpha,\varepsilon,\varphi,\Omega,w_{0})>0$ and a unique local solution $w$ satisfying
$\bullet w\in C([0,T_{0}];L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega))\cap C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega)),$
$\bullet\begin{cases}\partial_{t}w=\delta\Delta w+\nabla\cdot(m_{\varepsilon}(w)\nabla\varphi)\quad&\mathrm{a.e.\ in}\ \Omega\times(0,T_{0}],\\ \nabla w\cdot\boldsymbol{n}=0\quad&\mathrm{on}\ \partial\Omega\times(0,T_{0}],\\ w(0)=w_{0}\quad&\mathrm{in}\ L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega),\end{cases}$ (4.1)
$\bullet w(t)\geq 0,\ \|w(t)\|_{L^{1}(\Omega)}=\|w_{0}\|_{L^{1}(\Omega)}\quad t\in[0,T_{0}].$
This proposition is proved by the contraction mapping theorem. However, we need to take care with nonnegativity of functions because $m_{\varepsilon}(r)=(r+\varepsilon)^{\alpha}$ can not be definded for $r<-\varepsilon$. To overcome this, we inductively define a special contraction map depending a nonnegative function of a previous step (see Appendix). This idea is inspired by [14].
Proposition 4.2.
Let $w$ be a local solution in Proposition 4.1. Then $w$ can be extended globally in time.
Proof.
Let $w\in C([0,T_{0}];L^{2}\cap W^{1,p+1-\alpha}(\Omega))\cap C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega))$ be a nonnegative local solution to (4.1). Then multiplying the first equation of (4.1) by $w(t)$ for $t\in(0,T_{0}]$ and integrating in $\Omega$, we have
$\frac{d}{dt}\int_{\Omega}|w(t)|^{2}\,dx=\delta\int_{\Omega}(\Delta w(t))w(t)\,dx+\int_{\Omega}\nabla\cdot((w(t)+\varepsilon)^{\alpha}\nabla\varphi)w(t)\,dx.$
Note that thanks to $1+\alpha-2/d\leq p\leq 1+\alpha$ and the Sobolev embedding theorem, we have $w(t)\in W^{2,p+1-\alpha}(\Omega)\hookrightarrow H^{1}(\Omega)\hookrightarrow L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)$, thus the right hand side is well-defined. Since $\nabla w\cdot\boldsymbol{n}=0$ and $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega\times(0,T_{0}]$, we infer from integration by parts, Hölder’s inequality and Young’s inequality that
$\frac{d}{dt}\|w(t)\|_{L^{2}(\Omega)}^{2}$ $=-\delta\|\nabla w(t)\|_{L^{2}(\Omega)}^{2}-\int_{\Omega}(w(t)+\varepsilon)^{\alpha}\nabla\varphi\cdot\nabla w(t)\,dx$
$\leq-\delta\|\nabla w(t)\|_{L^{2}(\Omega)}^{2}+\|(w(t)+\varepsilon)^{\alpha}\nabla\varphi\|_{L^{2}(\Omega)}\|\nabla w(t)\|_{L^{2}(\Omega)}$
$\leq-\delta\|\nabla w(t)\|_{L^{2}(\Omega)}^{2}+\delta\|\nabla w(t)\|_{L^{2}(\Omega)}^{2}+\frac{1}{4\delta}\|(w(t)^{\alpha}+\varepsilon^{\alpha})\nabla\varphi\|_{L^{2}(\Omega)}^{2}$
$\leq\frac{1}{2\delta}\left(\|w(t)\|_{L^{2}(\Omega)}^{2\alpha}\|\nabla\varphi\|_{L^{\frac{2}{1-\alpha}}(\Omega)}^{2}+\varepsilon^{2\alpha}\|\nabla\varphi\|_{L^{2}(\Omega)}^{2}\right)$
$\leq\|w(t)\|_{L^{2}(\Omega)}^{2}+C(\delta,\alpha,\varepsilon,\varphi),$
where $C(\delta,\alpha,\varepsilon,\varphi)$ is a constant. Using Gronwall’s lemma, we obtain
$\|w(t)\|_{L^{2}(\Omega)}^{2}\leq\left[\|w(0)\|_{L^{2}(\Omega)}^{2}+C(\delta,\alpha,\varepsilon,\varphi)\right]e^{T_{0}}\quad\mathrm{for}\ t\in[0,T_{0}].$
Since $p+1-\alpha\leq 2$, it also follows
$\|w(t)\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}\leq|\Omega|^{\frac{1+\alpha-p}{2}}\left(\left[\|w(0)\|_{L^{2}(\Omega)}^{2}+C(\delta,\alpha,\varepsilon,\varphi)\right]e^{T_{0}}\right)^{\frac{p+1-\alpha}{2}}\quad\mathrm{for}\ t\in[0,T_{0}].$
Next, setting $f(w)\coloneq\nabla\cdot((w+\varepsilon)^{\alpha}\nabla\varphi)$, we have
$\|f(w)\|_{L^{p+1-\alpha}(\Omega)}$
$=\left\|\frac{\alpha\nabla w\cdot\nabla\varphi}{(w+\varepsilon)^{1-\alpha}}+(w+\varepsilon)^{\alpha}\Delta\varphi\right\|_{L^{p+1-\alpha}(\Omega)}$
$\leq\frac{\alpha\|\nabla\varphi\|_{L^{\infty}(\Omega)}}{\varepsilon^{1-\alpha}}\|\nabla w\|_{L^{p+1-\alpha}(\Omega)}+\|w\|_{L^{p+1-\alpha}(\Omega)}^{\alpha}\|\Delta\varphi\|_{L^{\frac{p+1-\alpha}{1-\alpha}}(\Omega)}+\varepsilon^{\alpha}\|\Delta\varphi\|_{L^{p+1-\alpha}(\Omega)}$
$\leq C(\alpha,\varepsilon,\varphi)(1+\|w\|_{W^{1,p+1-\alpha}(\Omega)}).$
By [16, Proposition 7.2.2], it follows that $\|w(t)\|_{W^{1,p+1-\alpha}(\Omega)}$ is also bounded in $0,T_{0}]$. Hence, combining this boundedness and Proposition 4.1, we can extend $w$ globally in time. ∎
In order to consider the Euler–Lagrange equation for $u_{\tau}^{k}$, we need to use the flow interchange lemma ([15]). In the following, we prepare some lemmas to adapt it in our case.
Fix $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ and $\delta>0$. Without loss of the generality, we assume $0\in\Omega$. Let $\{a_{n}\}_{n}$ be a monotonically decreasing sequence converging to $1$ as $n\to\infty$ and set $\Omega_{n}\coloneq\{a_{n}x:x\in\Omega\}$. By Lemma 2.8 with $\mu_{0}=u_{\tau}^{k-1}$ and $\mu_{1}=u_{\tau}^{k}$, there exist a vanishing sequence $\{b_{n}\}_{n}$ such that $\Omega_{[b_{n}]}\subset\Omega_{n}$ and a nonnegative function $\tilde{\rho}_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1])$ such that $\|\tilde{\rho}_{n}(t)\|_{L^{1}(\Omega_{n})}=1$ for $n\in\mathbb{N}$ and $t\in[0,1]$ and a function $\tilde{\phi}_{n}\in C^{\infty}(\overline{\Omega}_{n}\times[0,1])$. Notice that they satisfy
$W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}=\lim_{n\to\infty}\int_{0}^{1}\int_{\Omega_{n}}m_{\varepsilon}(\tilde{\rho_{n}})|\nabla\tilde{\phi}_{n}|^{2}\,dx\,dt.$
Then we define $\rho_{n}:\overline{\Omega}\times[0,1]\to[0,\infty)$ by $\rho_{n}(x,t)\coloneq\tilde{\rho}_{n}(a_{n}x,t)$ for $(x,t)\in\overline{\Omega}\times[0,1]$ and similarly, $\phi_{n}:\overline{\Omega}\times[0,1]\to\mathbb{R}$ by $\phi_{n}(x,t)\coloneq\tilde{\phi}_{n}(a_{n}x,t)$ for $(x,t)\in\overline{\Omega}\times[0,1]$. Note that $\rho_{n},\phi_{n}\in C^{\infty}(\overline{\Omega}\times[0,1])$, $\|\rho_{n}(t)\|_{L^{1}(\Omega)}=a_{n}^{-d}$ for $n\in\mathbb{N}$ and $t\in[0,1]$ and they satisfy
$W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}=\lim_{n\to\infty}a_{n}^{d}\int_{0}^{1}\int_{\Omega}m_{\varepsilon}(\rho_{n})|\nabla\phi_{n}|^{2}\,dx\,dt.$ (4.2)
Moreover, since $\tilde{\rho}_{n}(i)\to u_{\tau}^{k-1+i}$ in $L^{1}(\mathbb{R}^{d})$ for $i=0,1$ and $a_{n}\to 1$ as $n\to\infty$, we have $\rho_{n}(i)\to u_{\tau}^{k-1+i}$ in $L^{1}(\Omega)$ as $n\to\infty$ for $i=0,1$.
Fix $n\in\mathbb{N}$ and $t\in[0,1]$. Adapting Proposition 4.1 and Proposition 4.2 with $w_{0}=\rho_{n}(t)$, we have a solution $S_{z}\rho_{n}(t)$ satisfying
$\bullet S_{z}\rho_{n}(t)\in C^{\infty}(\overline{\Omega}\times[0,\infty)),$
$\bullet\partial_{z}(S_{z}\rho_{n}(t))=\delta\Delta(S_{z}\rho_{n}(t))+\nabla\cdot(m_{\varepsilon}(S_{z}\rho_{n}(t))\nabla\varphi)\quad\mathrm{in}\ \Omega\times[0,\infty),$ (4.3)
$\bullet\nabla S_{z}\rho_{n}(t)\cdot\boldsymbol{n}=0\quad\mathrm{on}\ \partial\Omega\times(0,\infty),$
$\bullet S_{z}\rho_{n}(t)\geq 0,\ \|S_{z}\rho_{n}(t)\|_{L^{1}(\Omega)}=\|\rho_{n}(t)\|_{L^{1}(\Omega)}=a_{n}^{-d}\quad\mathrm{for}\ z\in[0,\infty).$
Remark 4.3.
Define $\rho_{n}^{h}(t)\coloneq S_{ht}\rho_{n}(t)$ for $h\in(0,1)$. Due to the smoothness of $S_{z}\rho_{n}(t)$ and $\rho_{n}(t)$, $\rho_{n}^{h}(t)$ is $t$-differentiable in $(0,1)$. Since for each $n\in\mathbb{N}$, $\|\rho_{n}^{h}(t)\|_{L^{1}(\Omega)}=a_{n}^{-d}$ for all $t\in[0,1]$ and $h>0$, we have
$\int_{\Omega}\partial_{t}\rho_{n}^{h}(t)\,dx=0.$
Hence for fixed $n\in\mathbb{N}$ and $t\in[0,1]$, as in the proof of Lemma 2.8, we can find a unique solution $\phi_{n}^{h}(t)\in H^{1}(\Omega)$ satisfying
$\partial_{t}\rho_{n}^{h}(t)=-\nabla\cdot(m_{\varepsilon}(\rho_{n}^{h}(t))\nabla\phi_{n}^{h}(t))\quad\mathrm{in}\ \Omega,\quad\nabla\phi_{n}^{h}(t)\cdot\boldsymbol{n}=0\quad\mathrm{on}\ \partial\Omega.$ (4.4)
and the elliptic regularity theorem yields $\phi_{n}^{h}\in C^{\infty}(\overline{\Omega}\times[0,1])$. In addition, due to the smoothness of $\rho_{n}^{h}(t)=S_{ht}\rho_{n}(t)$ with respect to $h$, $\phi_{n}^{h}$ is $h$-differentiable.
Then we define
$\boldsymbol{A}_{n}^{h}(t)\coloneq\int_{\Omega}m_{\varepsilon}(\rho_{n}^{h}(t))|\nabla\phi_{n}^{h}(t)|^{2}\,dx,$
and $\boldsymbol{V}_{\delta}:\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})\to(-\infty,\infty]$ by
$\boldsymbol{V}_{\delta}(u)\coloneq\int_{\Omega}u\varphi\,dx+\delta\boldsymbol{U}_{\varepsilon}(u),$
where $\boldsymbol{U}_{\varepsilon}$ is defined in Lemma 2.9. Note that if $u\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega)$, then $\boldsymbol{V}_{\delta}(u)<\infty$.
Next, we establish the Gronwall type inequality, and it is obtained by the same argument for [15, Lemma 4.2, Proposition 4.6]. Note that we use Lemma 2.10 in the proof of the following lemma, thus the convexity of domain $\Omega$ is required here.
Lemma 4.4.
For $n\in\mathbb{N},t\in[0,1],h>0$, it holds
$\frac{1}{2}\partial_{h}\boldsymbol{A}_{n}^{h}(t)+\partial_{t}\boldsymbol{V}_{\delta}(\rho_{n}^{h}(t))\leq-\lambda_{\delta}t\boldsymbol{A}_{n}^{h}(t),$ (4.5)
where
$\lambda_{\delta}$ $\coloneq-\frac{1}{2\delta}\|\nabla\varphi\|_{L^{\infty}(\Omega)}^{2}\sup_{r\geq 0}|m_{\varepsilon}(r)m_{\varepsilon}^{\prime\prime}(r)|-\|\nabla^{2}\varphi\|_{L^{\infty}(\Omega)}\sup_{r\geq 0}|m_{\varepsilon}^{\prime}(r)|$
$=-\frac{1}{2\delta}\|\nabla\varphi\|_{L^{\infty}(\Omega)}^{2}\frac{\alpha(1-\alpha)}{\varepsilon^{2(1-\alpha)}}-\|\nabla^{2}\varphi\|_{L^{\infty}(\Omega)}\frac{\alpha}{\varepsilon^{1-\alpha}}\leq 0,$
and
$\|\nabla^{2}\varphi\|_{L^{\infty}(\Omega)}=\left\|\sum_{i,j=1}^{d}\left|\frac{\partial^{2}\varphi}{\partial_{i}x\partial_{j}x}\right|\right\|_{L^{\infty}(\Omega)}.$
The following lemma is the estimate like the evolution variational inequality for the weighted Wasserstein distance with respect to the functional $\boldsymbol{U}_{\varepsilon}$. The proof is based on [15, Lemma 3.3, Lemma 4.2] and similar to the proof of Lemma 4.7. We can easily check the required properties of $\boldsymbol{U}_{\varepsilon}$ (see [15, Definition 2]) due to the propertty of the Neumann heat semigroup, Lemma 2.8 and Lemma 2.9.
Lemma 4.5.
Let $u_{\tau}^{k}\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega)$. Then it holds
$\frac{1}{2}\limsup_{h\downarrow 0}\frac{W_{m_{\varepsilon}}(e^{h\Delta}u_{\tau}^{k},u_{\tau}^{k-1})^{2}-W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{h}\leq\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k-1})-\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k}),$
where $e^{h\Delta}$ is the Neumann heat semigroup on $\Omega$.
The regularity of the minimizer $(u_{\tau}^{k},v_{\tau}^{k})\in X$ is not enough to get the weak formulation (the Euler–Lagrange equation). Hence we need to improve the regularity of minimizers enough to converge to the weak formulation.
Lemma 4.6.
Let $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be the minimizer of (3.1). Then $(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in H^{1}(\Omega)$ and $\Delta v_{\tau}^{k}-v_{\tau}^{k}+u_{\tau}^{k}\in L^{2}(\Omega)$. If $1+\alpha-2/d<p<1+\alpha$, or $p=1+\alpha-2/d$ and $\chi>0$ is small enough, then $u_{\tau}^{k}\in L^{2}(\Omega)$. In addition, there exists a constant $C_{0}=C_{0}(\alpha,p,d,\chi)>0$ such that the following estimates hold
$\bullet\,$ $\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}^{k}-v_{\tau}^{k}+u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}$
$\leq\frac{\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k-1})-\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k})}{\tau}+\frac{\|v_{\tau}^{k-1}\|_{H^{1}(\Omega)}^{2}-\|v_{\tau}^{k}\|_{H^{1}(\Omega)}^{2}}{\tau}$
$\quad+\|\nabla v_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}+C_{0}\left(1+\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{\frac{p+1-\alpha}{p-\alpha}}\right),$ (4.6)
$\bullet\,$ $\|u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}\leq\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+C_{0}\left(1+\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{\frac{p+1-\alpha}{p-\alpha}}\right)$ (4.7)
$\hskip 256.0748pt\mathrm{if}\ 1+\alpha-\frac{2}{d}\leq p<1+\alpha.$
Proof.
The proof is almost same for [4, Proposition 8] although the energy functional and the distance (the weighted Wasserstein distance) are different from theirs. Hence, we only point out the key idea for proving $(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in H^{1}(\Omega)$. Considering the Neumann heat equation with initial data $u_{\tau}^{k}$, we have formally
$\frac{d}{dt}E(e^{t\Delta}u_{\tau}^{k},v_{\tau}^{k})$ $=\frac{p}{\chi(p-\alpha)}\int_{\Omega}(e^{t\Delta}u_{\tau}^{k})^{p-\alpha}(\Delta e^{t\Delta}u_{\tau}^{k})\,dx-\int_{\Omega}(\Delta e^{t\Delta}u_{\tau}^{k})v_{\tau}^{k}\,dx$
$=-\frac{p}{\chi}\int_{\Omega}(e^{t\Delta}u_{\tau}^{k})^{p-1-\alpha}|\nabla e^{t\Delta}u_{\tau}^{k}|^{2}\,dx+\int_{\Omega}\nabla e^{t\Delta}u_{\tau}^{k}\cdot\nabla v_{\tau}^{k}\,dx$
$=-\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(e^{t\Delta}u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}\nabla e^{t\Delta}u_{\tau}^{k}\cdot\nabla v_{\tau}^{k}\,dx.$
By Hölder’s inequality and Young’s inequality, the second term can be absorbed by the first term. Using (3.11) and Lemma 4.5, we obtain
$\|\nabla(e^{t\Delta}u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}\leq C\left(\frac{\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k-1})-\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k})}{\tau}+\|\nabla v_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}\right)<\infty\quad\mathrm{for}\ t\in(0,1).$
Since $e^{t\Delta}u_{\tau}^{k}\in C([0,1];L^{p+1-\alpha}(\Omega))$, that is, $(e^{t\Delta}u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in C([0,1];L^{2}(\Omega))$, combining this with the above boundedness for $t\in(0,1)$, we have $\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in L^{2}(\Omega)$, then $(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in H^{1}(\Omega)$. ∎
Next lemma is the estimate like the evolution variational inequality for the weighted Wasserstein distance with respect to the functional $\boldsymbol{V}_{\delta}$. The proof is fundamentally based on [15, Lemma 3.3, Proposition 4.6], but our minimizers have the lower regularity than theirs. Thus we check the required properties of $\boldsymbol{V}_{\delta}$ (see [15, Definition 2]). This lemma helps us to obtain the weak formulation (Lemma 4.8).
Lemma 4.7.
Let $1+\alpha-2/d\leq p\leq 1+\alpha$ and $u_{\tau}^{k}\in L^{p+1-\alpha}\cap\mathcal{P}(\Omega)$ with $(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in H^{1}(\Omega)$. Then it holds
$\frac{1}{2}\limsup_{h\downarrow 0}\frac{W_{m_{\varepsilon}}(S_{h}u_{\tau}^{k},u_{\tau}^{k-1})^{2}-W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{h}+\frac{\lambda_{\delta}}{2}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}+\boldsymbol{V}_{\delta}(u_{\tau}^{k})\leq\boldsymbol{V}_{\delta}(u_{\tau}^{k-1}),$
where $\lambda_{\delta}$ is defined in Lemma 4.4.
Proof.
If $\varphi=0$, that is, $\lambda_{\delta}=0$ then we complete the proof by Lemma 4.5. Thus we can assume that $\varphi\not\equiv 0$ and then $\lambda_{\delta}\neq 0$. Since $\rho_{n}^{0}(t)=S_{0}\rho_{n}(t)=\rho_{n}(t)$ in $\Omega$, the definition of $\boldsymbol{A}_{n}^{h}$ and (4.2) imply
$W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}=\lim_{n\to\infty}a_{n}^{d}\int_{0}^{1}\boldsymbol{A}_{n}^{0}(t)\,dt.$ (4.8)
Since $\|\rho_{n}^{h}(t)\|_{L^{1}(\Omega)}=\|\rho_{n}^{0}(t)\|_{L^{1}(\Omega)}=\|\rho_{n}(t)\|_{L^{1}(\Omega)}=a_{n}^{-d}$ for all $n\in\mathbb{N},h\in(0,1),t\in[0,1]$ and $\boldsymbol{U}_{\varepsilon}\geq 0$, we have
$\boldsymbol{V}_{\delta}(\rho_{n}^{h}(t))$ $=\int_{\Omega}\rho_{n}^{h}(t)\varphi\,dx+\delta\boldsymbol{U}_{\varepsilon}(\rho_{n}^{h}(t))$
$\geq-\|\varphi\|_{L^{\infty}(\Omega)}\|\rho_{n}^{h}(t)\|_{L^{1}(\Omega)}$
$=-a_{n}^{-d}\|\varphi\|_{L^{\infty}(\Omega)}.$
Since $a_{n}\to 1$ as $n\to\infty$, that is, $\{a_{n}\}_{n}$ is bounded, there exists a constant $L>0$ such that $\boldsymbol{V}_{\delta}(\rho_{n}^{h}(t))\geq-L$ for all $n\in\mathbb{N},h\in(0,1),t\in[0,1]$.
Multiplying (4.5) by $e^{2\lambda_{\delta}th}$ and integrating with respect to $t\in[0,1]$, further using integration by parts, we have
$\frac{1}{2}\partial_{h}\int_{0}^{1}e^{2\lambda_{\delta}th}\boldsymbol{A}_{n}^{h}(t)\,dt$ $\leq-\int_{0}^{1}e^{2\lambda_{\delta}th}\partial_{t}(\boldsymbol{V}_{\delta}(\rho_{n}^{h}(t))+L)\,dt$
$=\boldsymbol{V}_{\delta}(\rho_{n}^{h}(0))+L-e^{2\lambda_{\delta}h}(\boldsymbol{V}_{\delta}(\rho_{n}^{h}(1))+L)$
$\quad+2\lambda_{\delta}h\int_{0}^{1}e^{2\lambda_{\delta}th}(\boldsymbol{V}_{n,\delta}(\rho_{n}^{h}(t))+L)\,dt.$
Since $\boldsymbol{V}_{\delta}(\rho_{n}^{h}(t))+L\geq 0$ for any $t\in[0,1]$ and $\lambda_{\delta}<0$, it follows
$\frac{1}{2}\partial_{h}\int_{0}^{1}e^{2\lambda_{\delta}th}\boldsymbol{A}_{n}^{h}(t)\,dt\leq\boldsymbol{V}_{n,\delta}(\rho_{n}^{h}(0))+L-e^{2\lambda_{\delta}h}(\boldsymbol{V}_{n,\delta}(\rho_{n}^{h}(1))+L).$
Observe that $\rho_{n}^{h}(0)=\rho_{n}(0)$, integrating over $(0,h)$, we have
$\frac{1}{2}\int_{0}^{1}e^{2\lambda_{\delta}th}\boldsymbol{A}_{n}^{h}(t)\,dt\leq\frac{1}{2}\int_{0}^{1}\boldsymbol{A}_{n}^{0}\,dt+h(\boldsymbol{V}_{\delta}(\rho_{n}(0))+L)-\int_{0}^{h}e^{2\lambda_{\delta}s}(\boldsymbol{V}_{\delta}(\rho_{n}^{s}(1))+L)\,ds.$
Here the function $s\mapsto\boldsymbol{V}_{\delta}(\rho_{n}^{s}(1))+L$ is nonincreasing. Indeed, calculating the derivative and using (4.3) with $t=1$ and $U_{\varepsilon}^{\prime\prime}(r)m_{\varepsilon}(r)=1$ for $r\geq 0$, we obtain
$\frac{d}{ds}\boldsymbol{V}_{\delta}(\rho_{n}^{s}(1))$ $=\frac{d}{ds}\left(\int_{\Omega}\rho_{n}^{s}(1)\varphi\,dx+\delta\int_{\Omega}U_{\varepsilon}(\rho_{n}^{s}(1))\,dx\right)$
$=\int_{\Omega}(\nabla\cdot(m_{\varepsilon}(\rho_{n}^{s}(1))\nabla\varphi)+\delta\Delta\rho_{n}^{s}(1))\varphi\,dx$
$\quad+\delta\int_{\Omega}U_{\varepsilon}^{\prime}(\rho_{n}^{s}(1))(\nabla\cdot(m_{\varepsilon}(\rho_{n}^{s}(1))\nabla\varphi)+\delta\Delta\rho_{n}^{s}(1))\,dx$
$=-\int_{\Omega}(m_{\varepsilon}(\rho_{n}^{s}(1))\nabla\varphi+\delta\nabla\rho_{n}^{s}(1))\cdot\nabla\varphi\,dx$
$\quad-\delta\int_{\Omega}U_{\varepsilon}^{\prime\prime}(\rho_{n}^{s}(1))\nabla\rho_{n}^{s}(1)\cdot(m_{\varepsilon}(\rho_{n}^{s}(1))\nabla\varphi+\delta\nabla\rho_{n}^{s}(1))\,dx$
$=-\int_{\Omega}m_{\varepsilon}(\rho_{n}^{s}(1))|\nabla\varphi|^{2}\,dx-\delta\int_{\Omega}\nabla\rho_{n}^{s}(1)\cdot\nabla\varphi\,dx$
$\quad-\delta\int_{\Omega}\nabla\rho_{n}^{s}(1)\cdot\nabla\varphi\,dx-\delta^{2}\int_{\Omega}\frac{|\nabla\rho_{n}^{s}(1)|^{2}}{m(\rho_{n}^{s}(1))}\,dx$
$=-\int_{\Omega}\left|m_{\varepsilon}(\rho_{n}^{s}(1))^{\frac{1}{2}}\nabla\varphi+\frac{\delta\nabla\rho_{n}^{s}(1)}{m_{\varepsilon}(\rho_{n}^{s}(1))^{\frac{1}{2}}}\right|^{2}\,dx\leq 0.$
Thus it follows
$\frac{1}{2}\int_{0}^{1}e^{2\lambda_{\delta}th}\boldsymbol{A}_{n}^{h}(t)\,dt\leq\frac{1}{2}\int_{0}^{1}\boldsymbol{A}_{n}^{0}(t)\,dt+h(\boldsymbol{V}_{\delta}(\rho_{n}(0))+L)-\frac{1-e^{2\lambda_{\delta}h}}{-2\lambda_{\delta}}(\boldsymbol{V}_{\delta}(\rho_{n}^{h}(1))+L).$ (4.9)
On the other hand, for a decreasing function $\theta\in C^{1}([0,1])$ with $\theta>0$, we define a increasing function
$\tilde{\theta}(t)\coloneq\left[\int_{0}^{1}\frac{dz}{\theta(z)}\right]^{-1}\int_{0}^{t}\frac{dz}{\theta(z)}\quad\mathrm{for}\ t\in[0,1],$
and denote $\tilde{\theta}^{-1}$ an inverse function of $\tilde{\theta}$, that is, $\tilde{\theta}\circ\tilde{\theta}^{-1}(t)=\tilde{\theta}^{-1}\circ\tilde{\theta}(t)=t$ for $t\in[0,1]$. Then the pair $\left(\rho_{n}^{h}(\cdot,\tilde{\theta}^{-1}(\cdot)),m_{\varepsilon}(\rho_{n}^{h}(\cdot,\tilde{\theta}^{-1}(\cdot)))\nabla\{(\tilde{\theta}^{-1})^{\prime}(\cdot)\phi_{n}^{h}(\cdot,\tilde{\theta}^{-1}(\cdot))\}\right)$ belongs to $CE(0,1;\rho_{n}(0),\rho_{n}^{h}(1))$. Indeed, by (4.4), we have
$\partial_{t}[\rho_{n}^{h}(x,\tilde{\theta}^{-1}(t))]$ $=(\partial_{t}\rho_{n}^{h})(x,\tilde{\theta}^{-1}(t))(\tilde{\theta}^{-1})^{\prime}(t)$
$=\left[-\nabla\cdot\left\{m_{\varepsilon}\left(\rho_{n}^{h}(x,\tilde{\theta}^{-1}(t))\right)\nabla\phi_{n}^{h}(x,\tilde{\theta}^{-1}(t))\right\}\right](\tilde{\theta}^{-1})^{\prime}(t)$
$=-\nabla\cdot\left\{m_{\varepsilon}\left(\rho_{n}^{h}(x,\tilde{\theta}^{-1}(t))\right)\nabla\left((\tilde{\theta}^{-1})^{\prime}(t)\phi_{n}^{h}(x,\tilde{\theta}^{-1}(t))\right)\right\},$
hence they satisfy the continuity equation. In addition, since $\tilde{\theta}(0)=0$ and $\tilde{\theta}(1)=1$, it also holds $\tilde{\theta}^{-1}(0)=0$ and $\tilde{\theta}^{-1}(1)=1$, thus we have
$\rho_{n}^{h}(x,\tilde{\theta}^{-1}(0))=\rho_{n}^{h}(x,0)=\rho_{n}(x,0),\quad\rho_{n}^{h}(x,\tilde{\theta}^{-1}(1))=\rho_{n}^{h}(x,1)\quad\mathrm{for}\ x\in\Omega.$
Note that $\rho_{n}^{h}(1)=S_{h}\rho_{n}(1)$. By the definition of the weighted Wasserstein distance (see Definition 2.3) and the change of variables, it follows
$W_{m_{\varepsilon}}(\rho_{n}(0),S_{h}\rho_{n}(1))^{2}$ $\leq\int_{0}^{1}\int_{\Omega}\left((\tilde{\theta}^{-1})^{\prime}(z)\right)^{2}m_{\varepsilon}\left(\rho_{n}^{h}(x,\tilde{\theta}^{-1}(z))\right)|\nabla\phi_{n}^{h}(x,\tilde{\theta}^{-1}(z))|^{2}\,dx\,dz$
$=\int_{0}^{1}\int_{\Omega}\left((\tilde{\theta}^{-1})^{\prime}(\tilde{\theta}(t))\right)m_{\varepsilon}(\rho_{n}^{h}(x,t))|\nabla\phi_{n}^{h}(x,t)|^{2}\,dx\,dt$
$=\int_{0}^{1}\frac{dr}{\theta(r)}\int_{0}^{1}\theta(t)\boldsymbol{A}_{n}^{h}(t)\,dt,$
where we used
$(\tilde{\theta}^{-1})^{\prime}(\tilde{\theta}(t))=\frac{1}{\tilde{\theta}^{\prime}(t)}=\int_{0}^{1}\frac{dr}{\theta(r)}\theta(t).$
Hence, choosing $\theta(t)=e^{2\lambda_{\delta}th}$, we obtain
$W_{m_{\varepsilon}}(\rho_{n}(0),S_{h}\rho_{n}(1))^{2}\leq\frac{e^{-2\lambda_{\delta}h}-1}{-2\lambda_{\delta}h}\int_{0}^{1}e^{2\lambda_{\delta}th}\boldsymbol{A}_{n}^{h}(t)\,dt.$
Combining the above with (4.9), we have
$\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}W_{m_{\varepsilon}}(\rho_{n}(0),S_{h}\rho_{n}(1))^{2}$
$\leq\frac{1}{2}\int_{0}^{1}\boldsymbol{A}_{n}^{0}(t)\,dt+h(\boldsymbol{V}_{\delta}(\rho_{n}(0))+L)-\frac{1-e^{2\lambda_{\delta}h}}{-2\lambda_{\delta}}(\boldsymbol{V}_{\delta}(S_{\delta,h}\rho_{n}(1))+L).$ (4.10)
Let $S_{h}u_{\tau}^{k}$ be a solution in Propositions 4.1 and 4.2 with $w_{0}=u_{\tau}^{k}$. We will show
$S_{h}\rho_{n}(1)\to S_{h}u_{\tau}^{k}\ \mathrm{in}\ L^{2}(\Omega)\quad\mathrm{as}\ n\to\infty\ \mathrm{for}\ h\in(0,1).$
Since $y\coloneq S_{h}\rho_{n}(1)-S_{h}u_{\tau}^{k}$ satisfies the following equation
$\begin{cases}\partial_{h}y=\delta\Delta y+\nabla\cdot[(m_{\varepsilon}(S_{h}\rho_{n}(1))-m_{\varepsilon}(S_{h}u_{\tau}^{k}))\nabla\varphi]&\mathrm{in}\ \Omega\times(0,1),\\ \nabla y\cdot\boldsymbol{n}=0&\mathrm{on}\ \partial\Omega\times(0,1),\\ y(0)=\rho_{n}(1)-u_{\tau}^{k}&\mathrm{in}\ L^{2}(\Omega),\end{cases}$
multiplying the first equation by $y$ and integrating in $\Omega$, we have for $h\in(0,1)$
$\int_{\Omega}(\partial_{h}y)y\,dx=\int_{\Omega}[\delta\Delta y+\nabla\cdot\{(m_{\varepsilon}(S_{h}\rho_{n}(1))-m_{\varepsilon}(S_{h}u_{\tau}^{k}))\nabla\varphi\}]y\,dx.$
Since $\nabla y\cdot\boldsymbol{n}=0$ and $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega\times(0,1)$, we infer from integration by parts that
$\frac{1}{2}\partial_{h}\|y(h)\|_{L^{2}(\Omega)}^{2}$ $=-\delta\|\nabla y\|_{L^{2}(\Omega)}^{2}-\int_{\Omega}(m_{\varepsilon}(S_{h}\rho_{n}(1))-m_{\varepsilon}(S_{h}u_{\tau}^{k}))\nabla\varphi\cdot\nabla y\,dx$
$\leq-\delta\|\nabla y\|_{L^{2}(\Omega)}^{2}+\|\nabla\varphi\|_{L^{\infty}(\Omega)}\|\nabla y\|_{L^{2}(\Omega)}\|m_{\varepsilon}(S_{h}\rho_{n}(1))-m_{\varepsilon}(S_{h}u_{\tau}^{k})\|_{L^{2}(\Omega)}.$
Using the Lipschitz continuity of $m_{\varepsilon}$ and Young’s inequality, we obtain
$\frac{1}{2}\partial_{h}\|y(h)\|_{L^{2}(\Omega)}^{2}$ $\leq-\delta\|\nabla y\|_{L^{2}(\Omega)}^{2}+\|\nabla\varphi\|_{L^{\infty}(\Omega)}\|\nabla y\|_{L^{2}(\Omega)}\frac{\alpha}{\varepsilon^{1-\alpha}}\|S_{h}\rho_{n}(1)-S_{h}u_{\tau}^{k}\|_{L^{2}(\Omega)}$
$\leq C\|S_{h}\rho_{n}(1)-S_{h}u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}=C\|y(h)\|_{L^{2}(\Omega)}^{2},$
where $C=C(\alpha,\varepsilon,\delta,\varphi)$ is a constant. By Gronwall’s lemma, it follows
$\|y(h)\|_{L^{2}(\Omega)}^{2}\leq e^{2C}\|y(0)\|_{L^{2}(\Omega)}^{2}=e^{2C}\|\rho_{n}(1)-u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}\quad\mathrm{for}\ h\in(0,1).$
Since $\rho_{n}(1)\to u_{\tau}^{k}$ in $L^{2}(\Omega)$ as $n\to\infty$ (see Lemma 2.8), we conclude that $S_{h}\rho_{n}(1)$ converges to $S_{h}u_{\tau}^{k}$ in $L^{2}(\Omega)$ as $n\to\infty$ for $h\in(0,1)$. This convergence also implies that $S_{h}\rho_{n}(1)$ converges to $S_{h}u_{\tau}^{k}$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ as $n\to\infty$. Hence, combining this with the lower semicontinuity of $\boldsymbol{U}_{\varepsilon}$ (Lemma 2.9), we see
$\boldsymbol{V}_{\delta}(S_{h}u_{\tau}^{k})\leq\liminf_{n\to\infty}\boldsymbol{V}_{\delta}(S_{h}\rho_{n}(1)).$
Moreover by Lemma 2.8, we also have $\rho_{n}(0)\to u_{\tau}^{k-1}$ in $L^{1}\cap L^{2-\alpha}(\Omega)$ as $n\to\infty$, which thus yields
$\lim_{n\to\infty}\boldsymbol{V}_{\delta}(\rho_{n}(0))=\boldsymbol{V}_{\delta}(u_{\tau}^{k-1}).$
Therefore by Lemma 2.6 and (4.8), letting $n\to\infty$ in (4.1), we obtain
$\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}W_{m_{\varepsilon}}(u_{\tau}^{k-1},S_{h}u_{\tau}^{k})^{2}$
$\leq\frac{1}{2}W_{m_{\varepsilon}}(u_{\tau}^{k-1},u_{\tau}^{k})^{2}+h(\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})+L)-\frac{1-e^{2\lambda_{\delta}h}}{-2\lambda_{\delta}}(\boldsymbol{V}_{\delta}(S_{h}u_{\tau}^{k})+L),$
then
$\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}\frac{W_{m_{\varepsilon}}(S_{h}u_{\tau}^{k},u_{\tau}^{k-1})^{2}-W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{h}+\frac{1}{h}\left(\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}-\frac{1}{2}\right)W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}$
$\leq\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})+L-\frac{1-e^{2\lambda_{\delta}h}}{-2\lambda_{\delta}h}(\boldsymbol{V}_{\delta}(S_{h}u_{\tau}^{k})+L).$
Since
$\lim_{h\downarrow 0}\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}=\frac{1}{2},\quad\lim_{h\downarrow 0}\frac{1-e^{2\lambda_{\delta}h}}{-2\lambda_{\delta}h}=1,\quad\lim_{h\downarrow 0}\frac{1}{h}\left(\frac{-\lambda_{\delta}h}{e^{-2\lambda_{\delta}h}-1}-\frac{1}{2}\right)=\frac{\lambda_{\delta}}{2},$
and
$\boldsymbol{V}_{\delta}(u_{\tau}^{k})\leq\liminf_{h\downarrow 0}\boldsymbol{V}_{\delta}(S_{h}u_{\tau}^{k})$
due to $S_{h}u_{\tau}^{k}\to u_{\tau}^{k}$ in $L^{1}(\Omega)$ as $h\to 0$ (Proposition 4.1), we conclude that
$\frac{1}{2}\limsup_{h\downarrow 0}\frac{W_{m_{\varepsilon}}(S_{h}u_{\tau}^{k},u_{\tau}^{k-1})^{2}-W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}{h}+\frac{\lambda_{\delta}}{2}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}\leq\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})-\boldsymbol{V}_{\delta}(u_{\tau}^{k}).$
The proof is completed. ∎
4.2 A discrete type of weak formulations
First, we obtain the inequality like the Euler–Lagrange equation of the first equation of the Keller–Segel system (1.1).
Lemma 4.8.
Let $p=1+\alpha,(u_{\tau}^{k-1},v_{\tau}^{k-1})\in X$ and $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be a minimizer of (3.1). Then
$\frac{\boldsymbol{V}_{\delta}(u_{\tau}^{k})-\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})}{\chi}$ $\leq-\tau\left[\frac{p}{\chi}\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla u_{\tau}^{k}\cdot\nabla\varphi\,dx-\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\right]$
$\quad-\tau\lambda_{\delta}(E(u_{\tau}^{k-1},v_{\tau}^{k-1})-E(u_{\tau}^{k},v_{\tau}^{k}))+\tau\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx.$
Proof.
In this proof, we write $\|\cdot\|_{L^{q}(\Omega)}=\|\cdot\|_{q}$ for $q\in[1,\infty]$. Observe that when $p=1+\alpha$, $u_{\tau}^{k}\in H^{1}(\Omega)$ by Lemma 4.6. Let $S_{t}u_{\tau}^{k}$ be a nonnegative solution to (4.1) with $w_{0}=u_{\tau}^{k}$. Note that $S_{t}u_{\tau}^{k}\in C((0,T];H^{2}(\Omega))\cap C^{1}((0,T];L^{2}(\Omega))$ and $S_{t}u_{\tau}^{k}\to u_{\tau}^{k}$ in $L^{1}\cap H^{1}(\Omega)$ as $t\to 0$. Then for $t>0$, using (4.1) and integration by parts, we have
$\frac{d}{dt}E(S_{t}u_{\tau}^{k},v_{\tau}^{k})$ $=\frac{1}{\chi}\int_{\Omega}p(S_{t}u_{\tau}^{k})(\partial_{t}S_{t}u_{\tau}^{k})\,dx-\int_{\Omega}v_{\tau}^{k}(\partial_{t}S_{t}u_{\tau}^{k})\,dx$
$=\frac{p}{\chi}\int_{\Omega}S_{t}u_{\tau}^{k}[\nabla\cdot(m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla\varphi)+\delta\Delta S_{t}u_{\tau}^{k}]\,dx$
$\quad-\int_{\Omega}v_{\tau}^{k}[\nabla\cdot(m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla\varphi)+\delta\Delta S_{t}u_{\tau}^{k}]\,dx$
$=-\frac{p}{\chi}\int_{\Omega}(\nabla S_{t}u_{\tau}^{k})\cdot m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla\varphi\,dx-\frac{p\delta}{\chi}\int_{\Omega}|\nabla S_{t}u_{\tau}^{k}|^{2}\,dx$
$\quad+\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx+\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla S_{t}u_{\tau}^{k}\,dx$
$\leq-\frac{p}{\chi}\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla S_{t}u_{\tau}^{k}\cdot\nabla\varphi\,dx$
$\quad+\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx+\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla S_{t}u_{\tau}^{k}\,dx.$
By integrating over $(0,t)$, it follows
$E(S_{t}u_{\tau}^{k},v_{\tau}^{k})-E(u_{\tau}^{k},v_{\tau}^{k})$ $\leq-\frac{p}{\chi}\int_{0}^{t}\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla S_{t}u_{\tau}^{k}\cdot\nabla\varphi\,dx\,dt$ (4.11)
$\quad+\int_{0}^{t}\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\,dt+\delta\int_{0}^{t}\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla S_{t}u_{\tau}^{k}\,dx\,dt.$
Since $S_{t}u_{\tau}^{k}\to u_{\tau}^{k}$ in $H^{1}(\Omega)$ as $t\to 0$, it immediately follows that
$\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla S_{t}u_{\tau}^{k}\,dx\to\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx\quad\mathrm{as}\ t\to 0.$
We will show
$\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla S_{t}u_{\tau}^{k}\cdot\nabla\varphi\,dx\to\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla u_{\tau}^{k}\cdot\nabla\varphi\,dx\quad\mathrm{as}\ t\to 0,$ (4.12)
$\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\to\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\quad\mathrm{as}\ t\to 0.$ (4.13)
First, using the Lipschitz continuity of $m_{\varepsilon}$ and Hölder’s ineqality, we have
$\left|\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla S_{t}u_{\tau}^{k}\cdot\nabla\varphi\,dx-\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla u_{\tau}^{k}\cdot\nabla\varphi\,dx\right|$
$\leq\int_{\Omega}|m_{\varepsilon}(S_{t}u_{\tau}^{k})-m_{\varepsilon}(u_{\tau}^{k})||\nabla S_{t}u_{\tau}^{k}||\nabla\varphi|\,dx+\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})|\nabla S_{t}u_{\tau}^{k}-\nabla u_{\tau}^{k}||\nabla\varphi|\,dx$
$\leq\int_{\Omega}\frac{\alpha}{\varepsilon^{1-\alpha}}|S_{t}u_{\tau}^{k}-u_{\tau}^{k}||\nabla S_{t}u_{\tau}^{k}||\nabla\varphi|\,dx+\|m_{\varepsilon}(u_{\tau}^{k})\nabla\varphi\|_{2}\|\nabla S_{t}u_{\tau}^{k}-\nabla u_{\tau}^{k}\|_{2}$
$\leq\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|S_{t}u_{\tau}^{k}-u_{\tau}^{k}\|_{2}\|\nabla S_{t}u_{\tau}^{k}\|_{2}+\|m_{\varepsilon}(u_{\tau}^{k})\nabla\varphi\|_{2}\|\nabla S_{t}u_{\tau}^{k}-\nabla u_{\tau}^{k}\|_{2}\to 0\quad\mathrm{as}\ t\to 0,$
where we used $S_{t}u_{\tau}^{k}\to u_{\tau}^{k}$ in $H^{1}(\Omega)$. Next, the similar argument yields
$\left|\int_{\Omega}m_{\varepsilon}(S_{t}u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx-\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\right|$
$\leq\int_{\Omega}|m_{\varepsilon}(S_{t}u_{\tau}^{k})-m_{\varepsilon}(u_{\tau}^{k})||\nabla v_{\tau}^{k}||\nabla\varphi|\,dx$
$\leq\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|S_{t}u_{\tau}^{k}-u_{\tau}^{k}\|_{2}\|\nabla v_{\tau}^{k}\|_{2}\to 0\quad\mathrm{as}\ t\to 0.$
Since $(S_{t}u_{\tau}^{k},v_{\tau}^{k})\in X$, by (3.11), we have
$0\leq\frac{1}{2\tau\chi}[W_{m_{\varepsilon}}(S_{t}u_{\tau}^{k},u_{\tau}^{k-1})^{2}-W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}]+E(S_{t}u_{\tau}^{k},v_{\tau}^{k})-E(u_{\tau}^{k},v_{\tau}^{k}).$
Dividing by $t>0$ and letting $t\to 0$, we infer from Lemma 4.7, (4.11), (4.12) and (4.13) that
$0$ $\leq-\frac{\lambda_{\delta}}{2\tau\chi}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}+\frac{1}{\tau\chi}[\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})-\boldsymbol{V}_{\delta}(u_{\tau}^{k})]$
$\quad-\frac{p}{\chi}\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla u_{\tau}^{k}\cdot\nabla\varphi\,dx+\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx+\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx.$
Further using (3.11) with $(\tilde{u},\tilde{v})=(u_{\tau}^{k-1},v_{\tau}^{k-1})$:
$\frac{1}{2\tau\chi}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}\leq E(u_{\tau}^{k-1},v_{\tau}^{k-1})-E(u_{\tau}^{k},v_{\tau}^{k}),$
note that $\lambda_{\delta}\geq 0$, we conclude that
$\frac{\boldsymbol{V}_{\delta}(u_{\tau}^{k})-\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})}{\chi}$ $\leq-\tau\left[\frac{p}{\chi}\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla u_{\tau}^{k}\cdot\nabla\varphi\,dx-\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\right]$
$\quad-\tau\lambda_{\delta}(E(u_{\tau}^{k-1},v_{\tau}^{k-1})-E(u_{\tau}^{k},v_{\tau}^{k}))+\tau\delta\int_{\Omega}\nabla\ v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx.$
The proof is completed. ∎
Next, we obtain the Euler–Lagrange equation of the second equation of the Keller–Segel system (1.1). Moreover we see that $v_{\tau}^{k}$ satisfies the Neumann boundary condition.
Lemma 4.9.
Let $p\geq 1+\alpha-2/d$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Let $v_{\tau}^{k-1}\in H^{1}(\Omega)$ and $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be a minimizer of (3.1). Then it holds
$\int_{\Omega}\frac{v_{\tau}^{k}-v_{\tau}^{k-1}}{\tau}\zeta\,dx+\int_{\Omega}(\nabla v_{\tau}^{k}\cdot\nabla\zeta+v_{\tau}^{k}\zeta-u_{\tau}^{k}\zeta)\,dx=0\quad\mathrm{for\ all}\ \zeta\in H^{1}(\Omega).$
In additon, if $\Delta v_{\tau}^{k}\in L^{2}(\Omega)$ then it holds that $\nabla v_{\tau}^{k}\cdot\boldsymbol{n}=0$ on $\partial\Omega$.
Proof.
Let $\zeta\in H^{1}(\Omega)$ and $a>0$. Note that $v_{\tau}^{k}+a\zeta\in H^{1}(\Omega)$. By (3.11) with $(\tilde{u},\tilde{v})=(u_{\tau}^{k},v_{\tau}^{k}+a\zeta)$, it follows
$\frac{1}{2\tau}\|v_{\tau}^{k}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}+E(u_{\tau}^{k},v_{\tau}^{k})\leq\frac{1}{2\tau}\|v_{\tau}^{k}+a\zeta-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}+E(u_{\tau}^{k},v_{\tau}^{k}+a\zeta),$
then
$0$ $\leq\frac{1}{2\tau}\int_{\Omega}(|v_{\tau}^{k}+a\zeta-v_{\tau}^{k-1}|^{2}-|v_{\tau}^{k}-v_{\tau}^{k-1}|^{2})\,dx$
$\quad+\frac{1}{2}\int_{\Omega}(|\nabla v_{\tau}^{k}+a\nabla\zeta|^{2}-|\nabla v_{\tau}^{k}|^{2})\,dx+\frac{1}{2}\int_{\Omega}(|v_{\tau}^{k}+a\zeta|^{2}-|v_{\tau}^{k}|^{2})\,dx-a\int_{\Omega}u_{\tau}^{k}\zeta\,dx.$
Dividing by $a>0$ and letting $a\to 0$, by simple calculations, we have
$0\leq\int_{\Omega}\frac{v_{\tau}^{k}-v_{\tau}^{k-1}}{\tau}\zeta\,dx+\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla\zeta\,dx+\int_{\Omega}v_{\tau}^{k}\zeta\,dx-\int_{\Omega}u_{\tau}^{k}\zeta\,dx.$
Replacing $\zeta$ by $\zeta$, we obtain the opposite inequality.
Assume that $\Delta v_{\tau}^{k}\in L^{2}(\Omega)$. Letting $\psi\in C_{c}^{\infty}(\Omega)$ be arbitary, we have
$\int_{\Omega}\left(\frac{v_{\tau}^{k}-v_{\tau}^{k-1}}{\tau}-\Delta v_{\tau}^{k}+v_{\tau}^{k}-u_{\tau}^{k}\right)\psi\,dx=0.$
Hence it follows
$\frac{v_{\tau}^{k}-v_{\tau}^{k-1}}{\tau}-\Delta v_{\tau}^{k}+v_{\tau}^{k}-u_{\tau}^{k}=0\quad\mathrm{a.e.\ in}\ \Omega.$
Then, for all $\zeta\in H^{1}(\Omega)$, we conclude that
$0$ $=\int_{\Omega}\left(\frac{v_{\tau}^{k}-v_{\tau}^{k-1}}{\tau}-\Delta v_{\tau}^{k}+v_{\tau}^{k}-u_{\tau}^{k}\right)\zeta\,dx+\int_{\partial\Omega}\nabla v_{\tau}^{k}\cdot\boldsymbol{n}\zeta\,dS$
$=\int_{\partial\Omega}\nabla v_{\tau}^{k}\cdot\boldsymbol{n}\zeta\,dS,$
then $\nabla v_{\tau}^{k}\cdot\boldsymbol{n}=0$ on $\partial\Omega$. ∎
Corollary 4.10.
Let $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be a minimizer of (3.1) with $\Delta v_{\tau}^{k}\in L^{2}(\Omega)$ and $u_{\tau}^{k}\in L^{2}(\Omega)$. Then $v_{\tau}^{k}\in H^{2}(\Omega)$ and there exists a constant $C>0$ such that
$\|v_{\tau}^{k}\|_{H^{2}(\Omega)}^{2}\leq C(\|\Delta v_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}+\|v_{\tau}^{k}\|_{H^{1}(\Omega)}^{2}).$
Proof.
By Lemma 4.6, Lemma 4.9 and the elliptic regularity theorem, we can complete the proof immediately. ∎
5 Uniform estimates and convergences
Definition 5.1.
We define the following piecewise constant functions:
$\begin{cases}u_{\tau}(t)\coloneq u_{\tau}^{k}\quad\mathrm{if}\ t\in((k-1)\tau,k\tau]\quad\mathrm{for}\ k\in\mathbb{N},\ u_{\tau}(0)\coloneq u_{0},\\ v_{\tau}(t)\coloneq v_{\tau}^{k}\quad\mathrm{if}\ t\in((k-1)\tau,k\tau]\quad\mathrm{for}\ k\in\mathbb{N},\ v_{\tau}(0)\coloneq v_{0}.\end{cases}$
Notice that since the minimizers $(u_{\tau}^{k},v_{\tau}^{k})$ are nonnegative functions (Remark 3.5), the above functions are also nonnegative.
First, we establish the uniform estimates including time variables.
Lemma 5.2.
Let $T>0$, $1+\alpha-2/d\leq p\leq 1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then there exist positive constants $C_{1},C_{2}$ and $C_{3}$ depending on $\alpha,p,d,\chi,u_{0}$ and $v_{0}$ such that the following uniform estimates hold:
$\sup_{0\leq t\leq T}\left(\|u_{\tau}(t)\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\|v_{\tau}(t)\|_{H^{1}(\Omega)}^{2}\right)\leq C_{1},$ (5.1)
$\int_{0}^{T}\left(\|\nabla(u_{\tau}(t))^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}(t)-v_{\tau}(t)+u_{\tau}(t)\|_{L^{2}(\Omega)}^{2}\right)\,dt\leq C_{2}(1+T),$ (5.2)
$\int_{0}^{T}\|v_{\tau}(t)\|_{H^{2}(\Omega)}^{2}\,dt\leq C_{3}(1+T).$ (5.3)
Proof.
To simplify, we set $T=N\tau$ for $N\in\mathbb{N}$. From (3.12), summing up $i=1$ to $i=k$ for any $k\in\mathbb{N}$, we have
$\sum_{i=1}^{k}E(u_{\tau}^{i},v_{\tau}^{i})$ $\leq\sum_{i=1}^{k}E(u_{\tau}^{i-1},v_{\tau}^{i-1}),$
so that
$E(u_{\tau}^{k},v_{\tau}^{k})$ $\leq E(u_{0},v_{0}),$
then
$\frac{p}{\chi(p-\alpha)(p+1-\alpha)}\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}-\|u_{\tau}^{k}v_{\tau}^{k}\|_{L^{1}(\Omega)}+\frac{1}{2}\|v_{\tau}^{k}\|_{H^{1}(\Omega)}^{2}\leq E(u_{0},v_{0}).$
By using the inequality (3.3) or (3.1), it follows
$\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\|v_{\tau}^{k}\|_{H^{1}(\Omega)}^{2}\leq C_{1}\quad\mathrm{for}\ k\in\mathbb{N}$
for some constant $C_{1}=C_{1}(\alpha,p,d,\chi,u_{0},v_{0})$, which gives
$\sup_{0\leq t\leq T}\left(\|u_{\tau}(t)\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}+\|v_{\tau}(t)\|_{H^{1}(\Omega)}^{2}\right)\leq C_{1}.$
Combining the above uniform estimate with (4.6), we have
$\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}^{k}-v_{\tau}^{k}+u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}$
$\leq\frac{\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k-1})-\boldsymbol{U}_{\varepsilon}(u_{\tau}^{k})}{\tau}+\frac{\|v_{\tau}^{k-1}\|_{H^{1}(\Omega)}^{2}-\|v_{\tau}^{k}\|_{H^{1}(\Omega)}^{2}}{\tau}+C_{1}+C_{0}(1+C_{1}+C_{1}^{\frac{1}{p-\alpha}}).$
Hence it follows by integrating over $(0,T)$ that
$\int_{0}^{T}\left(\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(u_{\tau}(t))^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}(t)-v_{\tau}(t)+u_{\tau}(t)\|_{L^{2}(\Omega)}^{2}\right)\,dt$
$=\sum_{k=1}^{N}\int_{(k-1)\tau}^{k\tau}\left(\frac{4p}{\chi(p+1-\alpha)^{2}}\|\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}^{k}-v_{\tau}^{k}+u_{\tau}^{k}\|_{L^{2}(\Omega)}^{2}\right)\,dt$
$\leq\boldsymbol{U}_{\varepsilon}(u_{0})-\boldsymbol{U}_{\varepsilon}(u_{\tau}^{N})+\|v_{0}\|_{H^{1}(\Omega)}^{2}-\|v_{\tau}^{N}\|_{H^{1}(\Omega)}^{2}+\tilde{C}T$
$\leq\frac{1}{1-\alpha}\|u_{0}\|_{L^{2-\alpha}(\Omega)}^{2-\alpha}+\|v_{0}\|_{H^{1}(\Omega)}^{2}+\tilde{C}T,$
where $\tilde{C}=C_{1}+C_{0}(1+C_{1}+C_{1}^{\frac{1}{p-\alpha}})$ and we used Lemma 2.9:
$\boldsymbol{U}_{\varepsilon}(u_{0})\leq\frac{1}{1-\alpha}\|u_{0}\|_{L^{2-\alpha}(\Omega)}^{2-\alpha}\ \mathrm{and}\ \boldsymbol{U}_{\varepsilon}\geq 0.$
Thus there exists a constant $C_{2}=C_{2}(\alpha,p,d,\chi,u_{0},v_{0})>0$ such that
$\int_{0}^{T}\left(\|\nabla(u_{\tau}(t))^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}+\|\Delta v_{\tau}(t)-v_{\tau}(t)+u_{\tau}(t)\|_{L^{2}(\Omega)}^{2}\right)\,dt\leq C_{2}(1+T).$
Observe that if $p=1+\alpha$ then $p+1-\alpha=2$, that is, $\sup_{0\leq t\leq T}\|u_{\tau}(t)\|_{L^{2}(\Omega)}^{2}\leq C_{1}$. By Corollary 4.10, (5.1), (5.2) and (4.7), we can get the estimate (5.3) for some constant $C_{3}$. ∎
The following lemma is about estimates like the equi-continuity to use the refined Ascoli–Arzelà theorem ([2, Proposition 3.3.1]). Note that the weighted Wasserstein distance depends on $\varepsilon$.
Lemma 5.3.
Let $T>0$, $p\geq 1+\alpha-2/d$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then there exists $C_{4}=C_{4}(\alpha,p,d,\chi,u_{0},v_{0})>0$ satisfying for all $(t,s)\in[0,T]^{2}$ and $\tau\in(0,1)$ it holds
$W_{m_{\varepsilon}}(u_{\tau}(t),u_{\tau}(s))\leq C_{4}(\sqrt{|t-s|}+\sqrt{\tau}),$
$\|v_{\tau}(t)-v_{\tau}(s)\|_{L^{2}(\Omega)}\leq C_{4}(\sqrt{|t-s|}+\sqrt{\tau}).$
Proof.
We only prove the first inequality because the other inequality can be shown by the same argument. Let $0\leq s<t\leq T$ and define
$N\coloneq\bigg\lceil\frac{t}{\tau}\bigg\rceil,\ P\coloneq\bigg\lceil\frac{s}{\tau}\bigg\rceil,$
where $\lceil x\rceil$ denotes the superior integer part of the real number $x$. From (3.11) with $\tilde{u}=u_{\tau}^{k-1}$ and $\tilde{v}=v_{\tau}^{k-1}$, we have
$W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}+\chi\|v_{\tau}^{k}-v_{\tau}^{k-1}\|_{L^{2}(\Omega)}^{2}\leq 2\tau\chi(E(u_{\tau}^{k-1},v_{\tau}^{k-1})-E(u_{\tau}^{k},v_{\tau}^{k})),$
then
$\sum_{k=1}^{N}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}\leq 2\tau\chi(E(u_{0},v_{0})-E(u_{\tau}^{N},v_{\tau}^{N})).$
Because the functional $E$ is bounded below in $X$ (see Lemma 3.1), we see
$\sum_{k=1}^{N}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}\leq 2\tau\chi\left(E(u_{0},v_{0})-\inf_{(u,v)\in X}E(u,v)\right).$
Since $t\in((N-1)\tau,N\tau]$ and $s\in((P-1)\tau,P\tau]$ by the definition of $N$ and $P$, it follows
$W_{m_{\varepsilon}}(u_{\tau}(t),u_{\tau}(s))$ $=W_{m_{\varepsilon}}(u_{\tau}^{N},u_{\tau}^{P})\leq\sum_{k=P+1}^{N}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})$
$\leq\sqrt{N-P}\sqrt{\sum_{k=P+1}^{N}W_{m_{\varepsilon}}(u_{\tau}^{k},u_{\tau}^{k-1})^{2}}$
$\leq\sqrt{N-P}\sqrt{2\tau\chi\left(E(u_{0},v_{0})-\inf_{(u,v)\in X}E(u,v)\right)}$
$\leq\sqrt{2\chi}\sqrt{t-s+\tau}\left(E(u_{0},v_{0})-\inf_{(u,v)\in X}E(u,v)\right)^{\frac{1}{2}}$
$\leq C_{4}(\sqrt{|t-s|}+\sqrt{\tau}),$
where $C_{4}=C_{4}(\alpha,p,d,\chi,u_{0},v_{0})$ is a constant, and in the second ineqality, we used
$(x_{1}+\cdots+x_{n})^{2}\leq n(x_{1}^{2}+\cdots+x_{n}^{2})\quad\mathrm{for}\ x_{i}\geq 0,\ i=1,\cdots,n.$
The proof is completed. ∎
From the above lemmas, we obtain the convergences with respect to $\tau$.
Lemma 5.4.
Let $p=1+\alpha$ and $T>0$. There exist a subsequence $\{(u_{\tau_{n}},v_{\tau_{n}})\}_{n}$ with $\tau_{n}\to 0$ as $n\to\infty$ and a pair of functions $(u_{\varepsilon},v_{\varepsilon})\in X$ such that
$u_{\tau_{n}}\rightharpoonup u_{\varepsilon}\quad\mathrm{weakly\ in}\ L^{2}((0,T);H^{1}(\Omega))\ \mathrm{as}\ n\to\infty,$
$u_{\tau_{n}}(t)\rightharpoonup u_{\varepsilon}(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T],$
$v_{\tau_{n}}\rightharpoonup v_{\varepsilon}\quad\mathrm{weakly\ in}\ L^{2}((0,T);H^{2}(\Omega))\ \mathrm{as}\ n\to\infty,$
$v_{\tau_{n}}(t)\rightharpoonup v_{\varepsilon}\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T].$
In particular, $v_{\varepsilon}\in C^{\frac{1}{2}}([0,T];L^{2}(\Omega))$.
Proof.
By Lemma 5.2, $\{u_{\tau}\}_{\tau>0}$ is bounded in $L^{2}((0,T);H^{1}(\Omega))$, then there exist a subsequence $\{u_{\tau_{n}}\}$ and a function $u_{\varepsilon}\in L^{2}((0,T);H^{1}(\Omega))$ such that $u_{\tau_{n}}$ weakly converges to $u_{\varepsilon}$ in $L^{2}((0,T);H^{1}(\Omega))$. In addition, by Lemma 5.3 and the refined Ascoli–Arzelà theorem ([2, Proposition 3.3.1]), there exist a subsequence (not relabeled) and $\tilde{u}_{\varepsilon}:[0,T]\to\mathcal{P}(\Omega)$ such that
$u_{\tau_{n}}(t)\rightharpoonup\tilde{u}_{\varepsilon}(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T].$
Due to the uniqueness of limit, we have $u_{\varepsilon}=\tilde{u}_{\varepsilon}$ a.e. in $\Omega\times[0,T]$. Similarly, by Lemma 5.2, $\{v_{\tau}\}_{\tau>0}$ is bounded in $L^{2}((0,T);H^{2}(\Omega))$ and by Lemma 5.3 and the refined Ascoli–Arzelà theorem, we have
$v_{\tau_{n}}\rightharpoonup v_{\varepsilon}\quad\mathrm{weakly\ in}\ L^{2}((0,T);H^{2}(\Omega))\ \mathrm{as}\ n\to\infty,$
$v_{\tau_{n}}(t)\rightharpoonup v_{\varepsilon}\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T],$
$v_{\varepsilon}\in C^{\frac{1}{2}}([0,T];L^{2}(\Omega)).$
The proof is completed. ∎
In previous lemma, we derived the weak convergences with respect to $\tau$, hence we next obtain the strong convergence for $\tau$.
Lemma 5.5.
Let $p=1+\alpha$ and $T>0$. Then for the sequence $\{u_{\tau_{n}}\}_{n}$ in Lemma 5.4, it holds
$u_{\tau_{n}}\to u_{\varepsilon}\quad\mathrm{strongly\ in}\ L^{2}(\Omega\times(0,T))\ \mathrm{as}\ n\to\infty.$
Proof.
By the Rellich–Kondrachov theorem, $H^{1}(\Omega)$ is compactly embedded in $L^{2}(\Omega)$. Hence $L^{2}(\Omega)$ is compactly embedded in $(H^{1}(\Omega))^{*}$, where $(H^{1}(\Omega))^{*}$ is the dual space of $H^{1}(\Omega)$. By Lemma 5.2, $\|u_{\tau_{n}}(t)\|_{L^{2}(\Omega)}$ is bounded with respect to $\tau_{n}$ for all $t\in[0,T]$, thus there exist a subsequence (not relabeled) and $w_{t}\in(H^{1}(\Omega))^{*}$ such that $u_{\tau_{n}}(t)$ converges to $w_{t}$ strongly in $(H^{1}(\Omega))^{*}$. Now thanks to Lemma 5.4, we know that $u_{\tau_{n}}(t)$ converges to $u_{\varepsilon}(t)$ weakly in $L^{2}(\Omega)$. Due to the uniqueness of limit, we have $w_{t}=u_{\varepsilon}(t)$ a.e. in $\Omega$. Moreover, $\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{(H^{1}(\Omega))^{*}}^{2}\leq\sup_{0\leq t\leq T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{L^{2}(\Omega)}^{2}$ is bounded by Lemma 5.2 and Lemma 5.4. Hence we infer from Lebesgue’s dominated convergence theorem that
$\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{(H^{1}(\Omega))^{*}}^{2}\,dt\to 0\quad\mathrm{as}\ n\to\infty.$
On the other hand, using the interpolation inequality, we have
$\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{L^{2}(\Omega)}^{2}\,dt$
$\leq\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{H^{1}(\Omega)}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{(H^{1}(\Omega))^{*}}\,dt$
$\leq\left(\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{H^{1}(\Omega)}^{2}\,dt\right)^{\frac{1}{2}}\left(\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{(H^{1}(\Omega))^{*}}^{2}\,dt\right)^{\frac{1}{2}}.$
Combining the above result with Lemma 5.2, we conclude that $u_{\tau_{n}}$ converges to $u_{\varepsilon}$ strongly in $L^{2}(\Omega\times(0,T))$ as $n\to\infty$. ∎
6 Proof of Theorem 1.1
First, we establish weak formulations to (1.1)ε.
Lemma 6.1.
Let $p=1+\alpha$. Then $(u_{\varepsilon},v_{\varepsilon})$ in Lemma 5.4 satisfies the following weak formulation: for all $T>0$ and $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$, it holds
$\int_{\Omega}(u_{0}-u_{\varepsilon}(T))\varphi\,dx=\int_{0}^{T}\int_{\Omega}\nabla\varphi\cdot[m_{\varepsilon}(u_{\varepsilon})(p\nabla u_{\varepsilon}-\chi\nabla v_{\varepsilon})]\,dx\,dt.$ (6.1)
Proof.
Let $T>0$ and $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$. Let $\{\tau_{n}\}\subset(0,1)$ be a subsequence of $\{\tau\}$ which is obtained in Lemma 5.4 and Lemma 5.5, and set $\delta_{n}\coloneq\tau_{n}^{\frac{1}{2}}$. To simplify, we assume that $T=N\tau_{n}$ for some $N\in\mathbb{N}$. By Lemma 4.8, we have
$\frac{\boldsymbol{V}_{\delta_{n}}(u_{\tau_{n}}(T))-\boldsymbol{V}_{\delta_{n}}(u_{\tau_{n}}(0))}{\chi}$
$=\frac{\boldsymbol{V}_{\delta_{n}}(u_{\tau_{n}}^{N})-\boldsymbol{V}_{\delta_{n}}(u_{0})}{\chi}=\sum_{k=1}^{N}\frac{\boldsymbol{V}_{\delta_{n}}(u_{\tau_{n}}^{k})-\boldsymbol{V}_{\delta_{n}}(u_{\tau_{n}}^{k-1})}{\chi}$
$\leq-\sum_{k=1}^{N}\tau_{n}\left[\frac{p}{\chi}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}}^{k})\nabla u_{\tau_{n}}^{k}\cdot\nabla\varphi\,dx-\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}}^{k})\nabla v_{\tau_{n}}^{k}\cdot\nabla\varphi\,dx\right]$
$\quad-\tau_{n}\lambda_{\delta_{n}}\sum_{k=1}^{N}(E(u_{\tau_{n}}^{k-1},v_{\tau_{n}}^{k-1})-E(u_{\tau_{n}}^{k},v_{\tau_{n}}^{k}))+\delta_{n}\sum_{k=1}^{N}\tau_{n}\int_{\Omega}\nabla v_{\tau_{n}}^{k}\cdot\nabla u_{\tau_{n}}^{k}\,dx$
$=-\int_{0}^{T}\int_{\Omega}\frac{p}{\chi}m_{\varepsilon}(u_{\tau_{n}}(t))\nabla u_{\tau_{n}}(t)\cdot\nabla\varphi\,dx\,dt+\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}}(t))\nabla v_{\tau_{n}}(t)\cdot\nabla\varphi\,dx\,dt$
$\quad-\tau_{n}\lambda_{\delta_{n}}\left(E(u_{0},v_{0})-E(u_{\tau_{n}}^{N},v_{\tau_{n}}^{N})\right)+\delta_{n}\int_{0}^{T}\int_{\Omega}\nabla v_{\tau_{n}}(t)\cdot\nabla u_{\tau_{n}}(t)\,dx\,dt.$
Here, by the definition of $\lambda_{\delta_{n}}$, we see
$\tau_{n}|\lambda_{\delta_{n}}|$ $\leq\tau_{n}(\delta_{n}^{-1}+1)C(\varphi,\alpha,\varepsilon)\leq 2\tau_{n}^{\frac{1}{2}}C(\varphi,\alpha,\varepsilon),$
and by Hölder’s inequality and Lemma 5.2, it follows
$\int_{0}^{T}\int_{\Omega}|\nabla v_{\tau_{n}}(t)||\nabla u_{\tau_{n}}(t)|\,dx\,dt$ $\leq\left(\int_{0}^{T}\|\nabla v_{\tau_{n}}(t)\|_{L^{2}(\Omega)}^{2}\,dt\right)^{\frac{1}{2}}\left(\int_{0}^{T}\|\nabla u_{\tau_{n}}(t)\|_{L^{2}(\Omega)}^{2}\,dt\right)^{\frac{1}{2}}$
$\leq(C_{1}C_{2})^{\frac{1}{2}}(1+T).$
Hence taking account of the definition of $\boldsymbol{V}_{\delta_{n}}$ and the boundedness from below of $E$ in $X$ (Lemma 3.1), we obtain
$\int_{\Omega}(u_{\tau_{n}}(x,T)-u_{0}(x))\varphi(x)\,dx$
$\leq-\int_{0}^{T}\int_{\Omega}\frac{p}{\chi}m_{\varepsilon}(u_{\tau_{n}}(x,t))\nabla u_{\tau_{n}}(x,t)\cdot\nabla\varphi(x)\,dx\,dt$
$\quad+\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}}(x,t))\nabla v_{\tau_{n}}(x,t)\cdot\nabla\varphi(x)\,dx\,dt$
$\quad+\tau_{n}^{\frac{1}{2}}\left[2C(\varphi,\alpha,\varepsilon)\left(E(u_{0},v_{0})-\inf_{(\tilde{u},\tilde{v})\in X}E(\tilde{u},\tilde{v})\right)+(C_{2}C_{3})^{\frac{1}{2}}(1+T)\right]$
$\quad+\tau_{n}^{\frac{1}{2}}\left[\boldsymbol{U}_{\varepsilon}(u_{0})-\boldsymbol{U}_{\varepsilon}(u_{\tau_{n}}(T))\right].$ (6.2)
By Lemma 2.9, we have
$\boldsymbol{U}_{\varepsilon}(u_{0})\leq\frac{1}{1-\alpha}\|u_{0}\|_{L^{2-\alpha}(\Omega)}^{2-\alpha}\ \mbox{and}\ \boldsymbol{U}_{\varepsilon}(u_{\tau_{n}}(T))\geq 0.$
Thanks to Lemma 5.4, it is easy to check that
$\int_{\Omega}u_{\tau_{n}}(x,T)\varphi(x)\,dx\to\int_{\Omega}u_{\varepsilon}(x,T)\varphi(x)\,dx\quad\mathrm{as}\ n\to\infty.$
We will show
$\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla u_{\tau_{n}}\cdot\nabla\varphi\,dx\,dt\to\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla u_{\varepsilon}\cdot\nabla\varphi\,dx\,dt\quad\mathrm{as}\ n\to\infty$
and
$\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla v_{\tau_{n}}\cdot\nabla\varphi\,dx\,dt\to\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla v_{\varepsilon}\cdot\nabla\varphi\,dx\,dt\quad\mathrm{as}\ n\to\infty.$
First, we have
$\left|\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla u_{\tau_{n}}\cdot\nabla\varphi\,dx\,dt-\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla u_{\varepsilon}\cdot\nabla\varphi\,dx\,dt\right|$
$\leq\int_{0}^{T}\int_{\Omega}|m_{\varepsilon}(u_{\tau_{n}})-m_{\varepsilon}(u_{\varepsilon})||\nabla u_{\tau_{n}}||\nabla\varphi|\,dx\,dt+\left|\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})(\nabla u_{\tau_{n}}-\nabla u_{\varepsilon})\cdot\nabla\varphi\,dx\,dt\right|$
$\eqcolon\mathrm{I}_{1}+\mathrm{I}_{2}.$
Observe that $m_{\varepsilon}(r)=(r+\varepsilon)^{\alpha}$, we infer from (2.11), Hölder’s inequality and (5.2) that
$\mathrm{I}_{1}$ $\leq\int_{0}^{T}\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{\alpha}|\nabla u_{\tau_{n}}||\nabla\varphi|\,dx\,dt$
$\leq\|\nabla\varphi\|_{L^{\infty}(\Omega)}\|\nabla u_{\tau_{n}}\|_{L^{2}(\Omega\times(0,T))}\int_{0}^{T}\left(\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{2\alpha}\,dx\right)^{\frac{1}{2}}\,dt$
$\leq\|\nabla\varphi\|_{L^{\infty}(\Omega)}\sup_{n}\left\{\|\nabla u_{\tau_{n}}\|_{L^{2}(\Omega\times(0,T))}\right\}|\Omega|^{\frac{1-\alpha}{2}}\int_{0}^{T}\|u_{\tau_{n}}(t)-u_{\varepsilon}(t)\|_{L^{2}(\Omega)}^{\alpha}\,dt$
$\leq\|\nabla\varphi\|_{L^{\infty}(\Omega)}C_{2}(1+T)|\Omega|^{\frac{1-\alpha}{2}}T^{\frac{2-\alpha}{2}}\|u_{\tau_{n}}-u_{\varepsilon}\|_{L^{2}(\Omega\times(0,T))}^{\alpha}.$
By Lemma 5.5, we have $\mathrm{I}_{1}\to 0$ as $n\to\infty$. Further, since $\nabla u_{\tau_{n}}\rightharpoonup\nabla u_{\varepsilon}$ weakly in $L^{2}(\Omega\times(0,T))$ and $m_{\varepsilon}(u_{\varepsilon})\nabla\varphi\in L^{2}(\Omega\times(0,T))$, we obtain $\mathrm{I}_{2}\to 0$ as $n\to\infty$.
Next, we have
$\left|\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla v_{\tau_{n}}\cdot\nabla\varphi\,dx\,dt-\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla v_{\varepsilon}\cdot\nabla\varphi\,dx\,dt\right|$
$\leq\int_{0}^{T}\int_{\Omega}|m_{\varepsilon}(u_{\tau_{n}})-m_{\varepsilon}(u_{\varepsilon})||\nabla v_{\tau_{n}}||\nabla\varphi|\,dx\,dt+\left|\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})(\nabla v_{\tau_{n}}-\nabla v_{\varepsilon})\nabla\varphi\,dx\,dt\right|.$
Since $\sup_{0\leq t\leq T}\|\nabla v_{\tau_{n}}(t)\|_{L^{2}(\Omega)}$ is bounded and $\nabla v_{\tau_{n}}$ converges to $\nabla v_{\varepsilon}$ weakly in $L^{2}(\Omega\times(0,T))$, as in the above argument, we conclude that the right hand side converges to 0 as $n\to\infty$.
Hence by letting $n\to\infty$ in (6), it follows
$\frac{1}{\chi}\int_{\Omega}(u_{\varepsilon}(x,T)-u_{0}(x))\varphi(x)\,dx$
$\leq-\int_{0}^{T}\int_{\Omega}\frac{p}{\chi}m_{\varepsilon}(u_{\varepsilon}(x,t))\nabla u_{\varepsilon}(x,t)\cdot\nabla\varphi(x)\,dx\,dt+\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon}(x,t))\nabla v_{\varepsilon}(x,t)\cdot\nabla\varphi(x)\,dx\,dt.$
Replacing $\varphi$ with $\varphi$, we have
$\int_{\Omega}(u_{0}(x)-u_{\varepsilon}(x,T))\varphi(x)\,dx=\int_{0}^{T}\int_{\Omega}\nabla\varphi(x)\cdot m_{\varepsilon}(u_{\varepsilon})(p\nabla u_{\varepsilon}(x,t)-\chi\nabla v_{\varepsilon}(x,t))\,dx\,dt.$
The proof is completed. ∎
Lemma 6.2.
Let $T>0$. Then $(u_{\varepsilon},v_{\varepsilon})$ satisfies the following weak formulation: for all $\zeta\in H^{1}(\Omega)$, it holds
$\int_{0}^{T}\int_{\Omega}[\nabla v_{\varepsilon}\cdot\nabla\zeta+v_{\varepsilon}\zeta-u_{\varepsilon}\zeta]\,dx\,dt=\int_{\Omega}(v_{0}-v_{\varepsilon}(T))\zeta\,dx.$
Proof.
Let $T>0$, $\zeta\in H^{1}(\Omega)$ and $\{\tau_{n}\}$ be a subsequence in Lemma 5.4 and Lemma 5.5. To simplify, we assume $T=N\tau_{n}$. By Lemma 4.9, we have
$\sum_{k=1}^{N}\tau_{n}\int_{\Omega}(\nabla v_{\tau_{n}}^{k}\cdot\nabla\zeta+v_{\tau_{n}}^{k}\zeta-u_{\tau_{n}}^{k}\zeta)\,dx=\sum_{k=1}^{N}\int_{\Omega}(v_{\tau_{n}}^{k-1}-v_{\tau_{n}}^{k})\zeta\,dx$
then
$\int_{0}^{T}\int_{\Omega}(\nabla v_{\tau_{n}}(t)\cdot\nabla\zeta+v_{\tau_{n}}(t)\zeta-u_{\tau_{n}}(t)\zeta)\,dx\,dt=\int_{\Omega}(v_{0}-v_{\tau_{n}}^{N})\zeta\,dx.$
Hence letting $n\to\infty$, we infer from Lemma 5.4 that
$\int_{0}^{T}\int_{\Omega}(\nabla v_{\varepsilon}(t)\cdot\nabla\zeta+v_{\varepsilon}(t)\zeta-u_{\varepsilon}(t)\zeta)\,dx\,dt=\int_{\Omega}(v_{0}-v_{\varepsilon}(T))\zeta\,dx.$
The proof is completed. ∎
Next lemma is about the weak compactness of $\{u_{\varepsilon}(t)\}_{\varepsilon}$ for each $t\in[0,T]$. As in the Lemma 5.4, if we have the equi-continuity with respect to $W_{m}$, where $m(r)=r^{\alpha}$, then we can easily get the conclusion adapting the refined Ascoli–Arzelà theorem ([2, Proposition 3.3.1]). However, we only have the equi-continuity with respect to $W_{m_{\varepsilon}}$ depending on $\varepsilon$ (Lemma 5.3). To avoid this problem, we use not only the equi-continuity but also the lower semicontinuity of the weighted Wasserstein distance (see Lemma 2.6).
Lemma 6.3.
Let $p=1+\alpha$ and $T>0$. There exist a subsequence $\{u_{\varepsilon_{n}}\}_{n}$ with $\varepsilon_{n}\to 0$ as $n\to\infty$ and $u:[0,T]\to\mathcal{P}(\Omega)$ such that
$u_{\varepsilon_{n}}(t)\rightharpoonup u(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T].$
In particular, $W_{m}(u(t),u(s))\leq C_{4}\sqrt{|t-s|}\ \mathrm{for}\ t,s\in[0,T]$, where $m(r)=r^{\alpha}$.
Proof.
First, $(\mathcal{M}_{loc}^{+}(\mathbb{R}^{d}),W_{m})$ is complete ([8, Theorem 5.7]) and $m_{\varepsilon}(r)$ is decreasing with respect to $\varepsilon$ and pointwise converging to $m(r)$ as $\varepsilon\to 0$. Set $S\coloneq\{f\in L^{2}\cap\mathcal{P}(\Omega);\|f\|_{L^{2}(\Omega)}^{2}\leq C_{1}\}$ where $C_{1}$ is the constant in Lemma 5.2. Then $S$ is sequentially compact with respect to the weak topology of $L^{1}\cap L^{2}(\Omega)$. Indeed, let $\{f_{n}\}\subset S$, we can easily see that $\{f_{n}\}$ is bounded in $L^{1}\cap L^{2}(\Omega)$ and equi-integrable. Hence, taking a subsequence (not relabeled), there exists a function $f\in L^{1}\cap L^{2}(\Omega)$ such that
$f_{n}\rightharpoonup f\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ n\to\infty.$
Note that if $f_{n}$ converges to $f$ weakly in $L^{1}\cap L^{2}(\Omega)$ then $f_{n}$ converges to $f$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$. Since $f_{n}$ weakly converges to $f$ in $L^{2}(\Omega)$ as $n\to\infty$, we have $\|f\|_{L^{2}(\Omega)}^{2}\leq C_{1}$, then $f\in S$.
Since $u_{\varepsilon}(t)\in L^{2}\cap\mathcal{P}(\Omega)$ and $\|u_{\varepsilon}(t)\|_{L^{2}(\Omega)}^{2}\leq C_{1}$ for all $t\in[0,T]$ by Lemma 5.2 and Lemma 5.4, that is, $\{u_{\varepsilon}(t)\}_{\varepsilon}\subset S$ for all $t\in[0,T]$, using the diagonal argument, there exist a subsequence $\{u_{\varepsilon_{n}}\}_{n}$ and $u:\mathbb{Q}\cap[0,T]\to S$ such that
$u_{\varepsilon_{n}}(t)\rightharpoonup u(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in\mathbb{Q}\cap[0,T].$
By Lemma 2.6 and Lemma 5.3, we have
$W_{m}(u(t),u(s))$ $\leq\liminf_{\varepsilon_{n}\to 0}W_{\varepsilon_{n}}(u_{\varepsilon_{n}}(t),u_{\varepsilon_{n}}(s))\leq\limsup_{\varepsilon_{n}\to 0}W_{\varepsilon_{n}}(u_{\varepsilon_{n}}(t),u_{\varepsilon_{n}}(s))$
$\leq\limsup_{\varepsilon_{n}\to 0}\liminf_{\tau\to 0}W_{\varepsilon_{n}}(u_{\tau}(t),u_{\tau}(s))$
$\leq\limsup_{\varepsilon_{n}\to 0}\limsup_{\tau\to 0}W_{\varepsilon_{n}}(u_{\tau}(t),u_{\tau}(s))$
$\leq C_{4}\sqrt{|t-s|}.$
We will show that $(S,W_{m})$ is complete. Let $\{f_{n}\}\subset S$ be a Cauchy sequence, since $\{f_{n}\}\subset L_{loc}^{1}(\mathbb{R}^{d})\subset\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ and $(\mathcal{M}_{loc}^{+}(\mathbb{R}^{d}),W_{m})$ is complete, there exists a Radon measure $f\in\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ such that $f_{n}\to f$ in $W_{m}$. In particular, $f_{n}\rightharpoonup f$ weakly* in $\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ (Proposition 2.4). Note that since $\{f_{n}\}\subset L_{loc}^{1}(\mathbb{R}^{d})$ and $\{f_{n}\}$ is bounded in $L^{2}(\Omega)$, we can identify the measure $f\in\mathcal{M}_{loc}^{+}(\mathbb{R}^{d})$ with the density function $f\in L^{1}_{loc}(\mathbb{R}^{d})$.
On the other hand, since $\{f_{n}\}\subset S$, there exist a subsequence $\{f_{n_{k}}\}_{k}$ and $g\in S$ such that $f_{n_{k}}\rightharpoonup g$ weakly in $L^{1}\cap L^{2}(\Omega)$. For all $\zeta\in C_{c}^{\infty}(\Omega)$, we have
$\int_{\Omega}f_{n_{k}}\zeta\,dx\to\int_{\Omega}f\zeta\,dx\quad\mathrm{as}\ k\to\infty$
and
$\int_{\Omega}f_{n_{k}}\zeta\,dx\to\int_{\Omega}g\zeta\,dx\quad\mathrm{as}\ k\to\infty.$
Hence we obtain $f=g$ a.e. in $\Omega$. Thus it holds that $f\in S$ and
$W_{m}(f_{n},f)\leq W_{m}(f_{n},f_{n_{k}})+W_{m}(f_{n_{k}},f)\to 0\quad\mathrm{as}\ n,k\to\infty.$
Let $t\in[0,T]$, then there exists $\{t_{k}\}\subset\mathbb{Q}\cap[0,T]$ such that $t_{k}\to t$ as $k\to\infty$. Since $W_{m}(u(t_{k}),u(t_{l}))\leq C_{4}\sqrt{|t_{k}-t_{l}|}\to 0$ as $k,l\to\infty$ and $\{u(t_{k})\}\subset S$, we can uniquely define
$u(t)\coloneq\lim_{k\to\infty}u(t_{k})\quad\mathrm{in}\ (S,W_{m}).$
Hence we obtain $u:[0,T]\to S\subset\mathcal{P}(\Omega)$.
Finally, we show $u_{\varepsilon_{n}}(t)\rightharpoonup u(t)$ weakly in $L^{1}\cap L^{2}(\Omega)$ for $t\in[0,T]$. It is sufficient to prove that all subsequences of $\{u_{\varepsilon_{n}}(t)\}$ have a subsequence converging to $u(t)$ weakly in $L^{1}\cap L^{2}(\Omega)$. Fix $t\in[0,T]$ and let $\{u_{\varepsilon_{n}^{\prime}}(t)\}\subset\{u_{\varepsilon_{n}}(t)\}$. Since $S$ is sequentially compact, taking a subsequence (not relabeled), we have
$u_{\varepsilon_{n}^{\prime}}(t)\rightharpoonup\tilde{u}\quad\mathrm{weakly\ in}\ L^{1}\cap L^{2}(\Omega)\ \mathrm{as}\ \varepsilon_{n}^{\prime}\to 0$
for some $\tilde{u}\in S$. For all $s\in\mathbb{Q}\cap[0,T]$, we obtain
$W_{m}(\tilde{u},u(t))$ $\leq W_{m}(\tilde{u},u(s))+W_{m}(u(s),u(t))$
$\leq\liminf_{\varepsilon_{n}^{\prime}\to 0}W_{m_{\varepsilon_{n}^{\prime}}}(u_{\varepsilon_{n}^{\prime}}(t),u_{\varepsilon_{n}^{\prime}}(s))+W_{m}(u(s),u(t))$
$\leq C_{4}\sqrt{|t-s|}+W_{m}(u(s),u(t)).$
Letting $s\to t$, we have $W_{m}(\tilde{u},u(t))\leq 0$ and since $\tilde{u},u(t)\in S$, we see $\tilde{u}=u(t)$ in $S$. ∎
Remark 6.4.
Let $T>0$. Since the estimates in Lemma 5.2 are independent of $\varepsilon$, by the same arguments for $\tau$, we can easily see
$u_{\varepsilon_{n}}\rightharpoonup u\quad\mathrm{weakly\ in}\ L^{2}((0,T);H^{1}(\Omega))\ \mathrm{as}\ n\to\infty,$
$u_{\varepsilon_{n}}\to u\quad\mathrm{strongly\ in}\ L^{2}(\Omega\times(0,T))\ \mathrm{as}\ n\to\infty,$
$v_{\varepsilon_{n}}\rightharpoonup v\quad\mathrm{weakly\ in}\ L^{2}((0,T);H^{2}(\Omega))\ \mathrm{as}\ n\to\infty,$
$v_{\varepsilon_{n}}(t)\rightharpoonup v(t)\quad\mathrm{weakly\ in}\ H^{1}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T],$
$v\in C^{\frac{1}{2}}([0,T];L^{2}(\Omega)).$
Finally, we prove Theorem 1.1.
Proof of Theorem 1.1.
Let $T>0,\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ and $\zeta\in H^{1}(\Omega)$. By Lemma 6.3 and Remark 6.4, we can easily get
$\bullet u\in L^{\infty}((0,T);L^{2}(\Omega))\cap L^{2}((0,T);H^{1}(\Omega)),$
$\bullet\|u(t)\|_{L^{1}(\Omega)}=1\quad\forall t\in[0,T],$
$\bullet v\in L^{\infty}((0,T);H^{1}(\Omega))\cap L^{2}((0,T);H^{2}(\Omega))\cap C^{\frac{1}{2}}([0,T];L^{2}(\Omega)),$
$\bullet\lim_{t\to 0}W_{m}(u(t),u_{0})=0\ \mathrm{and}\ \lim_{t\to 0}\|v(t)-v_{0}\|_{L^{2}(\Omega)}=0.$
In addition, using Lemma 6.1, Lemma 6.2 and Remark 6.4, as in the proof of Lemma 6.1 and Lemma 6.2, we obtain that a pair of nonnegative functions $(u,v)$ satisfies the following weak formulations:
$\int_{0}^{T}\int_{\Omega}u^{\alpha}((1+\alpha)\nabla u-\chi\nabla v)\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u_{0}-u(\cdot,T))\varphi(x)\,dx,$
$\int_{0}^{T}\int_{\Omega}\left[\nabla v\cdot\nabla\zeta+v\zeta-u\zeta\right]\,dx\,dt=\int_{\Omega}(v_{0}-v(\cdot,T))\zeta\,dx.$
We complete the proof. ∎
7 Proof of Theorem 1.2 and Theorem 1.4
Even if $1+\alpha-2/d\leq p<1+\alpha$, the estimates and convergences of $v_{\tau}$ is same for $p=1+\alpha$. Therefore, we only consider about $u_{\tau}$. Most of arguments are not very different from the case for $p=1+\alpha$, but we need additional estimates of it in order to obtain the weak formulation in Theorem 1.2 and Theorem 1.4.
Lemma 7.1.
Let $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Let $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be a minimizer of (3.1). Then $u_{\tau}^{k}\in W^{1,p+1-\alpha}(\Omega)$.
Proof.
Set $w=(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}$, then by Lemma 4.6, it holds $w\in H^{1}(\Omega)$. Since $2/(p+1-\alpha)>1$, for $f\in C_{c}^{\infty}(\Omega)$, we have
$\int_{\Omega}u_{\tau}^{k}\nabla f\,dx=\int_{\Omega}w^{\frac{2}{p+1-\alpha}}\nabla f\,dx$ $=-\int_{\Omega}(\nabla w^{\frac{2}{p+1-\alpha}})f\,dx$
$=-\int_{\Omega}\left(\frac{2}{p+1-\alpha}w^{\frac{2}{p+1-\alpha}-1}\nabla w\right)f\,dx$
$=-\int_{\Omega}\left(\frac{2}{p+1-\alpha}(u_{\tau}^{k})^{\frac{1+\alpha-p}{2}}\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\right)f\,dx.$
Here, by Hölder’s inequality, it follows
$\int_{\Omega}|(u_{\tau}^{k})^{\frac{1+\alpha-p}{2}}\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}|^{p+1-\alpha}\,dx$
$\leq\left(\int_{\Omega}(u_{\tau}^{k})^{p+1-\alpha}\,dx\right)^{\frac{1+\alpha-p}{2}}\left(\int_{\Omega}|\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}|^{2}\,dx\right)^{\frac{p+1-\alpha}{2}}.$
Since $u_{\tau}^{k}\in L^{p+1-\alpha}(\Omega)$ and $\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in L^{2}(\Omega)$, we have $(u_{\tau}^{k})^{\frac{1+\alpha-p}{2}}\nabla(u_{\tau}^{k})^{\frac{p+1-\alpha}{2}}\in L^{p+1-\alpha}(\Omega)$. This means that $\nabla u_{\tau}^{k}\in L^{p+1-\alpha}(\Omega)$, that is, $u_{\tau}^{k}\in W^{1,p+1-\alpha}(\Omega)$. ∎
Next lemma is the modification of Lemma 4.8. When $1+\alpha-2/d\leq p<1+\alpha$, because of $p-\alpha<1$, we do not know whether $\nabla(u_{\tau}^{k})^{p-\alpha}$ is integrable. Hence we will obtain the modified inequality which does not have the term $\nabla(u_{\tau}^{k})^{p-\alpha}$.
Lemma 7.2.
Let $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Let $(u_{\tau}^{k-1},v_{\tau}^{k-1})\in X$ and $(u_{\tau}^{k},v_{\tau}^{k})\in X$ be a minimizer of (3.1). Then
$\frac{\boldsymbol{V}_{\delta}(u_{\tau}^{k})-\boldsymbol{V}_{\delta}(u_{\tau}^{k-1})}{\chi}$ $\leq\tau\left[\frac{p}{\chi(p-\alpha)}\int_{\Omega}(u_{\tau}^{k})^{p-\alpha}\nabla\cdot(m_{\varepsilon}(u_{\tau}^{k})\nabla\varphi)\,dx+\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\right]$
$\quad-\tau\lambda_{\delta}(E(u_{\tau}^{k-1},v_{\tau}^{k-1})-E(u_{\tau}^{k},v_{\tau}^{k}))+\tau\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx.$
Proof.
Let $w_{\delta}\in C([0,\infty);W^{1,p+1-\alpha}(\Omega))\cap C((0,\infty);W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,\infty);L^{p+1-\alpha}(\Omega))$ be a nonnegative solution to
$\begin{cases}\partial_{t}w_{\delta}=\delta\Delta w_{\delta}+\nabla\cdot(m_{\varepsilon}(w_{\delta})\nabla\varphi)\quad&\mathrm{a.e.\ in}\ \Omega\times(0,\infty),\\ \nabla w_{\delta}\cdot\boldsymbol{n}=0\ &\mathrm{on}\ \partial\Omega\times(0,\infty),\\ w_{\delta}(0)=u_{\tau}^{k}\ &\mathrm{in}\ L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega),\end{cases}$
where $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ (Propositions 4.1 and 4.2). Then, by the similar calculation of Lemma 4.8, we have
$\frac{d}{ds}E(w_{\delta}(s),v_{\tau}^{k})$
$=\int_{\Omega}\frac{p}{\chi(p-\alpha)}(w_{\delta}(s))^{p-\alpha}(\partial_{s}w_{\delta}(s))\,dx-\int_{\Omega}v_{\tau}^{k}(\partial_{s}w_{\delta}(s))\,dx$
$=\frac{p}{\chi(p-\alpha)}\int_{\Omega}(w_{\delta}(s))^{p-\alpha}[\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)+\delta\Delta w_{\delta}(s)]\,dx$
$\quad-\int_{\Omega}v_{\tau}^{k}[\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)+\delta\Delta w_{\delta}(s)]\,dx$
$=\frac{p}{\chi(p-\alpha)}\int_{\Omega}(w_{\delta}(s))^{p-\alpha}\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)\,dx+\frac{\delta p}{\chi(p-\alpha)}\int_{\Omega}(w_{\delta}(s))^{p-\alpha}\Delta w_{\delta}(s)\,dx$
$\quad+\int_{\Omega}m_{\varepsilon}(w_{\delta}(s))\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx+\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla w_{\delta}(s)\,dx.$
Note that it holds $v_{\tau}^{k}\in H^{2}(\Omega)$ by Corollary 4.10, and the Sobolev embedding theorem yields $H^{1}(\Omega)\hookrightarrow L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)$. Since $w_{\delta}(s)\to u_{\tau}^{k}$ in $W^{1,p+1-\alpha}(\Omega)$ as $s\to 0$, it is easy to check that the third term and the fourth term converge:
$\int_{\Omega}m_{\varepsilon}(w_{\delta}(s))\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\to\int_{\Omega}m_{\varepsilon}(u_{\tau}^{k})\nabla v_{\tau}^{k}\cdot\nabla\varphi\,dx\quad\mathrm{as}\ s\to 0,$
$\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla w_{\delta}(s)\,dx\to\delta\int_{\Omega}\nabla v_{\tau}^{k}\cdot\nabla u_{\tau}^{k}\,dx\quad\mathrm{as}\ s\to 0.$
Hence we will show
$\int_{\Omega}(w_{\delta}(s))^{p-\alpha}\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)\,dx\to\int_{\Omega}(u_{\tau}^{k})^{p-\alpha}\nabla\cdot(m_{\varepsilon}(u_{\tau}^{k})\nabla\varphi)\,dx\quad\mathrm{as}\ s\to 0,$
$\int_{\Omega}(w_{\delta}(s))^{p-\alpha}\Delta w_{\delta}(s)\,dx\leq 0\quad\mathrm{for}\ s>0.$
First, note that
$\bullet\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)=\frac{\alpha\nabla w_{\delta}(s)\cdot\nabla\varphi}{(w_{\delta}(s)+\varepsilon)^{1-\alpha}}+(w_{\delta}(s)+\varepsilon)^{\alpha}\Delta\varphi,$
$\bullet\|(w_{\delta}(s)+\varepsilon)^{\alpha}\Delta\varphi\|_{L^{p+1-\alpha}(\Omega)}\leq\tilde{C}(\varphi,\alpha,\varepsilon)(\|w_{\delta}(s)\|_{L^{p+1-\alpha}}^{\alpha}+1).$
Then,
$\left|\int_{\Omega}(w_{\delta}(s))^{p-\alpha}\nabla\cdot(m_{\varepsilon}(w_{\delta}(s))\nabla\varphi)\,dx-\int_{\Omega}(u_{\tau}^{k})^{p-\alpha}\nabla\cdot(m_{\varepsilon}(u_{\tau}^{k})\nabla\varphi)\,dx\right|$
$\leq\int_{\Omega}|(w_{\delta}(s))^{p-\alpha}-(u_{\tau}^{k})^{p-\alpha}|\left|\frac{\alpha\nabla w_{\delta}(s)\cdot\nabla\varphi}{(w_{\delta}(s)+\varepsilon)^{1-\alpha}}+(w_{\delta}(s)+\varepsilon)^{\alpha}\Delta\varphi\right|\,dx$
$\quad+\int_{\Omega}(u_{\tau}^{k})^{p-\alpha}\left|\frac{\alpha\nabla w_{\delta}(s)\cdot\nabla\varphi}{(w_{\delta}(s)+\varepsilon)^{1-\alpha}}-\frac{\alpha\nabla u_{\tau}^{k}\cdot\nabla\varphi}{(u_{\tau}^{k}+\varepsilon)^{1-\alpha}}\right|\,dx$
$\quad+\int_{\Omega}(u_{\tau}^{k})^{p-\alpha}|(w_{\delta}(s)+\varepsilon)^{\alpha}-(u_{\tau}^{k}+\varepsilon)^{\alpha}||\Delta\varphi|\,dx$
$\eqcolon\mathrm{I}_{1}+\mathrm{I}_{2}+\mathrm{I}_{3}.$
Since $|(w_{\delta}(s))^{p-\alpha}-(u_{\tau}^{k})^{p-\alpha}|\leq|w_{\delta}(s)-u_{\tau}^{k}|^{p-\alpha}$, by Hölder’s inequality, we have
$\mathrm{I}_{1}\leq\|w_{\delta}(s)-u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p-\alpha}C(\varphi,\alpha,\varepsilon)(\|\nabla w_{\delta}(s)\|_{L^{p+1-\alpha}(\Omega)}+\|w_{\delta}(s)\|_{L^{p+1-\alpha}(\Omega)}^{\alpha}+1).$
Since $w_{\delta}(s)\to u_{\tau}^{k}$ in $L^{p+1-\alpha}(\Omega)$ as $s\to 0$ and $\sup_{s\in[0,1]}\|w_{\delta}(s)\|_{W^{1,p+1-\alpha}(\Omega)}<\infty$, we obtain $\mathrm{I}_{1}\to 0$ as $s\to 0$. Further, by Hölder’s inequality and the mean value theorem, we have
$\mathrm{I}_{2}\leq\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p-\alpha}C(\varphi,\alpha,\varepsilon)\|w_{\delta}(s)-u_{\tau}^{k}\|_{W^{1,p+1-\alpha}(\Omega)}.$
Since $w_{\delta}(s)\to u_{\tau}^{k}$ in $W^{1,p+1-\alpha}(\Omega)$ as $s\to 0$, we obtain $\mathrm{I}_{2}\to 0$ as $s\to 0$. Similarly we have
$\mathrm{I}_{3}\leq\|u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{p-\alpha}C(\varphi,\alpha,\varepsilon)\|w_{\delta}(s)-u_{\tau}^{k}\|_{L^{p+1-\alpha}(\Omega)}^{\alpha}\to 0\quad\mathrm{as}\ s\to 0.$
Finally, set $w_{n}\coloneq w_{\delta}+1/n$, then $w_{n}$ still satisfies $\nabla w_{n}\cdot\boldsymbol{n}=0$ on $\partial\Omega$. By integration by parts, it follows
$\int_{\Omega}(\Delta w_{\delta}(s))w_{n}^{p-\alpha}\,dx$ $=-\int_{\Omega}\nabla w_{\delta}(s)\cdot\nabla(w_{n})^{p-\alpha}\,dx$
$=-\int_{\Omega}(p-\alpha)\frac{|\nabla w_{\delta}(s)|^{2}}{(w_{n})^{1+\alpha-p}}\,dx\leq 0\quad\mathrm{for}\ s>0.$
Since $(w_{n})^{p-\alpha}\to w_{\delta}^{p-\alpha}(s)$ a.e. in $\Omega$ as $n\to\infty$ and $|\Delta w_{\delta}(s)|w_{n}^{p-\alpha}\leq|\Delta w_{\delta}(s)|(w_{\delta}^{p-\alpha}(s)+1)\in L^{1}(\Omega)$ due to $\Delta w_{\delta}(s)\in L^{p+1-\alpha}(\Omega)$ and $w_{\delta}^{p-\alpha}(s)\in L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)$, we infer from Lebesgue’s dominated convergence theorem that
$\int_{\Omega}(\Delta w_{\delta}(s))w_{\delta}^{p-\alpha}(s)\,dx=\lim_{n\to\infty}\int_{\Omega}(\Delta w_{\delta}(s))w_{n}^{p-\alpha}\,dx\leq 0\quad\mathrm{for}\ s>0.$
The rest of the proof can be shown by the same argument in the proof of Lemma 4.8. ∎
Next, we will prepare some uniform estimates and convergences to prove Theorem 1.2 and Theorem 1.4.
Lemma 7.3.
Let $T>0$, $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then there exist positive constants $C_{5}$ and $C_{6}$ depending on $\alpha,p,d,\chi,u_{0}$ and $v_{0}$ such that
$\int_{0}^{T}\|u_{\tau}(t)\|_{L^{2}(\Omega)}^{2}\,dt\leq C_{5}(1+T),$ (7.1)
$\int_{0}^{T}\|u_{\tau}(t)\|_{W^{1,p+1-\alpha}(\Omega)}^{2}\,dt\leq C_{6}(1+T).$ (7.2)
Proof.
By (4.7), (5.1) and (5.2), we immediately obtain (7.1) for some constant $C_{5}$. On the other hand, observe that it follows from Lemma 7.1 that
$\nabla u_{\tau_{n}}(t)=\frac{2}{p+1-\alpha}u_{\tau_{n}}^{\frac{1+\alpha-p}{2}}(t)\nabla(u_{\tau_{n}}(t))^{\frac{p+1-\alpha}{2}}\quad\mathrm{a.e.\ in}\ \Omega\ \mathrm{for}\ t\in[0,T].$ (7.3)
Then we infer from (5.1) and (5.2) that
$\int_{0}^{T}\left(\int_{\Omega}|\nabla u_{\tau}|^{p+1-\alpha}\,dx\right)^{\frac{2}{p+1-\alpha}}\,dt$
$\leq\left(\frac{2}{p+1-\alpha}\right)^{2}\int_{0}^{T}\left(\int_{\Omega}(u_{\tau})^{p+1-\alpha}\,dx\right)^{\frac{1+\alpha-p}{p+1-\alpha}}\left(\int_{\Omega}|\nabla(u_{\tau})^{\frac{p+1-\alpha}{2}}|^{2}\,dx\right)\,dt$
$\leq\left(\frac{2}{p+1-\alpha}\right)^{2}\left(\sup_{0\leq t\leq T}\|u_{\tau}(t)\|_{L^{p+1-\alpha}(\Omega)}^{p+1-\alpha}\,dt\right)^{\frac{1+\alpha-p}{p+1-\alpha}}\left(\int_{0}^{T}\|\nabla(u_{\tau}(t))^{\frac{p+1-\alpha}{2}}\|_{L^{2}(\Omega)}^{2}\,dt\right)$
$\leq\left(\frac{2}{p+1-\alpha}\right)^{2}C_{1}^{\frac{1+\alpha-p}{p+1-\alpha}}C_{2}(1+T).$
Thus, combining this estimate with (5.1), we obtain (7.2) and complete the proof. ∎
From Lemma 7.3, Lemma 5.3 and the refined Ascoli–Arzelà theorem ([2, Proposition 3.3.1]), we immediately obtain next corollary.
Corollary 7.4.
Let $T>0$, $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then there exist a subsequence $\{u_{\tau_{n}}\}_{n}$ with $\tau_{n}\to 0$ as $n\to\infty$ and $u_{\varepsilon}:[0,T]\to\mathcal{P}(\Omega)$ such that
$u_{\tau_{n}}\rightharpoonup u_{\varepsilon}\quad\mathrm{weakly\ in}\ L^{2}((0,T);W^{1,p+1-\alpha}(\Omega))\ \mathrm{as}\ n\to\infty,$
$u_{\tau_{n}}(t)\rightharpoonup u_{\varepsilon}(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{p+1-\alpha}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T].$
We next get the strong convergence for $\tau$ as in the Lemma 5.5.
Lemma 7.5.
Let $T>0$, $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then for the sequence $\{u_{\tau_{n}}\}_{n}$ in Corollary 7.4, it holds that
$u_{\tau_{n}}\to u_{\varepsilon}\quad\mathrm{strongly\ in}\ L^{2}(\Omega\times(0,T))\ \mathrm{as}\ n\to\infty.$
Proof.
Since $p\geq 1+\alpha-2/d$ and $d\geq 2$ then
$(p+1-\alpha)^{*}\coloneq\frac{d(p+1-\alpha)}{d-(p+1-\alpha)}\geq 2,$
and by the Rellich–Kondrachov theorem, $W^{1,p+1-\alpha}(\Omega)$ is compactly embedded in $L^{2}(\Omega)$. Hence $L^{2}(\Omega)$ is compactly embedded in $(W^{1,p+1-\alpha}(\Omega))^{*}$, where $(W^{1,p+1-\alpha}(\Omega))^{*}$ is the dual space of $W^{1,p+1-\alpha}(\Omega)$. By Lemma 5.2, $\|u_{\tau_{n}}(t)\|_{L^{p+1-\alpha}(\Omega)}$ is bounded with respect to $\tau_{n}$ for all $t\in[0,T]$. In addtition, we know that $u_{\tau_{n}}(t)$ weakly converges to $u_{\varepsilon}(t)$ in $L^{p+1-\alpha}(\Omega)$. Thus by the same argument in the proof of Lemma 5.5, we conclude that $u_{\tau_{n}}$ strongly converges to $u_{\varepsilon}$ in $L^{2}(\Omega\times(0,T))$ as $n\to\infty$. ∎
Since $0<\alpha<1$, it is not easy to get the integrability of the derivative of $u_{\tau_{n}}^{\alpha}$. But we know that $|\nabla u_{\tau_{n}}^{\frac{p+1-\alpha}{2}}|^{2}$ is integrable in $\Omega$, so using this integrability, we obtain the estimate of the derivative of $u_{\tau_{n}}^{\alpha}$.
Lemma 7.6.
Let $T>0$, $(1+p)/3\leq\alpha<1$ and $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then for the sequence $\{u_{\tau_{n}}\}_{n}$ in Corollary 7.4, $\nabla m_{\varepsilon}(u_{\tau_{n}})=\nabla(u_{\tau_{n}}+\varepsilon)^{\alpha}$ weakly converges to $\nabla(u_{\varepsilon}+\varepsilon)^{\alpha}=\nabla m_{\varepsilon}(u_{\varepsilon})$ in $L^{q}(\Omega\times(0,T))$ as $n\to\infty$ where $q=4/(3\alpha+1-p)\in(1,2]$. In addition, $m_{\varepsilon}(u_{\varepsilon})\in L^{q}((0,T);W^{1,q}(\Omega))$ and there exists a constant $C_{7}>0$ depending on $\alpha,p,d,\chi,u_{0}$ and $v_{0}$ such that the following estimate holds:
$\int_{0}^{T}\int_{\Omega}|\nabla m_{\varepsilon}(u_{\tau_{n}})|^{q}\,dx\,dt\leq C_{7}(1+T).$ (7.4)
Proof.
Note that since $(1+p)/3\leq\alpha$, it hold $3\alpha-1-p\geq 0$ and $q\leq 2$. By Hölder’s inequality, (5.2) and (7.1), we have
$\int_{0}^{T}\int_{\Omega}|\nabla(u_{\tau_{n}}+\varepsilon)^{\alpha}|^{q}\,dx\,dt$
$=\int_{0}^{T}\int_{\Omega}\left|\frac{\alpha\nabla u_{\tau_{n}}}{(u_{\tau_{n}}+\varepsilon)^{1-\alpha}}\right|^{q}\,dx\,dt$
$=\int_{0}^{T}\int_{\Omega}\left|\frac{\alpha}{(u_{\tau_{n}}+\varepsilon)^{1-\alpha}}\frac{2}{p+1-\alpha}u_{\tau_{n}}^{\frac{1+\alpha-p}{2}}\nabla u_{\tau_{n}}^{\frac{p+1-\alpha}{2}}\right|^{q}\,dx\,dt$
$\leq\int_{0}^{T}\int_{\Omega}\left|\frac{2\alpha}{p+1-\alpha}u_{\tau_{n}}^{\frac{3\alpha-1-p}{2}}\nabla u_{\tau_{n}}^{\frac{p+1-\alpha}{2}}\right|^{q}\,dx\,dt$
$\leq\left(\frac{2\alpha}{p+1-\alpha}\right)^{q}\int_{0}^{T}\left(\int_{\Omega}|\nabla u_{\tau_{n}}^{\frac{p+1-\alpha}{2}}|^{2}\,dx\right)^{\frac{q}{2}}\left(\int_{\Omega}u_{\tau_{n}}^{2}\,dx\right)^{\frac{2-q}{2}}\,dt$
$\leq\left(\frac{2\alpha}{p+1-\alpha}\right)^{q}\left(\int_{0}^{T}\|\nabla u_{\tau_{n}}^{\frac{p+1-\alpha}{2}}(t)\|_{L^{2}(\Omega)}^{2}\,dt\right)^{\frac{q}{2}}\left(\int_{0}^{T}\|u_{\tau_{n}}(t)\|_{L^{2}(\Omega)}^{2}\,dt\right)^{\frac{2-q}{2}}$
$\leq\left(\frac{2\alpha}{p+1-\alpha}\right)^{q}(C_{2}(1+T))^{\frac{q}{2}}(C_{5}(1+T))^{\frac{2-q}{2}}.$
Thus the estimate (7.4) holds for some constant $C_{7}$, and there exist a subsequence (not relabeled) and $y_{\varepsilon}\in L^{q}(\Omega\times(0,T))$ such that $\nabla m_{\varepsilon}(u_{\tau_{n}})\rightharpoonup y_{\varepsilon}$ weakly in $L^{q}(\Omega\times(0,T))$ as $n\to\infty$. Let $f\in C_{c}^{\infty}(\Omega\times(0,T))$, then we have
$\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla f\,dx\,dt$ $=-\int_{0}^{T}\int_{\Omega}(\nabla m_{\varepsilon}(u_{\tau_{n}}))f\,dx\,dt$
$\to-\int_{0}^{T}\int_{\Omega}y_{\varepsilon}f\,dx\,dt\quad\mathrm{as}\ n\to\infty.$
On the other hand, we know that $u_{\tau_{n}}\to u_{\varepsilon}$ strongly in $L^{2}(\Omega\times(0,T))$ as $n\to\infty$ (Lemma 7.5), so that
$\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla f\,dx\,dt\to\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla f\,dx\,dt\quad\mathrm{as}\ n\to\infty.$
Here, since $u_{\varepsilon}\in L^{2}((\Omega\times(0,T)))$ by Lemma 7.5 and $\alpha q<2$, then we have $m_{\varepsilon}(u_{\varepsilon})\in L^{q}(\Omega\times(0,T))$. Hence, we obtain $m_{\varepsilon}(u_{\varepsilon})\in L^{q}((0,T);W^{1,q}(\Omega))$ and $y_{\varepsilon}=\nabla m_{\varepsilon}(u_{\varepsilon})$ a.e. in $\Omega\times(0,T)$, that is, $\nabla m_{\varepsilon}(u_{\tau_{n}})\rightharpoonup\nabla m_{\varepsilon}(u_{\varepsilon})$ weakly in $L^{q}(\Omega\times(0,T))$ as $n\to\infty$. ∎
Next, we establish the weak formulation of the first equation of (1.1)ε as in the Lemma 6.1.
Lemma 7.7.
Let $(1+p)/3\leq\alpha<1$ and $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then $(u_{\varepsilon},v_{\varepsilon})$ satisfies the following weak formulation: for all $T>0$ and $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$, it holds
$\int_{\Omega}(u_{0}(x)-u_{\varepsilon}(x,T))\varphi(x)\,dx$ $=-\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}u_{\varepsilon}^{p-\alpha}(x,t)\nabla\cdot[m_{\varepsilon}(u_{\varepsilon}(x,t))\nabla\varphi(x)]\,dx\,dt$ (7.5)
$\quad-\int_{0}^{T}\int_{\Omega}\chi m_{\varepsilon}(u_{\varepsilon}(x,t))\nabla v_{\varepsilon}(x,t)\cdot\nabla\varphi(x)\,dx\,dt.$
Proof.
Let $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$. Let $\{\tau_{n}\}\subset(0,1)$ be a subsequence of $\{\tau\}$ satisfying $\tau_{n}\to 0$ as $n\to\infty$ and obtained in Corollary 7.4, and set $\delta_{n}\coloneq\tau_{n}^{\frac{1}{2}}$. We can prove most of this lemma as in the proof of Lemma 6.1 using Lemma 7.2. By Corollary 7.4, it is easy to check that
$\int_{\Omega}u_{\tau_{n}}(x,T)\varphi(x)\,dx\to\int_{\Omega}u_{\varepsilon}(x,T)\varphi(x)\,dx\quad\mathrm{as}\ n\to\infty,$
$\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\tau_{n}})\nabla v_{\tau_{n}}\cdot\nabla\varphi\,dx\,dt\to\int_{0}^{T}\int_{\Omega}m_{\varepsilon}(u_{\varepsilon})\nabla v_{\varepsilon}\cdot\nabla\varphi\,dx\,dt\quad\mathrm{as}\ n\to\infty.$
Hence we will show
$\int_{0}^{T}\int_{\Omega}(u_{\tau_{n}})^{p-\alpha}\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]\,dx\,dt\to\int_{0}^{T}\int_{\Omega}(u_{\varepsilon})^{p-\alpha}\nabla\cdot[m_{\varepsilon}(u_{\varepsilon})\nabla\varphi]\,dx\,dt\quad\mathrm{as}\ n\to\infty.$
Indeed, we have
$=\left|\int_{0}^{T}\int_{\Omega}(u_{\tau_{n}})^{p-\alpha}\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]\,dx\,dt-\int_{0}^{T}\int_{\Omega}(u_{\varepsilon})^{p-\alpha}\nabla\cdot[m_{\varepsilon}(u_{\varepsilon})\nabla\varphi]\,dx\,dt\right|$
$\leq\int_{0}^{T}\int_{\Omega}|(u_{\tau_{n}})^{p-\alpha}-(u_{\varepsilon})^{p-\alpha}||\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]|\,dx\,dt$
$\quad+\left|\int_{0}^{T}\int_{\Omega}(u_{\varepsilon})^{p-\alpha}(\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]-\nabla\cdot[m_{\varepsilon}(u_{\varepsilon})\nabla\varphi])\,dx\,dt\right|$
$\eqcolon\mathrm{I}_{1}+\mathrm{I}_{2}.$
Note that $\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]=\nabla m_{\varepsilon}(u_{\tau_{n}})\cdot\nabla\varphi+m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi$. We infer from Hölder’s inequality and Minkowski’s inequality that
$\mathrm{I}_{1}$ $\leq\int_{0}^{T}\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{p-\alpha}|\nabla\cdot[m_{\varepsilon}(u_{\tau_{n}})\nabla\varphi]|\,dx\,dt$
$\leq\int_{0}^{T}\left(\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{2}\,dx\right)^{\frac{p-\alpha}{2}}\left(\int_{\Omega}|\nabla m_{\varepsilon}(u_{\tau_{n}})\cdot\nabla\varphi+m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi|^{\frac{2}{2-p+\alpha}}\,dx\right)^{\frac{2-p+\alpha}{2}}\,dt$
$\leq\int_{0}^{T}\left(\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{2}\,dx\right)^{\frac{p-\alpha}{2}}\left(\int_{\Omega}|\nabla m_{\varepsilon}(u_{\tau_{n}})\cdot\nabla\varphi|^{\frac{2}{2-p+\alpha}}\,dx\right)^{\frac{2-p+\alpha}{2}}\,dt$
$\quad+\int_{0}^{T}\left(\int_{\Omega}|u_{\tau_{n}}-u_{\varepsilon}|^{2}\,dx\right)^{\frac{p-\alpha}{2}}\left(\int_{\Omega}|m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi|^{\frac{2}{2-p+\alpha}}\,dx\right)^{\frac{2-p+\alpha}{2}}\,dt$
$\leq\|u_{\tau_{n}}-u_{\varepsilon}\|_{L^{2}(\Omega\times(0,T))}^{p-\alpha}\left(\int_{0}^{T}\int_{\Omega}|\nabla m_{\varepsilon}(u_{\tau_{n}})\cdot\nabla\varphi|^{\frac{2}{2-p+\alpha}}\,dx\,dt\right)^{\frac{2-p+\alpha}{2}}$
$\quad+\|u_{\tau_{n}}-u_{\varepsilon}\|_{L^{2}(\Omega\times(0,T))}^{p-\alpha}\left(\int_{0}^{T}\int_{\Omega}|m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi|^{\frac{2}{2-p+\alpha}}\,dx\,dt\right)^{\frac{2-p+\alpha}{2}}.$
Here, we see
$\frac{2}{2-p+\alpha}\leq q=\frac{4}{3\alpha+1-p}\Leftrightarrow 3\alpha+1-p\leq 2(2-p+\alpha)\Leftrightarrow\alpha+p\leq 3$
and $\alpha+p\leq 3$ is true, hence by using Lemma 7.6, it follows that $\|\nabla m_{\varepsilon}(u_{\tau_{n}})\|_{L^{\frac{2}{2-p+\alpha}}(\Omega\times(0,T))}$ is bounded with respect to $n$. In addition, since
$|m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi|=|(u_{\tau_{n}}+\varepsilon)^{\alpha}\Delta\varphi|\leq|u_{\tau_{n}}^{\alpha}\Delta\varphi|+\varepsilon^{\alpha}|\Delta\varphi|$
and
$\frac{2\alpha}{2-p+\alpha}\leq 2\Leftrightarrow p\leq 2,$
we infer from (7.1) that $\|m_{\varepsilon}(u_{\tau_{n}})\Delta\varphi\|_{L^{\frac{2}{2-p+\alpha}}(\Omega\times(0,T))}$ is also bounded with respect to $n$. Combining these boundedness with Lemma 7.5, we obtain $\mathrm{I}_{1}\to 0$ as $n\to\infty$.
On the other hand, we have
$\mathrm{I}_{2}$ $\leq\left|\int_{0}^{T}\int_{\Omega}(u_{\varepsilon})^{p-\alpha}(\nabla m_{\varepsilon}(u_{\tau_{n}})-\nabla m_{\varepsilon}(u_{\varepsilon}))\cdot\nabla\varphi\,dx\,dt\right|$
$\quad+\int_{0}^{T}\int_{\Omega}(u_{\varepsilon})^{p-\alpha}|m_{\varepsilon}(u_{\tau_{n}})-m_{\varepsilon}(u_{\varepsilon})||\Delta\varphi|\,dx\,dt.$
Since $u_{\varepsilon}^{p-\alpha}\nabla\varphi\in(L^{q}(\Omega\times(0,T)))^{*}$, where $(L^{q}(\Omega\times(0,T)))^{*}$ is the dual space of $L^{q}(\Omega\times(0,T))$, Lemma 7.6 implies that the first term converges to 0 as $n\to\infty$. Moreover, as in the above argument, Lemma 7.5 yields that the second term also converges to 0 as $n\to\infty$. Hence we have $\mathrm{I}_{2}\to 0$ as $n\to\infty$. The proof is completed. ∎
The following lemma is about the convergences with respect to $\varepsilon$.
Lemma 7.8.
Let $T>0$, $(1+p)/3\leq\alpha<1$ and $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then there exist a subsequence $\{u_{\varepsilon_{n}}\}_{n}$ with $\varepsilon_{n}\to 0$ as $n\to\infty$ and a function $u$ such that
$u_{\varepsilon_{n}}\rightharpoonup u\quad\mathrm{weakly\ in}\ L^{2}((0,T);W^{1,p+1-\alpha}(\Omega))\ \mathrm{as}\ n\to\infty,$
$u_{\varepsilon_{n}}(t)\rightharpoonup u(t)\quad\mathrm{weakly\ in}\ L^{1}\cap L^{p+1-\alpha}(\Omega)\ \mathrm{as}\ n\to\infty\ \mathrm{for}\ t\in[0,T],$
$u_{\varepsilon_{n}}\to u\quad\mathrm{strongly\ in}\ L^{2}(\Omega\times(0,T))\ \mathrm{as}\ n\to\infty,$
$\nabla m_{\varepsilon_{n}}(u_{\varepsilon_{n}})\rightharpoonup\nabla u^{\alpha}\quad\mathrm{weakly\ in}\ L^{q}(\Omega\times(0,T))\ \mathrm{as}\ n\to\infty,$
$W_{m}(u(t),u(s))\leq C_{4}\sqrt{|t-s|}\ \mathrm{for}\ t,s\in[0,T],$
where $q=4/(3\alpha+1-p)\in(1,2]$. In particular, $u^{\alpha}\in L^{q}((0,T);W^{1,q}(\Omega))$.
Proof.
This lemma is followed by the similar argument in Lemma 6.3 and the same arguments of Lemma 7.5 and Lemma 7.6 replaced $\tau_{n}$ by $\varepsilon_{n}$. ∎
We next show that the weak formulation of the first equation of (1.1) holds.
Lemma 7.9.
Let $(1+p)/3\leq\alpha<1$ and $1+\alpha-2/d\leq p<1+\alpha$ and assume that $\chi>0$ is small enough if $p=1+\alpha-2/d$. Then $(u,v)$ satisfies the following weak formulation: for all $T>0,\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$, it holds
$\int_{0}^{T}\int_{\Omega}(\nabla u^{p}-\chi u^{\alpha}\nabla v)\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u_{0}-u(\cdot,T))\varphi\,dx.$
Proof.
Let $T>0$ and $\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$. We infer from Lemma 7.7 and Lemma 7.8 that
$\int_{\Omega}(u_{0}(x)-u(x,T))\varphi(x)\,dx$ $=-\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}u^{p-\alpha}(x,t)\nabla\cdot[u^{\alpha}(x,t)\nabla\varphi(x)]\,dx\,dt$
$\quad-\int_{0}^{T}\int_{\Omega}\nabla\varphi(x)\cdot\chi u^{\alpha}(x,t)\nabla v(x,t)\,dx\,dt.$
By the same argument in the proof of Lemma 7.1, we have
$\nabla u=\frac{1}{\alpha}u^{1-\alpha}\nabla u^{\alpha}\quad\mathrm{a.e.\ in}\ \Omega\times(0,T),$
$\nabla u=\frac{2}{p+1-\alpha}u^{\frac{1+\alpha-p}{2}}\nabla u^{\frac{p+1-\alpha}{2}}\quad\mathrm{a.e.\ in}\ \Omega\times(0,T),$
so that
$u^{1-\alpha}\nabla u^{\alpha}=\frac{2\alpha}{p+1-\alpha}u^{\frac{1+\alpha-p}{2}}\nabla u^{\frac{p+1-\alpha}{2}}\quad\mathrm{a.e.\ in}\ \Omega\times(0,T).$
On the other hand, we see
$\nabla u^{p}=\frac{2p}{p+1-\alpha}u^{\frac{p-1+\alpha}{2}}\nabla u^{\frac{p+1-\alpha}{2}}\quad\mathrm{a.e.\ in}\ \Omega\times(0,T).$
Hence, by using these and integration by parts, it follows
$\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}u^{p-\alpha}(x,t)\nabla\cdot[u^{\alpha}(x,t)\nabla\varphi(x)]\,dx\,dt$
$=\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}\left[u^{p-1}(x,t)u^{1-\alpha}(x,t)\nabla u^{\alpha}(x,t)\cdot\nabla\varphi(x)+u^{p}(x,t)\Delta\varphi(x)\right]\,dx\,dt$
$=\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}\left[\frac{2\alpha}{p+1-\alpha}u^{\frac{p-1+\alpha}{2}}(x,t)\nabla u^{\frac{p+1-\alpha}{2}}(x,t)\varphi(x)-\nabla u^{p}(x,t)\cdot\nabla\varphi(x)\right]\,dx\,dt$
$=\int_{0}^{T}\int_{\Omega}\frac{p}{p-\alpha}\left(\frac{\alpha}{p}-1\right)\nabla u^{p}(x,t)\cdot\nabla\varphi(x)\,dx\,dt$
$=-\int_{0}^{T}\int_{\Omega}\nabla\varphi(x)\cdot\nabla u^{p}(x,t)\,dx\,dt.$
Thus we obtain
$\int_{\Omega}(u_{0}(x)-u(x,T))\varphi(x)\,dx=\int_{0}^{T}\int_{\Omega}\nabla\varphi(x)\cdot[\nabla u^{p}(x,t)-\chi u^{\alpha}(x,t)\nabla v(x,t)]\,dx\,dt.$
Finally, we prove Theorem 1.2 and Theorem 1.4.
Proof of Theorem 1.2 and Theorem 1.4.
Let $T>0,\varphi\in C^{\infty}(\overline{\Omega})$ with $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega$ and $\zeta\in H^{1}(\Omega)$. Note that (5.1) and (5.2), then by Lemma 7.8, we have
$\bullet u\in L^{\infty}((0,T);L^{p+1-\alpha}(\Omega)),\ u^{\frac{p+1-\alpha}{2}}\in L^{2}((0,T);H^{1}(\Omega)),$
$\bullet\|u(t)\|_{L^{1}(\Omega)}=1\quad\forall t\in[0,T],$
$\bullet\lim_{t\to 0}W_{m}(u(t),u_{0})=0.$
Moreover, even if $1+\alpha-2/d\leq p<1+\alpha$, we similarly obtain the convergence about $v$ in Remark 6.4, so that
$\bullet v\in L^{\infty}((0,T);H^{1}(\Omega))\cap L^{2}((0,T);H^{2}(\Omega))\cap C^{\frac{1}{2}}([0,T];L^{2}(\Omega)),$
$\bullet\lim_{t\to 0}\|v(t)-v_{0}\|_{L^{2}(\Omega)}=0.$
Using Lemma 7.9, Lemma 6.2 and Remark 6.4, as in the proof of Theorem 1.1, we conclude that a pair of nonnegative functions $(u,v)$ satisfies the following weak formulation:
$\int_{0}^{T}\int_{\Omega}(\nabla u^{p}-\chi u^{\alpha}\nabla v)\cdot\nabla\varphi\,dx\,dt=\int_{\Omega}(u_{0}-u(\cdot,T))\varphi\,dx,$
$\int_{0}^{T}\int_{\Omega}\left[\nabla v\cdot\nabla\zeta+v\zeta-u\zeta\right]\,dx\,dt=\int_{\Omega}(v_{0}-v(\cdot,T))\zeta\,dx.$
The proof is completed. ∎
Appendix A Appendix
Proof of Proposition 4.1.
We devide the proof into four steps. To simplify, we write $\|\cdot\|_{L^{q}(\Omega)}=\|\cdot\|_{q}$ for $q\in[1,\infty]$ and $\|\cdot\|_{W^{l,p+1-\alpha}(\Omega)}=\|\cdot\|_{W^{l,p+1-\alpha}}$ for $l\in\mathbb{N}$.
Step 1: Existence of a local soluiton.
Set $M_{0}\coloneq\|w_{0}\|_{W^{1,p+1-\alpha}}$ and
$Y\coloneq\{y\in C([0,T_{0}];W^{1,p+1-\alpha}(\Omega));\|y\|_{Y}\leq 4M_{0}\},$
where $\|y\|_{Y}\coloneq\sup_{0\leq t\leq T_{0}}\|y(t)\|_{W^{1,p+1-\alpha}}$ and $T_{0}\in(0,\infty)$ will be fixed later. We define a function $w_{1}$ by $w_{1}=e^{\delta t\Delta}w_{0}$, where $e^{\delta t\Delta}$ is the Neumann heat semigroup, that is, $w_{1}$ is a solution to
$\begin{cases}\partial_{t}w_{1}=\delta\Delta w_{1}\quad&\mathrm{in}\ \Omega\times(0,\infty),\\ \nabla w_{1}\cdot\boldsymbol{n}=0\quad&\mathrm{on}\ \partial\Omega\times(0,\infty),\\ w_{1}(0)=w_{0}\quad&\mathrm{in}\ L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega).\end{cases}$
Then $w_{1}$ is nonnegative and $w_{1}$ belongs to $C([0,\infty);L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega))\cap C^{\infty}(\Omega\times(0,\infty))$. We also define a function
$\Phi[w](t)\coloneq e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\left(\frac{\alpha\nabla w(s)}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}(s)+\varepsilon)^{\alpha}\Delta\varphi\right)\,ds$
$\hskip 284.52756pt\mathrm{for}\ w\in Y,\ t\in[0,T_{0}].$
Then $\Phi$ belongs to $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$ due to the property of the heat semigroup.
In this proof, we often use the following estimate:
$\|(y+\varepsilon)^{\alpha}\Delta\varphi\|_{p+1-\alpha}$ $\leq\|(y^{\alpha}+\varepsilon^{\alpha})\Delta\varphi\|_{p+1-\alpha}\leq\|y\|_{p+1-\alpha}^{\alpha}\|\Delta\varphi\|_{\frac{p+1-\alpha}{1-\alpha}}+\varepsilon^{\alpha}\|\Delta\varphi\|_{p+1-\alpha}$
$\leq C(\|y\|_{p+1-\alpha}^{\alpha}+1)\quad\mathrm{for}\ y\in Y\ \mathrm{and}\ y\geq 0\ \mathrm{a.e.},$ (A.1)
where $C\coloneq\max\{\|\Delta\varphi\|_{\frac{p+1-\alpha}{1-\alpha}},\varepsilon^{\alpha}\|\Delta\varphi\|_{p+1-\alpha}\}$ is a constant.
First, we show that $\Phi$ is a contraction map on $Y$ if $T_{0}$ is small enough. Let $w\in Y$ and $t\in[0,T_{0}]$ then using $L^{p}$-$L^{q}$ estimates and (A), we have
$\|\Phi[w](t)\|_{p+1-\alpha}$ $\leq\|w_{0}\|_{p+1-\alpha}+\int_{0}^{t}\left\|\frac{\alpha\nabla w(s)}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}(s)+\varepsilon)^{\alpha}\Delta\varphi\right\|_{p+1-\alpha}\,ds$
$\leq M_{0}+\int_{0}^{t}\frac{C^{\prime}\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|\nabla w(s)\|_{p+1-\alpha}+C(\|w_{1}(s)\|_{p+1-\alpha}^{\alpha}+1)\,ds$
$\leq M_{0}+C_{1}(M_{0}+1)T_{0},$
where $C^{\prime}$ is a constant by $L^{p}$-$L^{q}$ estimates and $C_{1}(\alpha,\varepsilon,\varphi,\Omega)$ is also a constant. If $T_{0}\leq M_{0}/C_{1}(M_{0}+1)$ then $\sup_{0\leq t\leq T_{0}}\|\Phi[w](t)\|_{p+1-\alpha}\leq 2M_{0}$. Similarly, it follows
$\|\nabla\Phi[w](t)\|_{p+1-\alpha}$
$\leq\|\nabla w_{0}\|_{p+1-\alpha}+\int_{0}^{t}\frac{\tilde{C}}{(t-s)^{\frac{1}{2}}}\left\|\frac{\alpha\nabla w(s)}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}(s)+\varepsilon)^{\alpha}\Delta\varphi\right\|_{p+1-\alpha}\,ds$
$\leq M_{0}+C_{2}(M_{0}+1)T_{0}^{\frac{1}{2}},$
where $\tilde{C}$ is a constant by $L^{p}$-$L^{q}$ estimates and $C_{2}=C_{2}(\alpha,\varepsilon,\varphi,\Omega)$ is also a constant. If $T_{0}\leq M_{0}^{2}/C_{2}^{2}(M_{0}+1)^{2}$ then $\sup_{0\leq t\leq T_{0}}\|\nabla\Phi[w](t)\|_{p+1-\alpha}\leq 2M_{0}$. Thus we have $\|\Phi[w]\|_{Y}\leq 4M_{0}$ and $\Phi[w]\in Y$.
Let $w,y\in Y$ then we infer from $L^{p}$-$L^{q}$ estimates that
$\|\Phi[w](t)-\Phi[y](t)\|_{W^{1,p+1-\alpha}}$ $\leq\int_{0}^{t}(C^{\prime}+\tilde{C}(t-s)^{-\frac{1}{2}})\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|\nabla w(s)-\nabla y(s)\|_{p+1-\alpha}\,ds$
$\leq C_{3}(T_{0}+T_{0}^{\frac{1}{2}})\|w-y\|_{Y},$
where $C_{3}=C_{3}(\alpha,\varepsilon,\varphi,\Omega)$ is a constant. If $T_{0}\leq\min\{1/4C_{3},1/16C_{3}^{2}\}$ then $\|\Phi[w]-\Phi[y]\|_{Y}\leq 1/2\|w-y\|_{Y}$. Therefore choosing
$T_{0}\leq\min\left\{\frac{M_{0}}{C_{1}(M_{0}+1)},\frac{M_{0}^{2}}{C_{2}^{2}(M_{0}+1)^{2}},\frac{1}{4C_{3}},\frac{1}{16C_{3}^{2}}\right\},$
we see that $\Phi$ is a contraction map on $Y$. By Banach’s fixed point theorem, there is a function $w_{2}\in Y$ such that $\Phi[w_{2}]=w_{2}$. In other words, we obtain
$w_{2}(t)=e^{\delta t\Delta}v_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\left(\frac{\alpha\nabla w_{2}(s)}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}(s)+\varepsilon)^{\alpha}\Delta\varphi\right)\,ds.$ (A.2)
Step 2: Regularity and nonnegativity.
Set
$f(t)\coloneq\frac{\alpha\nabla w_{2}(t)}{(w_{1}(t)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}(t)+\varepsilon)^{\alpha}\Delta\varphi\quad\mathrm{for}\ t\in[0,T_{0}],$
then we have $f\in L^{\infty}((0,T_{0});L^{p+1-\alpha}(\Omega))$ because $w_{1},w_{2}\in C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$ and $\varphi$ belongs to $C^{\infty}(\bar{\Omega})$. Moreover, setting
$F(t)\coloneq\int_{0}^{t}e^{\delta(t-s)\Delta}f(s)\,ds\quad\mathrm{for}\ t\in[0,T_{0}],$
by [16, Lemma 7.1.1], we have $F\in C^{\frac{1}{2}}([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$. Since $e^{\delta t\Delta}w_{0}\in C^{\infty}(\Omega\times(0,\infty))$, we also obtain $w_{2}\in C^{\frac{1}{2}}((0,T_{0}];W^{1,p+1-\alpha}(\Omega))$. Then it holds $f\in C^{\frac{1}{2}}((0,T_{0}];L^{p+1-\alpha}(\Omega))$. Indeed, for $t,s\in(0,T_{0}]$, we have
$\|f(t)-f(s)\|_{p+1-\alpha}$
$\leq\left\|\frac{\alpha\nabla w_{2}(t)\cdot\nabla\varphi}{(w_{1}(t)+\varepsilon)^{1-\alpha}}-\frac{\alpha\nabla w_{2}(s)\cdot\nabla\varphi}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\right\|_{p+1-\alpha}+\left\|\left[(w_{1}(t)+\varepsilon)^{\alpha}-(w_{1}(s)+\varepsilon)^{\alpha}\right]\Delta\varphi\right\|_{p+1-\alpha}$
$\leq\frac{\alpha(1-\alpha)}{\varepsilon^{2-\alpha}}\|\nabla\varphi\|_{\infty}\|w_{2}(t)-w_{2}(s)\|_{W^{1,p+1-\alpha}}+\frac{\alpha}{\varepsilon^{1-\alpha}}\|\Delta\varphi\|_{\infty}\|w_{1}(t)-w_{1}(s)\|_{p+1-\alpha},$
where we used the mean value theorem as follows
$\left|\frac{\nabla w_{2}(t)}{(w_{1}(t)+\varepsilon)^{1-\alpha}}-\frac{\nabla w_{2}(s)}{(w_{1}(s)+\varepsilon)^{1-\alpha}}\right|\leq\frac{1-\alpha}{\varepsilon^{2-\alpha}}\left(|\nabla w_{2}(t)-\nabla w_{2}(s)|+|w_{1}(t)-w_{1}(s)|\right),$
$|(w_{1}(t)+\varepsilon)^{\alpha}-(w_{1}(s)+\varepsilon)^{\alpha}|\leq\frac{\alpha}{\varepsilon^{1-\alpha}}|w_{1}(t)-w_{1}(s)|.$
Here, since $w_{1},w_{2}\in C^{\frac{1}{2}}((0,T_{0}];W^{1,p+1-\alpha}(\Omega))$ we have $f\in C^{\frac{1}{2}}((0,T_{0}];L^{p+1-\alpha}(\Omega))$.
By [16, Theorem 4.3.4], we obtain $w_{2}\in C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega))$ and $w_{2}$ satisfies
$\begin{dcases}\partial_{t}w_{2}=\delta\Delta w_{2}+\frac{\alpha\nabla w_{2}}{(w_{1}+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi+(w_{1}+\varepsilon)^{\alpha}\Delta\varphi&\mathrm{a.e.\ in}\ \Omega\times(0,T_{0}),\\ (\nabla w_{2}\cdot\boldsymbol{n})|_{\partial\Omega}=0&\mathrm{for}\ t\in(0,T_{0}],\\ w_{2}(0)=w_{0}&\mathrm{in}\ W^{1,p+1-\alpha}(\Omega),\end{dcases}$ (A.3)
where $(\nabla w_{2}\cdot\boldsymbol{n})|_{\partial\Omega}$ is defined by the trace operator on $W^{2,p+1-\alpha}(\Omega)$ and the Neumann boundary condition is satisfies because of the definition of $w_{2}$ and the property of the Neumann heat semigroup.
We will show $w_{2}\geq 0$ a.e. in $\Omega\times[0,T_{0}]$. Let $w_{2}^{-}\coloneq\min\{w_{2},0\}$ and $t\in(0,T_{0}]$. Then we choose arbitary $a\in(0,t)$ and fix it. Multiplying the first equation of (A.3) by $w_{2}^{-}$ and integrating in $\Omega$, we have
$\frac{1}{2}\frac{d}{dt}\|w_{2}^{-}(t)\|_{2}^{2}=-\delta\|\nabla w_{2}^{-}(t)\|_{2}^{2}+\int_{\Omega}\frac{\alpha\nabla w_{2}^{-}(t)w_{2}^{-}(t)}{(w_{1}(t)+\varepsilon)^{1-\alpha}}\cdot\nabla\varphi\,dx+\int_{\Omega}(w_{1}(t)+\varepsilon)^{\alpha}w_{2}^{-}(t)\Delta\varphi\,dx,$
where we used integration by parts and the condition $(\nabla w_{2}\cdot\boldsymbol{n})|_{\partial\Omega}=0$ for $t\in(0,T_{0}]$. Note that thanks to $1+\alpha-2/d\leq p\leq 1+\alpha$ and the Sobolev embedding theorem, it holds that $W^{2,p+1-\alpha}(\Omega)\hookrightarrow H^{1}(\Omega)\hookrightarrow L^{\frac{p+1-\alpha}{p-\alpha}}(\Omega)$, thus the right hand side is well-defined. Since $w_{1}\geq 0$, it follows from Hölder’s inequality and (A) that
$\frac{1}{2}\frac{d}{dt}\|w_{2}^{-}(t)\|_{2}^{2}$
$\leq-\delta\|\nabla w_{2}^{-}(t)\|_{2}^{2}+\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|\nabla w_{2}^{-}(t)\|_{2}\|w_{2}^{-}(t)\|_{2}+\|(w_{1}(t)+\varepsilon)^{\alpha}\Delta\varphi\|_{2}\|w_{2}^{-}(t)\|_{2}$
$\leq-\delta\|\nabla w_{2}^{-}(t)\|_{2}^{2}+\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|\nabla w_{2}^{-}(t)\|_{2}\|w_{2}^{-}(t)\|_{2}+C(\|w_{1}(t)\|_{2}^{\alpha}+1)\|w_{2}^{-}(t)\|_{2}.$
Moreover by $\|w_{1}(t)\|_{2}\leq M_{0}$ and Young’s inequality, we have
$\left(\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}}\|\nabla w_{2}^{-}(t)\|_{2}+C(M_{0}^{\alpha}+1)\right)\|w_{2}^{-}(t)\|_{2}$ $\leq C^{\prime}(\|\nabla w_{2}^{-}(t)\|_{2}+1)\|w_{2}^{-}(t)\|_{2}$
$\leq\frac{\theta}{2}(\|\nabla w_{2}^{-}(t)\|_{2}+1)^{2}+\frac{(C^{\prime})^{2}}{2\theta}\|w_{2}^{-}(t)\|_{2}^{2}$
$\leq\theta(\|\nabla w_{2}^{-}(t)\|_{2}^{2}+1)+\frac{(C^{\prime})^{2}}{2\theta}\|w_{2}^{-}(t)\|_{2}^{2},$
where
$C^{\prime}\coloneq\max\left\{\frac{\alpha\|\nabla\varphi\|_{\infty}}{\varepsilon^{1-\alpha}},C(M_{0}^{\alpha}+1)\right\}$
and $\theta>0$ satisfies
$\theta<\min_{t\in[a,T_{0}]}\frac{\delta\|\nabla w_{2}^{-}(t)\|_{2}^{2}}{\|\nabla w_{2}^{-}(t)\|_{2}^{2}+1}.$
Note that if $\|\nabla w_{2}^{-}(t)\|_{2}=0$ then $w_{2}^{-}(t)=0$ a.e. in $\Omega$ obviously. Hence we have
$\frac{1}{2}\frac{d}{dt}\|w_{2}^{-}(t)\|_{2}^{2}\leq\frac{(C^{\prime})^{2}}{2\theta}\|w_{2}^{-}(t)\|_{2}^{2}.$
By Gronwall’s lemma and $w_{0}\geq 0$ a.e. in $\Omega\times[0,T_{0}]$, it follows
$\|w_{2}^{-}(t)\|_{2}=0\quad\mathrm{for}\ t\in[a,T_{0}],$
then
$w_{2}^{-}(t)=0\quad\mathrm{a.e.\ in}\ \Omega,\ \mathrm{for}\ t\in[a,T_{0}].$
Since $a>0$ is arbitary, we see $w_{2}\geq 0$ a.e. in $\Omega\times(0,T_{0}]$, that is, $w_{2}\geq 0$ a.e. in $\Omega\times[0,T_{0}]$.
Step 3: Convergences and properties.
Repeating the above process, we can construct a sequence $\{w_{k}\}$ such that
$w_{k}\in Y\cap C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega)),$
$w_{k}\geq 0\ \mathrm{a.e.\ in}\ \Omega\times[0,T_{0}].$
We will show that $\{w_{k}\}_{k\geq 2}$ is a Cauchy sequence in $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$ if $T_{0}$ is small enough. First, for $t\in[0,T_{0}]$ by $L^{p}$-$L^{q}$ estimates and the mean value theorem, we have
$\|w_{k+1}(t)-w_{k}(t)\|_{p+1-\alpha}$
$\leq\int_{0}^{t}\left\|e^{\delta(t-s)\Delta}\left\{\alpha\left(\frac{\nabla w_{k+1}(s)}{(w_{k}(s)+\varepsilon)^{1-\alpha}}-\frac{\nabla w_{k}(s)}{(w_{k-1}(s)+\varepsilon)^{1-\alpha}}\right)\cdot\nabla\varphi\right\}\right\|_{p+1-\alpha}\,ds$
$\quad+\int_{0}^{t}\left\|e^{\delta(t-s)\Delta}\left\{((w_{k}(s)+\varepsilon)^{\alpha}-(w_{k-1}(s)+\varepsilon)^{\alpha})\Delta\varphi\right\}\right\|_{p+1-\alpha}\,ds$
$\leq\int_{0}^{t}C^{\prime}\alpha\|\nabla\varphi\|_{\infty}\left\|\frac{\nabla w_{k+1}(s)}{(w_{k}(s)+\varepsilon)^{1-\alpha}}-\frac{\nabla w_{k}(s)}{(w_{k-1}(s)+\varepsilon)^{1-\alpha}}\right\|_{p+1-\alpha}\,ds$
$\quad+\int_{0}^{t}C^{\prime}\|\Delta\varphi\|_{\infty}\|(w_{k}(s)+\varepsilon)^{\alpha}-(w_{k-1}(s)+\varepsilon)^{\alpha}\|_{p+1-\alpha}\,ds$
$\leq\int_{0}^{t}C_{4}\left(\|\nabla w_{k+1}(s)-\nabla w_{k}(s)\|_{p+1-\alpha}+\|w_{k}(s)-w_{k-1}(s)\|_{p+1-\alpha}\right)\,ds$
$\leq C_{4}T_{0}\left(\sup_{0\leq s\leq T_{0}}\|w_{k+1}(s)-w_{k}(s)\|_{W^{1,p+1-\alpha}}+\sup_{0\leq s\leq T_{0}}\|w_{k}(s)-w_{k-1}(s)\|_{W^{1,p+1-\alpha}}\right),$
where $C_{4}=C_{4}(\alpha,\varepsilon,\varphi,\Omega)$ is a constant. Similarly, it follows
$\|\nabla w_{k+1}(t)-\nabla w_{k}(t)\|_{p+1-\alpha}$
$\leq\int_{0}^{t}\left\|\nabla e^{\delta(t-s)\Delta}\left\{\alpha\left(\frac{\nabla w_{k+1}(s)}{(w_{k}(s)+\varepsilon)^{1-\alpha}}-\frac{\nabla w_{k}(s)}{(w_{k-1}(s)+\varepsilon)^{1-\alpha}}\right)\nabla\varphi\right\}\right\|_{p+1-\alpha}\,ds$
$\quad+\int_{0}^{t}\left\|\nabla e^{\delta(t-s)\Delta}\left\{((w_{k}(s)+\varepsilon)^{\alpha}-(w_{k-1}(s)+\varepsilon)^{\alpha})\Delta\varphi\right\}\right\|_{p+1-\alpha}\,ds$
$\leq C_{5}T_{0}^{\frac{1}{2}}\left(\sup_{0\leq s\leq T_{0}}\|w_{k+1}(s)-w_{k}(s)\|_{W^{1,p+1-\alpha}}+\sup_{0\leq s\leq T_{0}}\|w_{k}(s)-w_{k-1}(s)\|_{W^{1,p+1-\alpha}}\right),$
where $C_{5}=C_{5}(\alpha,\varepsilon,\varphi,\Omega)$ is a constant. Hence, choosing $T_{0}$ such that
$T_{0}<\min\left\{\frac{1}{4C_{4}},\frac{1}{16C_{5}^{2}}\right\},$
we have
$\sup_{0\leq t\leq T_{0}}\|w_{k+1}(t)-w_{k}(t)\|_{W^{1,p+1-\alpha}}\leq\frac{1}{2}\sup_{0\leq t\leq T_{0}}\|w_{k}(t)-w_{k-1}(t)\|_{W^{1,p+1-\alpha}}.$
Thus, for $m,n\in\mathbb{N}$ with $n>m$, it follows
$\sup_{0\leq t\leq T_{0}}\|w_{n}(t)-w_{m}(t)\|_{W^{1,p+1-\alpha}}$ $\leq\sum_{k=m}^{n-1}\sup_{0\leq t\leq T_{0}}\|w_{k+1}(t)-w_{k}(t)\|_{W^{1,p+1-\alpha}}$
$\leq\sum_{k=m}^{n-1}\left(\frac{1}{2}\right)^{k}\left(\sup_{0\leq t\leq T_{0}}\|w_{1}(t)\|_{W^{1,p+1-\alpha}}+\|w_{0}\|_{W^{1,p+1-\alpha}}\right)$
$\leq\left(\frac{1}{2}\right)^{m-1}5M_{0}\ \rightarrow 0\quad\mathrm{as}\ n,m\rightarrow\infty.$
Since $(C([0,T_{0}];W^{1,p+1-\alpha}(\Omega)),\sup_{0\leq t\leq T_{0}}\|\cdot\|_{W^{1,p+1-\alpha}})$ is complete, there exists a function $w\in C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$ such that $w_{k}\rightarrow w$ in $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$. In addition, we see $w\geq 0$ a.e. in $\Omega\times[0,T_{0}]$. Indeed, for all $\psi\in C_{c}^{\infty}(\Omega\times[0,T_{0}])$ with $\psi\geq 0$, we have
$\left|\int_{0}^{T_{0}}\int_{\Omega}w_{k}(x,t)\psi(x,t)\,dx\,dt-\int_{0}^{T_{0}}\int_{\Omega}w(x,t)\psi(x,t)\,dx\,dt\right|$
$\leq\int_{0}^{T_{0}}\|w_{k}(t)-w(t)\|_{p+1-\alpha}\|\psi(t)\|_{\frac{p+1-\alpha}{p-\alpha}}\,dt$
$\leq\int_{0}^{T_{0}}\|\psi(t)\|_{\frac{p+1-\alpha}{p-\alpha}}\,dt\sup_{0\leq t\leq T_{0}}\|w_{k}(t)-w(t)\|_{p+1-\alpha}\to 0\quad\mathrm{as}\ k\to\infty.$
Hence we obtain
$0\leq\int_{0}^{T_{0}}\int_{\Omega}w_{k}(x,t)\psi(x,t)\,dx\,dt\to\int_{0}^{T_{0}}\int_{\Omega}w(x,t)\psi(x,t)\,dx\,dt\quad\mathrm{as}\ k\to\infty$
$\hskip 199.16928pt\forall\psi\in C_{c}^{\infty}(\mathbb{R}^{d}\times[0,T_{0}])\ \mathrm{with}\ \psi\geq 0,$
then
$w(x,t)\geq 0\ \mathrm{a.e.\ in}\ \Omega\times[0,T_{0}].$
Next we will prove that $w$ satisfies
$w(t)$ $=e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\left(\frac{\alpha\nabla w(s)}{(w(s)+\varepsilon)^{1-\alpha}}\nabla\varphi+(w(s)+\varepsilon)^{\alpha}\Delta\varphi\right)\,ds$
$=e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\{\nabla\cdot((w(s)+\varepsilon)^{\alpha}\nabla\varphi)\}\,ds.$
By setting
$W(t)\coloneq e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\left(\frac{\alpha\nabla w(s)}{(w(s)+\varepsilon)^{1-\alpha}}\nabla\varphi+(w(s)+\varepsilon)^{\alpha}\Delta\varphi\right)\,ds\quad t\in[0,T_{0}],$
it follows from $L^{p}$-$L^{q}$ estimates and the mean value theorem that
$\|W(t)-w_{k}(t)\|_{W^{1,p+1-\alpha}}$
$\leq\int_{0}^{t}(C^{\prime}+\tilde{C}(t-s)^{-\frac{1}{2}})\left\|\frac{\nabla w(s)}{(w(s)+\varepsilon)^{1-\alpha}}-\frac{\nabla w_{k}(s)}{(w_{k-1}(s)+\varepsilon)^{1-\alpha}}\right\|_{p+1-\alpha}\,ds$
$\quad+\int_{0}^{t}(C^{\prime}+\tilde{C}(t-s)^{-\frac{1}{2}})\|w(s)-w_{k-1}(s)\|_{p+1-\alpha}\,ds$
$\leq C_{6}(T_{0}+T_{0}^{\frac{1}{2}})\left(\sup_{0\leq s\leq T_{0}}\|\nabla w(s)-\nabla w_{k}(s)\|_{p+1-\alpha}+\sup_{0\leq s\leq T_{0}}\|w(s)-w_{k-1}(s)\|_{p+1-\alpha}\right),$
where $C_{6}=C_{6}(\alpha,\varepsilon,\varphi,\Omega)$ is a constant. Hence we have $w_{k}\rightarrow W$ in $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$. Due to the uniquness of limit, it follows $w=W$ in $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$. Moreover, we see $w\in C([0,T_{0}];L^{1}\cap L^{2}(\Omega))$. Indeed, for $t,s\in[0,T_{0}]$, by Hölder’s inequality, we have
$\|w(t)-w(s)\|_{1}\leq|\Omega|^{\frac{p+1-\alpha}{p-\alpha}}\|w(t)-w(s)\|_{p+1-\alpha}.$
On the other hand, by the Sobolev embedding theorem, it follows
$\|w(t)-w(s)\|_{2}\leq C_{s}\|w(t)-w(s)\|_{W^{1,p+1-\alpha}},$
where $C_{s}$ is a constant. Since $w\in C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))$, we obtain $w\in C([0,T_{0}];L^{1}\cap L^{2}(\Omega))$. Adapting the same argument for $w_{2}$ in Step 2, we have $w\in C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega))$ and $w$ satisfies
$\begin{cases}\partial_{t}w=\delta\Delta w+\nabla\cdot((w+\varepsilon)^{\alpha}\nabla\varphi)&\mathrm{a.e.\ in}\ \Omega\times(0,T_{0}),\\ (\nabla w\cdot\boldsymbol{n})|_{\partial\Omega}=0&\mathrm{for}\ t\in(0,T_{0}],\\ w(0)=w_{0}&\mathrm{in}\ L^{1}\cap L^{2}\cap W^{1,p+1-\alpha}(\Omega).\end{cases}$ (A.4)
Now, we will show $\|w(t)\|_{1}=\|w_{0}\|$ for $t\in[0,T_{0}]$. Let $t\in[0,T_{0}]$. Integrating the first equation of (A.4) in $\Omega$, we have
$\frac{d}{dt}\int_{\Omega}w(t)\,dx$ $=\delta\int_{\Omega}\Delta w(t)\,dx+\int_{\Omega}\nabla\cdot((w(t)+\varepsilon)^{\alpha}\nabla\varphi)\,dx.$
Since $\nabla w\cdot\boldsymbol{n}=0$ and $\nabla\varphi\cdot\boldsymbol{n}=0$ on $\partial\Omega\times(0,T_{0}]$, we infer from integration by parts that
$\frac{d}{dt}\int_{\Omega}w(t)\,dx$ $=\delta\int_{\partial\Omega}\nabla w(t)\cdot\boldsymbol{n}\,dS+\int_{\Omega}(w(t)+\varepsilon)^{\alpha}\nabla\varphi\cdot\boldsymbol{n}\,dS=0.$
By integrating over $[0,t]$, it follows
$\int_{\Omega}w(t)\,dx=\int_{\Omega}w(0)\,dx=\int_{\Omega}w_{0}\,dx.$
Step 4: Uniqueness.
Let $y_{1},y_{2}\in C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))\cap C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega))$ be solutions to (A.4). Then by [16, Proposition 4.1.2], they are mild solutions to (A.4):
$y_{1}(t)=e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\{\nabla\cdot((y_{1}(s)+\varepsilon)^{\alpha}\nabla\varphi)\}\,ds,$
$y_{2}(t)=e^{\delta t\Delta}w_{0}+\int_{0}^{t}e^{\delta(t-s)\Delta}\{\nabla\cdot((y_{2}(s)+\varepsilon)^{\alpha}\nabla\varphi)\}\,ds.$
Define
$t_{0}\coloneq\max\{t\in[0,T_{0}];y_{1}(s)=y_{2}(s)\ \mathrm{for}\ 0\leq s\leq t\},$
and set $y_{0}\coloneq y_{1}(t_{0})=y_{2}(t_{0})$. If $t_{0}<T_{0}$, the problem
$\partial_{t}w(t)=\delta\Delta w(t)+\nabla\cdot((w(t)+\varepsilon)^{\alpha}\nabla\varphi),\ t>t_{0},\ w(t_{0})=y_{0},$ (A.5)
has a unique mild solution in a set
$Y^{\prime}=\left\{y\in C([t_{0},t_{0}+a];W^{1,p+1-\alpha}(\Omega));\sup_{t_{0}\leq t\leq t_{0}+a}\|y(t)\|_{W^{1,p+1-\alpha}}\leq R\right\},$
provided $R$ is large enough and $a$ is small enough. Since $y_{1}$ and $y_{2}$ are bounded with value in $W^{1,p+1-\alpha}(\Omega)$, there exists $R$ such that $\|y_{i}(t)\|_{W^{1,p+1-\alpha}}\leq R$ for $t_{0}\leq t\leq T_{0},i=1,2$. Thus it follows $y_{1}=y_{2}$ in $Y^{\prime}$. On the other hand, $y_{1}$ and $y_{2}$ are two different solutions of (A.4) in $t_{0},t_{0}+a]$, for every $a\in(0,T_{0}-t_{0}]$. This is a contradiction. Hence $t_{0}=T_{0}$, and the solution of (A.4) is unique in $C([0,T_{0}];W^{1,p+1-\alpha}(\Omega))\cap C((0,T_{0}];W^{2,p+1-\alpha}(\Omega))\cap C^{1}((0,T_{0}];L^{p+1-\alpha}(\Omega))$. The proof is completed. ∎
Acknowledgements
The author is greatly indebted to Professor Kentaro Fujie for his helpful comments during the preparation of the paper.